A Complete Bibliography of Publications in Mathematics of Computation, 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

    E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
               WWW URL: https://www.math.utah.edu/~beebe/

22 September 2025 Version 1.17

Title word cross-reference

(0, 2) [HMXY23]. 1 [KMNN21]. 1.5 [LNW25]. 2 [BBL20, FS20]. 2^{n} + p [DDDR20]. 2^{t} [MS20b]. 2m [CH20]. 3 [BBL20, KK21]. 4 [FS24]. 5 [KK21]. α [SSX26]. B [LL20]. C^{0} [Ley21]. C_{4} × C_{4} × C_{4} [BBERPC24]. D_{4} [ABG^{+}24]. E [BRS24]. E_{2} [GPS24]. ℓ_{p} [TJW25]. G(p^{f}) [LMP20]. Γ [MS20a]. H(div) [EV20]. H^{1} [EV20, HHZ23, WH24]. H^{2} [WH24]. hp [BCFM25, CDG22]. j [BHK^{+}25b]. k [MOT21]. L [AFPwaabAVS22, BMOR21, DMV23, GdLL25, LV22]. L^{2} [GK26]. L^{∞}(L^{2}) [WY22a]. λ [BS23, DK23]. 𝔽_{2} [AKPW25]. ℚ [PT22]. ℝ [LL20]. ℝ^{d} [GKS22, NS23]. ℝ^{n} [CH20]. 𝕋 [WY22b]. ℤ_{2} [Pag22]. H(curl) [CFEV22]. P_{2} − P_{1} [ORZ21]. H^{2} [Bör25]. FI [WG20]. N [Fre23]. \overline{M}_{0,6} [BKR20]. P [DMM24, BCM21, BLN^{+}24, CFEV22, GM25a, LV22, LMP20, MR20, MT20, MR24, PS21, SiTiY20, WD24]. π(N) [HKM25]. ψ(x) − x [GKPR25]. q [QH22]. qd [Hua20]. θ(x) [BKL^{+}21]. Õ(\sqrt{N}) [HKM25]. U_{n} − b^{m} = c [HTVZ23]. x^{3} + y^{3} + 1 − kxy [TG25]. X_{0}(N)

[AKMJ^{+}24, BKS23, NO24].

-action [ABG^{+}24]. -adic [BCM21, BLN^{+}24, GM25a, LV22, MT20, MR24, SiTiY20, WD24]. -algebras [DMV23]. -analogue [QH22]. -convergence [MS20a]. -Curves [PT22]. -extensions [Pag22]. -free [MOT21]. -functions [BMOR21, BRS24, GdLL25, LV22]. -groups [PS21]. -invariants [BHK^{+}25b, DK23]. -matrices [Bör25]. -module [WG20]. -numbers [GPS24]. -Poisson [BS23]. -polynomials [AFPwaabAVS22]. -recursive [DMM24]. -robust [CFEV22]. -schemes [LL20]. -stable [SSX26]. -subgroups [LMP20]. -type [Hua20, GK26]. -version [CDG22, MR20].

120-cell [MZ21]. 1D [WLL21, WZZ22].

2-torsion [HSS21]. 2D [MIJR21].

3-term [GKW24]. 30750-bit [GKL^{+}21]. 3D [ST23].

4-manifolds [MZ21].

5-chromatic [BBKP24].

Abel [AAP22]. abelian [CCG20, DK23, Dem22, FR22, HS21a, Mar21]. abstract [EG24]. acceleration [CT22a, LZ25, Rie23]. Accuracy [DSW22, HS21b, HS25b, Hua20, Sun25, PO23, PO24]. Accurate [BPT21, CLT22, DLL21]. achieving [DGS22, NS23]. acoustics [BLM25]. action [ABG^{+}24]. actions [FJRR21]. active [GG24]. adapted [JZCH24]. Adaptive [HW20, HPR24, Bör25, Car20, DGM20, FMP21, GHPS21, HXDM24, KG21, Li21, lLZ21]. additive [AHPS20, BBERPC24, GGS23]. ADI [BDF21]. adic [BCM21, BLN^{+}24, GM25a, LV22, MT20, MR24, SiTiY20, WD24]. ADMM [ZWZ20]. ADRT [LRR23]. advection [FM20]. affects [Sun25]. affine [BLH22]. after [FS21]. agent [CEHT23]. agglomeration [AHPS20]. agglomeration-based [AHPS20]. aggregation [CFS21]. Alfeld [CZ25, FGN20]. Algebra [CDLT25, ZL20]. algebraic [AMP22, BRS24, Fla21, Rüd22, Sgo23, WWW21]. algebraically [JKK21]. algebras [BGO24, CEFO20, CSS24, DMV23]. algorithm [Bla22, FP22, FP25, GCL23, Har21, LR21, LS20a, LdCS^{+}21, MM21, MT20, MR24, OR23, TJW25, ZC21]. algorithmic [BLH22, SY23]. Algorithms [CL26, CC25, HB22, PS21, BHK^{+}25a, BST21, BEL25, BV24, BP24, CKNS21, HRR25, JCWH21, KL23, SiTiY20, SLN23, SW20, WD24]. Aliasing [Gos20]. Allen [HJQ23, HS23]. almost [DGS22, WZSP21]. along [HSS21]. alternating [GSW23, JZCH24]. amongst [Ros22]. Ampère [GT23, QT20]. amplification [AGY21]. analog [Fla20]. analogue [QH22]. analyses [KBL21]. Analysis [ACS21, DSS23, FRS25, FSTW25, HHZ23, KHR23, LS24, LvBV25, LZMW20, lLZ21, Reu25, Not20, Reu22, YEHS25, BBP24, BM22, BRT23, BCFM25, BDF21, BLM23, BDM24, CLY23, CNV21, CH24a, CHW20, DM24, DL22, DGSW22, ELM^{+}24, EPR21, GM24, GK26, GS20, GWWZ24, HW21, HXDM24, HLX22, HKLP23, HS23, Kop21, LS23, LQW21, LSL22, LQT22, LLZ23, LMS25, LL24, Wan25, MIJR21, Nic24, NS23, PYAT21, SW21, VSB26, WZZ22, Wan24, XLL24, YLLO24]. analytic [SH25, Wan24]. analytical [DS20a]. analyticity [KK20]. Analyzing [WpL20]. anchoring [HXDM24]. anisothermal [FOS22]. anisotropic [BL25b, BBM21]. annealing [BP24]. ANOVA [GKS22, KPW21]. Anti [Not22]. Anti-Gaussian [Not22]. AP [BRS23]. application [DD23, KR25, WGMZ25]. applications [CH25b, CEHT23, CX25, DS20a, DOS20, GPW21, HH20, HKLP23, LR21, LLZ23, PS21, SLN23, ZWZ20, ZC21]. applied [DES22, GAHW21]. approach [BLH22, EG24, FMZ23, HKM25, JST22, KZ22, LZ25, Sog20, SY23, SJ24, Zim21]. Approximate [JCWH21, Nak20]. Approximating [FLMSS22]. Approximation [ACEL21, AK24, CX25, MKL21, AL20, ABP25, BDOT24, BHK^{+}25a, BR21, BKRS20, Ben22, BKKS24, BNPP20, BBL20, CM23, CKNS21, EG24, EGH25, FOS22, GN25, GHS23, KNZ25, KMNN21, Ley21, Tra25, Sog20, WZZ22, YEHS25, ZYYZ25]. Approximations [DMM20, AV25, AP24, AE25, BEKU21, BM22, BAL21, BHL24, CNV21, DM24, MS22, SLN23, Wan24]. arbitrarily [CDG22, ZL20]. arbitrarily-shaped [CDG22]. arbitrary [FM20, FS21, GN25, Han23, LMP20]. arbitrary-order [Han23]. archimedean [MX26]. arising [Fre23]. arithmetic [AF20, CJdL21, EHP22, FKRS21, GKW24, LM26, Ric22, vBHM20]. arithmeticity [BKK24]. aromas [BCM^{+}24]. aromatic [LV20]. arrangements [BK23a]. artificial [PYAT21]. assimilation [BFO20, DSW22]. assisted [CSS24]. associated [KBL21]. assuming [EHP22]. Asymptotic [DOP23, HM22, CL25, DMM24, LMY21, SW25]. Asymptotics [CT22b]. augmented [HMXY23, LDHS23, WX22]. automorphisms [OS21]. Average [FS20, lLW25, FJRR21]. AVF [CHS22]. avoidable [Ros20]. Avoiding [Ros22, BR21]. aware [GS23]. away [Ros20].

B [LV20]. B-series [LV20]. Backward [De 20, ABD20, HS21b, HPW20]. Baillie [BFW21]. balance [BAL21]. balanced [CWS26, DOP23, Gos20]. balanced-norm [CWS26]. band [CHV24]. barotropic [HLS21]. Barrett [BL25a]. barrier [GM25b]. base [GJN20]. based [AB20, AHPS20, DGSW22, GMS21, HPR24, HLLM25, MR20, TJW25]. bases [BSW23]. basis [BK24, HR25, HP21]. Baxter [AMV22]. Bayes’ [SJ24, COS21]. Bayesian [HKS21, ZCL^{+}22]. BDDC [WZSP21]. BDF [HS21b, HJQ23]. BDF2 [GNT24, lLZ21]. be [BBGS22, Ros20]. beams [LRY20]. become [FS21]. behavior [WG20]. behaviors [LL20]. bending [ABP25, FH25]. Bernstein [SWZ21b]. Besov [HKS21]. best [BDM25]. between [BLH22, GKS22, HJ20, Mil20]. Bézout [GLNY20]. BGG [BHKS25]. Bielliptic [BKS23, NV23]. biharmonic [CX24, CH25b, Ley21, ZMZW23]. binary [GKL^{+}21]. biomembranes [CYB^{+}21]. Biot [GCL23, KZ22, KMRB25]. birational [BGMS23]. Birch [KMwaabAVS23]. Bit [HG26, GKL^{+}21]. blending [FW23]. blob [CEHT23]. Block [WZSP21, HLLM25]. Bonse [MT22]. Bonse-type [MT22]. Boolean [ZL20]. Boosted [HNP22]. Borcherds [PSY20]. Bose [HY25a]. both [BLM24, HM22]. bottlenecks [DEG22]. bound [AMP22, Doe22, EHP22, GM25b, HJQ23]. boundary [ASZ23, ACS21, Ben22, BBV21, BHL24, Col20, DLL21, FSTW25, GS22, KR25, LNW25, MIJR21, Nic24, PYAT21, RSM22, WZZ22]. bounded [BNWX23]. Bounds [BDH^{+}21, Avd21, BRS23, BFS22, BCF23, BKL^{+}21, CDM^{+}20, CT22c, DMM24, EGH25, FMS24, Fla21, Gal23, GG24, HR25, KZ23, LP25, Nak20, OSS25, SH22, SH25]. Boussinesq [JLOS25]. BPX [BNWX23]. braces [BBERPC24]. bracketing [GPW21]. branching [HS25a]. Breaking [GM25b]. Bregman [ZZY23]. broken [CFEV22, EV20]. Brownian [LX25]. Bruijn [AB20]. BSDEs [CT20]. Butson [LÖS20].

Cahn [BV23, GWWZ24, HJS24, HJQ23, HS23, LQW21, LQT22, Li22b]. calculation [CHV24]. calculations [GPS24, KL23]. calculus [AFGS23, HLX22, LdCS^{+}21]. Calderón [GH22]. camera [CDLT25]. can [BBGS22]. cardinality [PP20]. Carlo [HKS21, BRK25, CLY23, HW21, KPW21, MS25b, SSX26]. Cartan [FJRR21, Chu25, MM21, MS20c]. Cartier [AFPwaabAVS22]. case [CWS26, FP25, Not20].

Cauchy [KR25]. cell [MZ21]. Cellular [AE25]. census [AKPW25]. centered [SJ24]. certain [Rüd22]. Certified [ZCL^{+}22]. chamber [BKR20]. change [GJN20]. changes [GKPR25]. changing [FS21]. character [MS25b]. characteristic [CL26, WDM20]. characterization [LRR23]. Characterizing [XXX25]. characters [LMP20, PT22]. Chebyshev [BKL^{+}21, Not22]. chemotaxis [WLF25]. Chevalley [BLH22, LMP20]. Cholesky [COS25]. chromatic [BBKP24]. circle [QESS24]. Class [BK22, Bar21, CLT22, EH20, GM21, GKMV25, GM25b, HH20, HLMT21, JST22, KNRR22, KK21, LV20, MT22, TG25]. classes [Wal99, Wal21]. classical [RS23]. classification [BS22, BHK^{+}25b, Rak24]. Clenshaw [WGMZ25]. Clifford [BJT20]. clusters [CDM^{+}20, HJLZ20]. CM [OS21, TG25]. cochain [FW23]. codes [BDH^{+}21, BGPW21]. coefficients [CM23, GS20, IGMS23]. Coleman [BT20]. Collective [Car20]. colliding [AGY21]. color [QMSZ25]. common [QESS24]. commuting [MR20]. compact [BJT20, BKKS24]. companion [De 20]. complementarity [ZC21]. complete [GAHW21, HLR23, IL23]. Completely [AB20, AK24]. completion [QMSZ25]. complex [CH22a, DD23, DIS23, Han23, LPPV24, VSB26, ZL20]. complexes [CH24b, HL21]. Complexity [Wan25, Bar20, BKM20, CT20, HG26, PP20]. Component [DGS22, CKNS21, God26]. Component-by-component [DGS22, CKNS21, God26]. composite [FS21, RSM22, WW24, ZWZ20]. composition [BCCET22, JW25]. compressible [BLMM^{+}23, HLS21]. Computation [CS23, GKL^{+}21, MS25a, BMR24, BKM20, BKS25, ERS21, FW20, GM24, HH20, LX25, MR23, MX26, Ric22, vBHM20]. Computational [PV20, GHPS21, Chu25]. compute [LR21, MM21]. Computing [AKMJ^{+}24, AFPwaabAVS22, BK23a, BKR20, DMM24, GM25a, GP24, GCO22, HKM25, HJ20, LV22, Mar21, Mil20, RS24, CL26, CC25, DDDR20, Hua20, JST22, LS20a, WG20]. Concurrent [vLW23]. condensates [HY25a]. condition [EPR21, Li25, MIJR21, XXX25, ZZY23]. Conditional [DGS26]. conditioned [BBGS22].

Conditioning [ABD20, De 20]. conditions [BBV21, BK22, BHL24, Col20, FSTW25, GS22, GGS23, HLZ23, Nic24, PYAT21, Sgo23, WZZ22]. Conforming [HL21, COSS26, CYB^{+}21, CH22b, CH25b, HHZ23, CYB^{+}21]. congruence [BK22]. Congruences [Fre23]. congruential [MS20b]. conjecture [BK23a, CMR25, CGT25, KMwaabAVS23, Pag22, Pul20, QH22]. conjugacy [BMP21]. conjugate [BCCET22]. connections [GK26]. consecutive [JT22]. conservation [AV25, DM24, HJLW20, PSW25]. conservative [BLM25]. Conservativity [HLMT21]. conserving [KHR23, WLF25]. Consistency [BAL21, COS21, ABP25, HLMT21]. consistent [GGKO23, KL23]. constant [AL20, Avd21, DK21, FM20]. constants [LM26]. constituted [FOS22]. constrained [JW25]. constraint [ABP25]. constraints [ACEL21, BLM24, EH20, GAHW21, LDHS23, MKL21]. Constructing [FR22, Kho20]. Construction [BRT23, FW23, Ula24, BS23, CKNS21, DGS22, KBL21, LLZ23, God26]. contact [BDM24]. continua [ZL20]. continued [BST21, DKS25, MR24, SiTiY20, WD24]. continuous [CH25a, GKW24, GHS23, LS21, MT22]. Contractivity [GDST26]. contrast [PV20]. control [CT20, CEHT23]. controlled [DSW22]. controlling [Li21]. convected [BRT23]. convection [CHEF24, CX24, CJS23, FP20, GLFN20, WLSZ23, YCQ20]. convection-diffusion [CX24, GLFN20]. Convergence [AV25, ASZ23, BL25a, BGH^{+}23, BDF21, BL21, BP24, CNV21, CS25, CT22a, CHW20, DGSW22, FP25, GT23, GGS23, GWWZ24, HW21, HXDM24, KÇR23, KG21, LQW21, Wan24, WD24, AL20, BCFM25, BEL25, BBDM26, FMP21, HJS24, HS23, Lav21, LMY21, LQT22, LL24, MLR22, NS23, PYAT21, SH21, SWZ21b, YCQ20, YLLO24, MS20a]. Convergent [KZ23, LD20, LWW^{+}21, LPP21, RSS23, ZWZ20]. convergents [DKS25]. converges [Zen23]. convex [Bar21, BK23b, BDF21, CHW20, GG24, HMXY23, LL21, LN22, LZ25, ZWZ20, ZYYZ25, Zha22]. convexity [ZZY23]. convolution [RSM22]. coordinate [WpL20]. corners [ACS21]. correction [GO20, LSL22, MLR22]. Correlation [QH22, CMR25]. correlations [GdLL25]. Corrigendum [CHL24, KG21, Wal21]. costs [GHPS21]. count [FS20]. Counting [BMOR21, HSW22, KS20, GEL24, KRRZ20]. coupled [AMU21]. Coupler [GEL24]. Coupling [HLZ23]. covers [GPW21]. Coxeter [KS20]. Crank [BBV21]. criterion [Sao22]. Crouzeix [CS22, HM22, ST23]. Crouzeix’s [CGT25]. crystalline [CHV24]. Cubature [CT20]. cubic [AKPW25, Wan25]. cubic-regularized [Wan25]. curl [CFEV22, CF23]. Curtis [WGMZ25]. curvature [DK21, EGL24, GN25]. curve [BL25a, FS24, HS25a, KNRR22, NO24]. Curves [PT22, AAP22, AKMJ^{+}24, ABG^{+}24, AFPwaabAVS22, BT20, BNP24, BS22, BKS23, BR21, Box21, BDS20, DGJ20, Fre23, GM25a, GJN20, GEL24, HSS21, HJ24, IL23, Kay22, Mil20, NV23, OS21, Rak24]. curvilinear [LQXY24]. cusp [PSY20]. CutFEM [BHL24]. Cutting [Pul20]. cycles [GMZ20, GJL^{+}24]. Cyclic [BNP24, WpL20]. cyclotomic [BK24, Pag22].

damping [BBP24, KR22]. Darmon [BCDF25]. data [BFO20, DSW22, FKRT24, LS25]. database [ELM^{+}24]. datum [CIK^{+}25]. Davidson [LX22]. Decomposition [WDM20, BKR20, GHM24, Lap20, Chu25]. decompositions [AC22, Füh22, Zim21]. decoupled [GCL23]. decrease [HRR25]. Dedekind [HSW22]. Deep [HPW20, DM24, FMZ23]. defined [LNW25]. definite [MM21]. degenerate [CNV21]. Degree [AMP22, CF23, EV20, FM20, FS20, Sun25, vLW23]. degree- [FS20]. degrees [GJL^{+}24]. del [vLW23]. Delay [WY22a]. Delay-dependent [WY22a]. delta [Avd21]. Deninger [LR21]. Density [CHS22, HJS24, BGH^{+}23, DDDR20]. dependent [BEL25, FMZ23, FSTW25, KBPS21, KHR23, WY22a]. depth [BBK^{+}24]. derivative [DES22, HRR25, JST22]. derivative-free [HRR25]. derivatives [GM24, TXS20]. descendant [BAFG20]. descent [CS25, JZCH24, LM22, WpL20, CHW20]. descents [Cre20, KS20]. design [BLM23]. desingularization [FKRS21]. detection [BS24, QESS24]. determinants [WY23]. deterministic [Har21, HH22, Hit21, Kir20]. developed [WLL21]. devising [DS21]. DG [IGMS23]. diagonal [AE25, Ula24]. diagrams [BLM23, CHV24]. difference [Ben22, BDF21, Coe24, HJS24, Sog20]. differences [BBM21]. Differential [CH20, AGdS22, BCCET22, BKKS24, BBM21, BRS24, DE25, FMZ23, GAHW21, KNZ25, LMY21, LZ25, SJ24]. differentiation [HS21b]. diffraction [QESS24]. diffusion [BL25b, BCHR20, CHEF24, CX24, CJS23, CWS26, CEHT23, Gos20, GLFN20, HLZZ21, lLZ21, OW20, WLSZ23]. diffusion-reaction [HLZZ21]. digit [FS21]. dimension [BBL20, BGMS23, DE25, GN25, YCQ20, ZCL^{+}22]. dimensional [BDM24, CX24, CWS26, Dua21, GT23, Gos20, HPW20, PO23, WZSP21, WGMZ25, ZL20, dCIDV24]. dimensions [CH22b, CH22a, CH24b, EV20, HHY20, HL21, LS25, GSY23]. diminishing [FTY24]. Diophantine [HTVZ23, KLM20, PT22]. Dirac [BCY20, SWZ21a]. Dirichlet [BMOR21, BBV21]. discontinuous [AHPS20, Bar21, CDG22, CH24a, CX24, CJS23, FSTW25, GKMV25, KHR23, KÇR23, KG21, LZMW20, Sun25, MLR22, PSW25, SH21, SX21, TXS20, TCLS25, WLSZ23, YCQ20]. discrepancy [Doe22, GPW21]. Discrete [BHKS25, BF24, DD23, FP20, GKL^{+}21, GKW24, Han23, HY25a, HJS24, HL21, HS23, LL21, LQXY24, MS22, RS23, SH21, TFD25, WY22a]. discrete-ordinate [SH21]. discretised [GGS23]. Discretization [MS20a, AFKL21, AMU21, BBM21, DL22, GT23, GGKO23, GGO20, KZ22, LNW25]. discretizations [BCFM25, CH24a, CDZ22, DSS23, Lav21, WLSZ23, WZSP21]. discretized [CLLZ25]. discriminants [Bar20]. disk [HLZZ21, HJ21]. dispersive [CRS22]. dissipation [lLW25]. dissipative [CLT22]. distance [DGS26]. distributed [AB20]. Distribution [MS20b, MS24, CT22b, MOT21]. div [CH22b, CH25b]. div-conforming [CH25b]. divergence [HHZ23, LvBV25]. divergence-free [HHZ23]. divergence-preserving [LvBV25]. Divisibility [Sgo23]. domain [LNW25, PYAT21, RSM22]. domains [ACS21, BNWX23, BY24, KK20, Ric22, RS24]. dominated [JKK21]. dominating [GKS22, GHS23]. Dörfler [PP20]. Double [XB24]. Double-variable [XB24]. Doubly [ABG^{+}24]. DP/BDDC [WZSP21]. drift [Gos20]. drift-diffusion [Gos20]. Drinfeld [CL26]. driven [SSX26]. DtN [WZZ22]. DtN-type [WZZ22]. Dual [CT22c, JCWH21, JW25, LD20, Wan25, ZYYZ25]. duality [Bar21, KNT21]. Durand [RSS23]. Dyer [KMwaabAVS23]. dynamic [YLLO24]. Dynamical [KNZ25, BEKU21, Lav21, YEHS25]. Dynamics [FS24, BFS22, Chu25]. Dynkin [MM21].

ECM [BS22]. ECM-friendly [BS22]. Effective [Avd21, LPPV24]. Efficient [LX25, MX26, BK23b, MNST20, TJW25]. eigenproblems [Reu22]. eigenschemes [BGV24]. eigenvalue [ALPS25, BT22, CSX23, CLLZ25, Gal23, GSTZ22, GDST26, LX22, ZC21]. eigenvalues [ABD20, CDM^{+}20, HM22, PNR25, RS24]. eigenvector [BBGS22]. eigenvectors [CDM^{+}20, OR23]. Einstein [HY25a]. elastic [BR21, DLL21]. elasticity [AP24, CH25a, CH22a, DS20a, KBPS21, LS23, LS24, WZSP21]. elastodynamics [DS20a, MIJR21]. electrical [BZ24]. electromagnetic [Nic24]. electronic [KL23]. electrostatic [RS23]. Element [CH20, AP24, AFGS23, BL25a, BL25b, BCFM25, BFO20, COSS26, CNV21, Car20, CH25a, CH22a, CH24b, CH25b, FOS22, FP20, GT23, GHPS21, GN25, GM21, GSTZ22, GS20, HY25b, HLX22, JKK21, KK20, KMRB25, LvBV25, LL21, Li21, Li22a, LYZ24, LQXY24, LMS25, LYZ25, Tra25, Reu25, ORS24, Reu22, WLL21, ZMZW23]. element/holomorphic [GSTZ22]. elementary [HKM25]. elements [BHKS25, CDG22, CS22, CH22b, DS21, FH25, GSW23, HM22, JT22, KS20, LS23, LS24, MIJR21, ORZ21, ST23]. eliminating [PSW25]. ellipse [SH25]. elliptic [AMU21, BNP24, BS22, BKKS24, BHL24, CSX23, CM23, DGJ20, DES22, Fre23, GJN20, HP23, HKS21, KZ23, KNRR22, Tra25, WY22a]. elliptic-parabolic [AMU21]. Embedded [FKRS21]. empirical [COS21]. Energy [CEOR20, FTY24, LS23, CNV21, GWWZ24, lLW25, LWW^{+}21, WLF25]. Energy-preserving [CEOR20]. energy-stable [CNV21]. enriched [HM22]. ensemble [CT22a, LS21]. entropy [FJRR21, GM21, Zha22]. Enumeration [AMV22, BBERPC24, DMV23]. envelope [AMP22]. envelopes [LN22]. equation [AMV22, AFKL21, BBP24, BV23, BCY20, BFS22, BCF23, BW24, BRT23, BDF21, BCDF25, BY24, BFO20, CHEF24, CH25b, Coe24, CKMS26, CHS22, DE25, EN24, GT23, GGKO23, GS20, GMS21, HLZZ21, HTVZ23, HW22, HPR24, HJS24, HJQ23, IGMS23, JLOS25, KR25, Kir20, LQW21, LQT22, Li22b, LYZ24, LZ25, Reu25, MV24b, PYAT21, SW25, TCLS25, WZZ22, WZ22, WLSZ23, WY22b, WGMZ25]. Equations [BGV24, CH20, AGdS22, AHHT23, BGH^{+}23, BKRS20, BCHR20, BBK^{+}24, BCCET22, BF24, BKKS24, BRS24, CRS22, CNV21, CH24a, CFS21, CX24, DGM20, DES22, DL22, FS22, FMS24, FMZ23, GS22, Gos20, GLFN20, HHZ23, HLMT21, HS21b, HS23, JKK21, KNZ25, KBPS21, KHR23, KÇR23, LZMW20, LL20, LMY21, LSL22, LMS25, lLZ21, MLR22, ORS24, PT22, PSS25, QT20, RSM22, SWZ21a, SH21, Sog20, SJ24, SX21, Wal99, Wal21, YCQ20, YEHS25, Zha22, AK24]. Equiangular [GSY23]. equilateral [BA22]. equilibration [CFEV22, CF23]. equilibrium [GCO22, RS23]. equipowerful [BGPW21]. Equivalence [GKS22]. equivariants [HB22]. Eratosthenes [Hel20]. ergodic [LV20]. Error [BW24, Bar21, BLMM^{+}23, CH24a, CT20, DM24, Kop21, MS22, MT22, ORS22, SH25, AGY21, BM22, BFS22, BCF23, BK23b, CLY23, CNV21, CFEV22, CF23, CX24, CWS26, De 20, DGS22, DMM24, DL22, EGH25, FMS24, GK26, GS20, HR25, HLZZ21, JLOS25, KBPS21, Ley21, LSL22, MV24b, Nak20, NS23, PT21, SSX26, SX21, SW25, WY22a, WZZ22, WLF25, XKLW23, XLL24]. errors [ABD20, GGO20, Li21, LRY20]. Escott [CMSV24]. essentially [CDG22]. estimate [CWS26, SW25, WLF25]. estimates [BV23, BW24, Bar21, BLMM^{+}23, BHL24, CFEV22, CX24, CHL22, CHL24, HLZZ21, JLOS25, Ley21, Li25, MS22, MT22, MV24b, ORS22, SSX26, SX21, WY22a, XKLW23]. estimating [FRS25]. estimation [BK23b, BPT21, COS21, HKS21, KBPS21, THW25]. estimators [CF23]. Euclidean [DGS26, DK21, GSY23]. Euler [AL20, AF20, ERS21, FLMSS22, HLMT21, LM26, Zha22]. Eulerian [ORS24]. evaluation [BKM20, BLM24, CT22b]. eventual [WG20]. everywhere [Dem22]. evolutionary [JKK21]. evolving [ORS24]. Exact [FGN20, GLN22, Lap20]. exactly [KHR23, LS23]. exceptional [BJT20]. exhibiting [OR23]. Existence [BEKU21, Dem22]. Exotic [LV20]. expansion [CT20, CEFO20]. expansions [DMM24, HM22, KZ23, LPP21, XKLW23]. Expected [HRR25, HW21]. Explicit [AGdS22, BT20, BRS23, BK23b, CHL22, EHP22, GPS24, Kay22, Rüd22, vBHM20, AMU21, BCFM25, CIK^{+}25, Cre20, FTY24, HJLW20, HS21b, HS25b, LMY21, LQT22, LLZ23, lLW25, SSX26, WLSZ23, CHL24]. exponent [BST21, Har21, HH22, KK21, TY25]. exponential [DL22, LS25, lLW25]. exponential-type [LS25]. extended [CIK^{+}25, GI23, HLX22]. Extending [HR25]. Extensible [THW25]. extensions [CFEV22, EV20, FW23, FR22, Pag22]. exterior [AFGS23, HLX22]. extrapolation [GNT24]. extreme [XB24].

Facial [Lap20]. factorisation [Har21, HH22].

factorization

[COS25, DH20, Hit21, LX25, QMSZ25]. Faltings [Avd21, Dua21]. families [BS22, MSSS23, Pul20]. family [GP24]. fans [BKR20]. far [Ros20]. Farin [GLN22]. Fast [CKNS21, ERS21, WX22, LR21, LS20a, Bör25, YLLO24, BNPP20, CHW20]. Faster [Har24]. Faulhaber [GPW21]. fault [GO20]. FEM [HHZ23]. FEMs [MS22, SW21]. Fermat [BCDF25]. Feshbach [DSS23]. FETI [WZSP21]. FETI-DP [WZSP21]. FETI-DP/BDDC [WZSP21]. Few [GJL^{+}24, GMZ20]. Fiedler [De 20]. field [BCM21, BK24, GKL^{+}21, JT22, YLLO24]. fields [BNP24, BS24, DGJ20, DK23, JST22, KK21, LV22, LC22, Mar21, OW20, Pag22, Rüd22, WH24]. fifth [Har21, HH22]. film [GGKO23]. Filon [GI23, WGMZ25]. filter [LS21]. Filtered [JLOS25, GGO20]. filters [EGL24]. find [SW20]. finder [RSS23]. Finding [PSY20, Kho20, Wan25]. FindStat [ELM^{+}24]. Finite [CH22b, CH24b, FOS22, GN25, GSTZ22, Tra25, AP24, AFGS23, BL25a, BL25b, BLM25, BLMM^{+}23, BAL21, Ben22, BCFM25, BCHR20, BDF21, BHKS25, BBM21, BFO20, COSS26, CNV21, Car20, CH25a, CH22a, CC25, CH25b, Coe24, DMV23, EGH25, FP20, FH25, GT23, GHPS21, GS20, HLMT21, HY25b, HLX22, HJS24, HB22, JT22, JKK21, KS20, KK20, LMP20, LS23, LS24, LvBV25, LL21, Li21, Li22a, LYZ24, LQXY24, LMS25, LYZ25, Reu25, Mar21, MIJR21, ORS24, Reu22, Sog20, WLL21, XLL24, dSKJV23, GM21]. finite-size [XLL24]. finiteness [MT20, WD24]. first [CT22b, GKS22, JCWH21, WY22b]. first-order [GKS22, JCWH21, WY22b]. fixed [AF20, BNP24]. Flory [GWWZ24]. flow [ASZ23, BL25a, COS21, CLLZ25, FOS22, HY25b, lLW25, SW21]. flows [BBDM26, CDZ22, FTY24, GNT24, HLS21, MS20a]. fluid [COSS26, CZ25]. fluid-structure [CZ25]. fluids [FOS22]. flux [AV25, BRT23, CHEF24, CWS26, EV20, EGH25, MS20a]. flux-limited [MS20a]. fluxes [CX24, LZMW20, SX21]. Fock [XLL24]. Fokker [SW25, AHHT23, LL20]. forced [Ros20]. form [CIK^{+}25, DDDR20, Tra25]. forms [AC22, BMR24, BJT20, BV24, BK22, CT22c, DPRT24, LwaabAL21, MR23, MS20c, MV24a, PSY20]. formula [CIK^{+}25]. formulae [Not22]. formulas [FS22, HS21b]. formulation [BRT23, CZ25]. Four [JT22, Ros22]. fourfolds [AKPW25]. Fourier [BLMM^{+}23, CMR25, DSS23, Li25, QH22, WY22b]. fourth [EH20, FSTW25]. fourth-order [FSTW25]. fraction [BST21, DKS25, SiTiY20]. Fractional [KR25, AHHT23, BBP24, BM22, BKKS24, BNWX23, FMP21, HLZZ21, KR22, Kop21, MSS24]. fractional-order [Kop21]. fractions [MR24, WD24]. fracture [BDM24, HY25b]. framelets [DH20]. framework [BHK^{+}25a, BF24, BDH^{+}21, HKLP23, KNT21, Chu25]. Frankl [Pul20]. free [ABP25, CZ25, HHZ23, HRR25, KR25, Mar21, MIJR21, MOT21]. Freeness [BDS20, BK23a]. frequency [BY24, PV20]. frequency-time [BY24]. frictionless [BDM24]. Fried [FJRR21]. friendly [BS22]. Frobenius [FS20]. Fuchsian [Ric22]. Full [DL22, AFKL21]. Fully [CLLZ25, DD23, Han23, HJS24, HS23, WY22a]. Function [KMNN21, AF20, Avd21, BPT21, BKL^{+}21, BLM24, CT22b, CHS22, GSTZ22, HR25, LR21, SH22, WY23]. Functional [MSSS23]. functions [BDOT24, BKPU24, BMOR21, BKM20, BRS24, CS23, CC25, FRS25, GAHW21, GdLL25, GHS23, HSW22, KNS22, LV22, MX26, PNR25, SH25, Wan24]. fundamental [HB22, Ric22]. fusion [PS21].

Galerkin [KG21, AHPS20, Bar21, BKRS20, CDG22, CH24a, CX24, CJS23, FSTW25, GKMV25, HLZZ21, HW20, HLX22, KHR23, KÇR23, LZMW20, Sun25, MLR22, PSW25, SH21, SW21, SX21, TXS20, TCLS25, WLSZ23, YCQ20]. Galerkin-mixed [SW21]. Galois [Fre23, Mas20]. gamma [LR21]. gap [GP24]. gaps [ALPS25]. Garcke [BL25a]. Gauss [SH25]. Gaussian [COS25, FKRT24, GKS24, KOG21, LRY20, Not22, WW24]. GCD [HG26]. Gearhart [Rie23]. GenCol [FP25]. General [LRY20, BVA20, CL26, Han23, HJQ23, KBL21, LDHS23, MR23, QT20]. Generalized [BM25, Cre20, DH20, Rie23, WH24, BGO24, BCDF25, CS23, CX24, GPW21, Hua20, LZMW20, MS22, SX21, XB24, EHP22]. generally [RSS23]. generated [OW20]. generating [GM25b]. generation [GSW23, LÖS20]. generator [MS24]. generators [BAFG20, DH20, JST22]. generic [BS23]. genetics [CT22b]. genus [ABG^{+}24, FS24, HSS21, Mil20, Rak24]. genus-2 [ABG^{+}24]. geodesic [GNT24]. geometric [AL20, LS20a, XB24, Zim21]. geometries [VSB26]. Geometry [AGY21, AAP22, CDLT25]. Georges [BGM22]. Gilbert [AFKL21]. Ginibre [CGT25]. GIT [BKR20]. GIT-fans [BKR20]. given [PSY20]. gives [GKS24]. global [Sun21]. globally [LD20]. Gluing [HSS21]. GMRES [CGT25]. Gonality [NO24]. good [Dem22, JLOS25]. Gordon [BFS22, WZ22]. graded [Kop21]. Gradgrad [HL21]. Gradgrad-complexes [HL21]. gradient [CS25, CLLZ25, FTY24, GNT24, JZCH24, lLW25, LM22, MS20a, YLLO24]. grandchildren [BA24]. graph [BGM22]. Graphs [GMZ20, BKKS24, GJL^{+}24, GEL24, LS20b, MSSS23]. great [BA24]. great-grandchildren [BA24]. greedy [BEL25]. Greenberg [Pag22]. grid [Not20, THW25]. grids [AHPS20, BLM25, Gos20, HY25b]. Gross [CLLZ25, CIK^{+}25, CKMS26, HW22]. ground [HY25a]. group [AMP22, BGO24, BBERPC24, Box21, BMP21, GM25b, KK21]. groups [BGO24, BJT20, BKK24, CC25, GJN20, GSW23, HB22, Hul22, KS20, KLM20, LMP20, MS25b, PS21, Ric22]. Growth [GJN20, CLY23]. Guaranteed [CDM^{+}20].

Hadamard [LÖS20]. half [Coe24]. half-line [Coe24]. Halpern [CL25, HXDM24]. Halpern-type [CL25]. Hamilton [Sog20]. Hamiltonian [BRK25, CDZ22, GMZ20, GJL^{+}24, JO23]. Hamiltonians [Chu25]. harmonic [ABP25, Li25]. Hartree [XLL24]. Hasse [BVA20, HKY22]. HDG [BRT23, CZ25, DS20a, DS21]. heat [BDF21]. Hecke [PT22]. heights [GM25a, vBHM20]. Helmholtz [BRT23, BCFM25, GGS23, GS20, LX22, WGMZ25]. Hensel [BV24, Mas20]. Hensel-lifting [Mas20]. heterogeneous [PV20]. heuristics [DK23]. hierarchical [COS21, SJ24]. High [HJLW20, LYZ24, AP24, AHPS20, DE25, EMS26, HHZ23, Hua20, HPW20, PV20, TXS20]. high-contrast [PV20]. high-dimensional [HPW20]. High-order [LYZ24, AHPS20]. Higher [AFKL21, FJRR21, GdLL25, HR25, NS23]. higher-level [GdLL25]. Higher-order [AFKL21, NS23]. higher-rank [FJRR21]. highly [DLL21, GI23, WX22, WW24]. Hilbert [KOG21, LM22]. Hilliard [GWWZ24, BV23, HJS24, LQW21, LQT22, Li22b]. Hodge [Bla22]. holomorphic [GSTZ22]. homogeneous [ABD20, BJT20, SLN23]. homology [DEG22, LPPV24]. Homomesies [ELM^{+}24]. homotopies [BSW23]. homotopy [LdCS^{+}21]. Hood [Reu25, ORZ21]. Hopf [FP20, LX25]. house [Fla21]. Huber [Avd21].

Huggins [GWWZ24]. hybrid [BBDM26, BY24]. Hybridization [AFGS23]. hybridized [KHR23, KÇR23]. hydrostatic [BL21]. hyperbolic [BAL21, Ben22, DM24, DGSW22, GK26, GM21, HJLW20, HS21b, HS25b, HLZ23, MZ21, PSW25]. hyperelliptic [DIS23, GM25a, Kay22]. hypersurfaces [ES24, LPPV24]. Hypocoercivity [BCHR20, BF24]. Hypothesis [EHP22].

Ideal [CMSV24, BK24]. idealized [YEHS25].

ideals [Bla22]. idempotents [BGO24].

Identifying [EN24, OR23]. II [BCDF25, CWS26]. iid [KNS22]. images [BLH22]. Imaginary [KK21, JST22]. imaging [KR22]. IMEX [HS21b, HS25b, HHZ23]. IMEX-BDF [HS21b]. IMEX-RK [HS25b]. IMEX-SAV [HHZ23]. impedance [BZ24, GGS23]. implementation [FKRS21]. Implicit [LLZ23, AFKL21, FTY24, HS21b, HS25b, HS23, LQW21, LQT22, SSX26, WLSZ23]. Implicit-explicit [LLZ23, HS21b, HS25b, LQT22, WLSZ23]. Implicitization [DS20b]. implicitly [FOS22]. implicitly-constituted [FOS22]. implies [Fei22]. implying [Pul20]. Improved [BCF23, Ric22, BKRS20, GPW21, Hel20]. inchworm [CLY23]. incomplete [KZ23]. incompressible [BGH^{+}23, BBDM26, KÇR23, MLR22, SW21, WZSP21]. indefinite [ZWZ20]. inequalities [EG24, MT22, WH24]. inequality [BM25, GAHW21, GPW21, LDHS23]. Inexact [BLM24, BKM20, Wan25]. inextensible [BR21]. Inf [Fei22, ORZ21, CS22, ST23]. Inf-sup [Fei22, ORZ21, CS22, ST23]. infinite [Box21]. infinitely [HJ24, Ula24]. Influence [BEL25]. inhomogeneous [CEHT23]. inner [Sun25]. inner-stage [Sun25]. inpainting [QMSZ25]. integer [ALPS25, Har21, HH22, Har24, Hit21]. integers [DDDR20, Fla21, WWW21]. integral [ASZ23, DLL21, FMP21, GM24, HLZZ21, KZ23, LPP21, RSM22]. integrals [BCM^{+}24, GI23, WX22, WW24, WGMZ25]. Integration [KOG21, BT20, BCCET22, CH24a, Kay22, KMNN21, LS20b, NS23]. integrator [WZ22, WY22b]. integrators [BCJ23, CRS22, DL22, FMS24, JO23, LV20, LS25, SWZ21a]. interaction [COSS26, CZ25]. interfaces [BDM24]. interfacial [ASZ23]. interior [BY24, Ley21, LD20, ZMZW23]. interior-point [LD20]. interpolant [FP20]. Interpolation [DST23, HR25, MR20]. intersection [IL23, vBHM20]. interval [Ben22, CHL22, CHL24]. intimate [GK26]. intrinsically [BKM20]. invariable [GSW23]. invariance [BKS25]. invariant [MS20c]. invariantizable [LL24]. invariants [BHK^{+}25b, BDS20, DK23, HW22, HB22]. inverse [GLFN20, HKS21, KR22, LM22, SJ24, ZCL^{+}22]. inversion [CT22a, CT22b]. inversions [KS20]. inversive [MS20b]. irrationals [BCM21]. Isogenies [SF25, BNP24, Mil20]. Isogenous [DIS23, ABG^{+}24]. isogeny [KNRR22]. isogeometric [MNST20, VSB26, WZSP21]. isolated [BHK^{+}25b]. isolation [HJLZ20]. isometries [ABP25]. isomorphisms [HJ20]. isoparametric [LQXY24]. Isotropic [LC22, MIJR21, SH21]. Iteration [BKM20, DGM20, GGO20, HXDM24, Zen23, ZYYZ25]. iterations [CL25, ZZY23]. Iterative [GHM24, GCL23, HW20]. Iwasawa [DK23].

Jacobi [LX22, MT20, SLN23, Sog20, SH25, XKLW23]. Jacobi-type [SLN23]. Jacobians [Cre20, DIS23, Mas20, Mil20, OS21]. jittered [Doe22]. jump [LL20].

K3 [BVA20]. Kac [FS22]. Kaczmarz [Rie23, Ste21]. Kahan [BCM^{+}24]. Kalman [CT22a, LS21]. KdV [LZMW20]. Keating [CIK^{+}25]. Keller [WLF25]. Kelmans [BGM22]. kernel [COS21, KOG21]. kernels [AK24, GCO22, KOG21]. Kerner [RSS23]. Kerr [MV22]. Kerr-type [MV22]. Khovanskii [BSW23]. kind [CT22b, KZ23]. kinetic [BCHR20, BF24, BRK25, BL21, HS21b, YEHS25]. Kirchhoff [DD23, FH25, KMRB25]. Kissing [DK21]. kit [MV24a]. Klein [BFS22, BJT20, WZ22]. Kleinian [BKK24]. knots [BR21]. Korn [WH24]. Koshy [Rie23]. Krylov [CS23, PSS25]. Kummer [SH22, SF25]. Kutta [BRK25, HS25b, FTY24, LLZ23, lLW25, Sun25, WLSZ23].

L2 [Kop21]. L2-type [Kop21]. Lagrange [DMM20]. Lagrangian [FM20, HMXY23, LDHS23, YCQ20]. Laguerre [Wan24]. Landau [AFKL21]. Langevin [BRK25, BP24, CHS22]. Langevin-simulated [BP24]. Laplace [GLFN20, Reu22]. Laplace’s [KR25]. Laplacian [FMP21, HLZZ21, RS24]. Laplacians [BNWX23]. Large [LL20, MS25b, Li22b]. Large-scale [MS25b]. lattice [BKS25, CKNS21, DGS22, God26, NS23]. lattices [HJ20, KMNN21]. Laurent [PNR25]. law [GCO22]. laws [AV25, BAL21, DM24, HJLW20, PSW25]. Lawson [ZYYZ25]. Lax [Chu25]. layer [CWS26, DE25]. layer-upwind [CWS26]. LDG [CWS26]. leading [CF23]. leapfrog [GMS21]. Learning [FKRT24, FMZ23]. Least [BNPP20, Car20, HNP22, HX24, MS25a, MNST20]. Least-Squares [BNPP20, Car20, HNP22]. left [BBERPC24]. Legendre [CX25]. Lehman [Hit21]. Lemma [CD23]. lemniscates [LP25]. less [Fla21]. letters [Ros20]. level [GdLL25, LX22, PSY20]. Levenshtein [BDH^{+}21]. Levin [WX22]. Lévy [SSX26, SW25]. liar [FS20]. Lie [BJT20, BCJ23, CEFO20, CSS24, JLOS25]. Lifshitz [AFKL21]. lifted [BDH^{+}21]. lifting [BV24, Mas20]. like [ZZY23]. limit [BCHR20, HLS21, LS21]. limited [LYZ24, MS20a]. Line [dSKJV23, BBK^{+}24, Coe24]. Line-search [dSKJV23]. Linear [CL25, AP24, AMP22, BLM25, BCHR20, BBM21, BGMS23, CH24a, CT22c, DGSW22, FSTW25, HJQ23, HLZ23, KBPS21, LS23, LQW21, LM22, Tra25, MIJR21, MV24a, PP20, SSX26, Ste21, WLSZ23, YCQ20, dCIDV24]. linearization [HW20]. linearizations [De 20]. linearized [LZMW20, ZZY23]. linearly [AFKL21, ZWZ20]. lines [GSY23, vLW23]. Lipschitz [BNWX23, LC22, ZZY23]. Lipschitz-like [ZZY23]. lists [Ros22]. Littlewood [BGPW21, HJ21, LdCS^{+}21]. Local [Sun21, CH24a, CJS23, FSTW25, GMS21, HJLW20, LZMW20, TXS20, TCLS25, WLSZ23, XKLW23]. localization [HP23, OR23, PNR25]. Locally [LL24, CSX23, Li22a, MS22]. locus [BK23a]. log [HH22]. log-log [HH22]. logarithm [GKL^{+}21]. logarithmic [KZ23]. logarithms [MV24a]. lognormal [CM23]. Long [FMS24, BFS22, BCF23, CRS22, LV20]. Long-time [FMS24, BFS22, BCF23]. loss [WY22b]. losses [TFD25]. Love [DD23, FH25]. Low [BHL24, GHS23, HLS21, JLOS25, LS25, QMSZ25, SWZ21a, BEKU21, FMS24, KNZ25, ORS22, WZ22, YEHS25]. Low-rank [GHS23, QMSZ25, BEKU21, KNZ25, YEHS25]. Low-regularity [LS25, SWZ21a, FMS24, WZ22]. lower [Gal23, OSS25]. LSOS [dSKJV23]. lumping [VSB26]. Lyapunov [BST21, GAHW21].

MAC [BGH^{+}23]. Mach [HLS21]. Magnus [CEFO20]. Mahler [FP22, LP25, TG25]. majorized [ZWZ20]. Mandelbrot [LwaabAL21]. manifold [HLR23]. manifolds [CEOR20, MZ21, SLN23]. Manin [AFPwaabAVS22]. many [HJ24, OS21, Ula24]. map [AE25, HS25a]. maps [AAP22, ABP25, BGMS23, ELM^{+}24]. Marcinkiewicz [EG24]. marginal [ACEL21, FP25]. marking [Car20, PP20]. Markov [DMM20, LX25]. Markov-modulated [LX25]. Mass [VSB26, KHR23, WLF25]. mass-conserving [WLF25]. massively [AHPS20]. materials [CHV24]. matrices [ALPS25, AFPwaabAVS22, CGT25, FRS25, GP24, LÖS20, Bör25, Li25]. Matrix [PSS25, ABD20, BDM24, CS23, DK23, DH20, Hua20, MM21, MS24, OW20, PNR25, XB24, Zim21]. matrix-fracture [BDM24]. matrix-valued [OW20]. maximal [BA22, DIS23, KK20]. maximization [XB24]. Maximum [Li22a, HJQ23, LL21, LQXY24]. Maximum-norm [Li22a]. Maxwell [LX22]. mean [EGL24, YLLO24]. means [PO23, PO24]. measure [BHK^{+}25a, LP25]. measure-transport [BHK^{+}25a]. measures [BCM^{+}24, Fla21, GCO22, TG25]. measuring [HR25]. media [BCFM25, MV22, PV20, RSM22, SW21]. median [EGL24, PO23, PO24]. median-of-means [PO24]. mediants [DKS25]. membrane [COSS26]. meromorphic [BNPP20]. mesh [Li22a]. meshes [DD23, Han23, HLR23, Kop21, LYZ25, SX21]. metabelian [KLM20]. Method [CH20, AGdS22, ASZ23, ACS21, AHHT23, BL25a, BBP24, BL25b, BLM25, BLMM^{+}23, BCM^{+}24, BFO20, CLY23, COSS26, CHV24, CFS21, CS25, CLT22, CT20, CHW20, CSX23, CJS23, CWS26, CZ25, CEHT23, DES22, DLL21, Dua21, DSS23, FMP21, FP20, FSTW25, GI23, GS22, GSTZ22, GMS21, GG24, HLZZ21, HMXY23, Hit21, HY25b, HJS24, Hua20, IGMS23, JZCH24, JW25, JKK21, KHR23, KÇR23, Kop21, Li22a, LQT22, LX22, LD20, LDHS23, Sun25, Reu25, MIJR21, MNST20, ORS24, OW20, PSW25, Rie23, SH21, SSX26, Ste21, TXS20, TCLS25, WLL21, WLF25, XB24, YLLO24, ZMZW23]. methods [AHPS20, BCY20, BFS22, BCF23, BW24, Bar21, BT22, BCCET22, BRK25, CDG22, CH24a, Car20, CS23, CEOR20, CYB^{+}21, CX24, CM23, FRS25, FTY24, Gal23, GHPS21, GM21, GS23, GKMV25, GO20, HW20, HLMT21, HP21, HJLW20, HLLM25, Kho20, KMRB25, Kir20, KG21, KPW21, LvBV25, LL21, LZMW20, Li21, LSL22, LN22, Li22b, LLZ23, LYZ24, LQXY24, LMS25, LYZ25, lLW25, Wan25, MSS24, Not20, PSS25, QESS24, Reu22, SX21, WX22, WLSZ23, YCQ20, dSKJV23]. metric [BKKS24, HLR23, TJW25]. metrical [DKS25]. million [Hul22]. Minimal [MSS24, CW23, FJRR21, PP20]. Minimality [AC22, BGM22]. minimax [ZYYZ25]. Minimization [BRS24, ES24, GAHW21, BK23b, BKM20, BLM23, BLM24]. minimum [DH20]. miscible [SW21]. Mixed [FH25, Gal23, Bar20, BDM24, BHL24, CH25a, CZ25, GKS22, GHS23, HY25b, LS23, LS24, LvBV25, Li21, SW21, WZSP21]. mixed-dimensional [BDM24]. mobility [HJQ23]. Möbius [ABD20]. mod [Fre23]. model [BDM24, GCL23, HPR24, HS21a, KZ22, WLF25, YEHS25]. Modelling [FMZ23]. models [FKRT24, GHM24, HY25b]. modifications [Zim21]. Modified [CHV24]. Modified-operator [CHV24]. Modular [HJ24, MS20c, AKMJ^{+}24, BS22, BKS23, BMR24, Box21, CCG20, CT22c, LwaabAL21, MR23, NV23, NO24, Rak24]. modulated [LX25]. modulation [EG24]. module [WG20]. modules [CL26]. moduli [BK23a]. modulo [MS20b, MS24]. Møller [XLL24]. moment [Kir20]. moments [ACEL21]. Monge [GT23, QT20]. monic [BDOV20]. Monochromatic [GKW24]. monomial [BGO24]. monotone [AK24, BBM21, CHEF24]. Monte [BRK25, CLY23, HW21, HKS21, KPW21, MS25b, SSX26]. Montgomery [CMR25]. Mordell [Box21]. Mori [BKR20]. morphisms [BLH22]. Motion [QESS24, EGL24, LX25]. moving [GEL24]. multi [CX24, CEHT23, ERS21, HLLM25, WGMZ25]. multi-agent [CEHT23]. multi-block [HLLM25]. multi-dimensional [CX24, WGMZ25]. multi-precision [ERS21]. Multidimensional [SiTiY20, BST21, GPS24, PO24]. Multilevel [Füh22]. Multiple [BY24, HJLZ20]. Multiple-scattering [BY24]. multiplication [DIS23, Har24, Bör25]. Multiplicative [LwaabAL21, BP24, HS23]. Multiplicities [CDM^{+}20]. Multiplicity [GdLL25]. multipliers [LDHS23]. multiprojective [HLRS21]. multiquadratic [KK21]. multirate [AGdS22]. multiresolution [KBL21]. Multiscale [Li25, MV22, HP23]. multivariate [CKNS21]. Müntz [CX25]. must [Ros20].

Nagao’s [KMwaabAVS23]. natural [FP20, Li21]. Navier [BLMM^{+}23, BGH^{+}23, GS22, HHZ23, KHR23, KÇR23, LSL22, MLR22, ORS24]. Near [CM23, AGY21, KZ23]. near-colliding [AGY21]. Near-optimal [CM23]. nearness [TJW25]. negative [DST23, Füh22, MSS24, SvV20]. neither [Sun25]. NEPv [LL24]. Nernst [LWW^{+}21]. Néron [vBHM20]. nested [DGM20, JW25]. nets [PO23, PO24]. network [DM24, GHM24]. networks [DE25]. Neumann [Li22a]. neural [DM24, DE25, GDST26]. neutrino [YEHS25]. Newman [HJ21]. Newton [CSX23, ZC21]. Newton-type [ZC21]. Newtonian [FOS22]. Newton’s [Zen23]. Nicolson [BBV21]. Nitsche [GS22]. NLS [ORS22]. no [Wal99, Wal21]. nodal [AAP22]. node [BLM25]. node-conservative [BLM25]. nodes [AGY21]. noise [BP24, GGKO23, HS23]. Non [SJ24, AHPS20, CYB^{+}21, DIS23, DK21, FOS22, JO23, LC22, Li25, MX26, MS20c]. non-archimedean [MX26]. Non-centered [SJ24]. non-conforming [CYB^{+}21]. non-Euclidean [DK21]. non-harmonic [Li25]. non-hyperelliptic [DIS23]. non-Lipschitz [LC22]. non-Newtonian [FOS22]. non-overlapping [AHPS20]. non-separable [JO23]. non-split [MS20c]. Nonconforming [CH20]. noncongruence [BMR24]. nonconvex [JZCH24, JW25, Li22a, dSKJV23]. nondivergence [Tra25]. nonergodic [LZ25]. Nonexistence [BJT20]. nonfree [BK23a]. nonisolated [Zen23]. nonisothermal [COSS26]. nonlinear [BFS22, BCF23, BW24, BAL21, CT22a, COS25, FMS24, GKMV25, HW20, HX24, HPW20, JKK21, KR22, LMY21, LD20, MV22, Nic24, PNR25, SWZ21a, WZ22, WY22b, ZCL^{+}22]. nonlinearity [BFS22, BW24]. Nonlocal [AV25, LQW21, WZZ22]. nonmatching [HY25b]. nonnormal [Not20]. nonrelativistic [BCY20]. nonsmooth [Bar21, JZCH24]. nonsurjective [Fre23]. nonsymmetric [KZ22]. nontrivial [GdLL25]. nor [Sun25]. Norm [HKY22, CWS26, FW20, HHY20, LS23, Li21, Li22a, TJW25]. Normal [CH25a, BNPP20, DES22]. Normal-normal [CH25a]. norms [CL26, FR22, Füh22, HR25, MSS24]. note [DS21, QT20]. novel [LDHS23, QMSZ25]. Novikov [TCLS25]. null [BGPW21]. number [BS24, DK23, DH20, DK21, EPR21, EHP22, FS22, HHZ23, HJ21, HLS21, JST22, KK21, LR21, Li25, MS24, PT21, Rüd22]. number-theoretic [LR21]. numbers [BCM21, BK22, CT22b, CW23, CHL22, GPS24, MS20b, MOT21, Rüd22, Sgo23, TG25, WG20, XXX25, CHL24]. Numerical [BBP24, BDM24, BBL20, BSW23, CLY23, CYB^{+}21, CKMS26, DE25, GLFN20, Kir20, LdCS^{+}21, LS20b, MIJR21, Nic24, AV25, AHHT23, BCCET22, BKKS24, BBDM26, BAFG20, BA24, CFS21, CHEF24, CLT22, CT22b, CX24, CHS22, GK26, GWWZ24, HLRS21, LZMW20, LQW21, LMY21, LWW^{+}21, MS20a]. Nürnberg [BL25a]. Nyström [BRK25].

objective [BLM24]. oblique [BBV21]. obstacle [DLL21]. obstructions [BVA20]. octagon [BA22].

Odd [BVA20, FM20]. ODEs [GDST26]. OEDG [PSW25]. one [GJL^{+}24, Har21, HH22, HHY20, HKY22, Sun25, PO23, YCQ20, Zim21, vLW23]. one-fifth [Har21, HH22]. onto [NS23]. operator [CHV24, DST23, GSTZ22]. operators [BNPP20, BBM21, Col20, DSS23, Sun25]. Optimal [BZ24, CX24, CWS26, EGH25, LMS25, SX21, WLF25, YCQ20, ACEL21, Car20, CSX23, CM23, DGS22, FMP21, FP25, GKS24, HNP22, KZ22, Lav21, Li21, MS22, TY25, WY22a, XXX25]. optimality [GHPS21]. optimization [CMR25, CHW20, DGS26, GDST26, HLLM25, JZCH24, JW25, LDHS23, LZ25, QH22, SLN23, dSKJV23]. Order [CH20, FMZ23, AP24, AFKL21, AHPS20, BKRS20, BVA20, BBM21, CH24a, CFS21, EH20, FSTW25, Füh22, GKS22, GSW23, HR25, Han23, HPR24, HJLW20, HJQ23, Hul22, JCWH21, Kop21, LYZ24, LNW25, Tra25, Sun25, Wan25, NS23, PS21, Sgo23, SvV20, TXS20, WZZ22, WY22b, XLL24, dSKJV23]. Orderly [LÖS20]. ordinate [SH21]. Orientable [MZ21]. Orthogonal [OX20, NS23, OW20, RS23, Zim21]. orthogonality [Fei22]. Oscillation [PSW25]. Oscillation-eliminating [PSW25]. oscillatory [GI23, WX22, WW24, WGMZ25]. outlier [VSB26]. overall [GHPS21]. overlapping [AHPS20]. overpartition [WY23].

packing [CT22c, dCIDV24]. packings [BKK24]. Padeé [BNPP20]. pair [CMR25, XB24, QH22]. pairs [Hua20, TY25, WDM20]. parabolic [AMU21, BEKU21, BBV21, CNV21, DSW22, KK20, Kir20, Kop21, MNST20, WY22a]. parallel [AHPS20, FKRS21, HMXY23]. parameter [BEL25, COS21, FMZ23, KBPS21]. parameter-dependent [BEL25, KBPS21]. parameters [HXDM24]. parametric [BL25a, BL25b, DM24, HPR24, SJ24]. parametrized [FS22]. paramodular [DPRT24, PSY20]. Part [YEHS25]. Partial [CH20, DE25, FMZ23, SJ24]. particle [CS25, FKRT24, HPR24]. particle-based [HPR24]. partition [WY23]. partitions [BGPW21]. patchwise [CFEV22]. patterns [SW20]. PDE [Tra25]. PDEs [BBV21, COS25, CM23, EH20, HPW20, Kir20, LLZ23, ORZ21, SSX26, TXS20]. Pell [Wal21, Wal99]. penalty [Ley21, ZMZW23]. perfect [Hul22]. Periodic [BCM21, Col20, NS23, XLL24]. periodicity [MT20, WD24]. periods [LPPV24]. permutation [CC25]. permutations [ELM^{+}24, WpL20]. Perron [MT20]. perturbation [XLL24]. perturbed [CJS23, HKLP23]. Pezzo [vLW23]. phase [XXX25]. phaseless [Sun21]. Picard [AFPwaabAVS22]. piecewise [BCFM25, FGN20, IGMS23, WH24]. piecewise-smooth [IGMS23]. Pitaevskii [CLLZ25, CKMS26, HW22]. PL [CYB^{+}21]. planar [AP24, HLR23, HS25a]. Planck [AHHT23, LL20, LWW^{+}21, SW25]. plane [BDS20]. planes [Pul20]. planning [YLLO24]. plate [FH25]. plates [DD23]. Plesset [XLL24]. POD [CKNS21]. Pohst [BM25]. point [EH20, EGH25, HKLP23, JCWH21, LD20]. points [AKMJ^{+}24, BKPU24, Box21, FS24, HJ24, Wan25, Mas20, NV23, Rak24, TG25, Ula24]. Pointwise [Ley21, XKLW23]. Poisson [ACS21, BS23, BCJ23, DE25, HP21, HPR24, LvBV25, LWW^{+}21]. polarized [Mar21]. polygonal [DD23]. polygons [Li22a]. polyhedra [LL21, LQXY24]. polyhedral [DS21, Han23]. Polynomial [BDOT24, LYZ25, BT22, CFEV22, CF23, EPR21, EV20, FW23, FS22, FGN20, HJLZ20, HG26, KRRZ20, Lap20, Sun25, PSS25, PO23, PO24, SLN23, WDM20]. polynomial-degree-robust [CF23, EV20]. polynomial-time [KRRZ20]. polynomials [ABD20, AFPwaabAVS22, BGPW21, CL26, CX25, FS24, HJ21, KNT21, Kho20, LwaabAL21, MSSS23, OX20, RS23, SWZ21b]. Polytope [BDM25]. polytopes [HLLM25]. polytopic [AHPS20]. population [CT22b]. poroelasticity [KMRB25]. poromechanical [BDM24]. porous [SW21]. posedness [AV25]. positive [Fla21, MM21, WWW21]. Positivity [WY23, GGKO23, GWWZ24, LWW^{+}21, WLF25]. positivity-preserving [GGKO23, GWWZ24, LWW^{+}21, WLF25]. Post [BKRS20, CEFO20]. post-Lie [CEFO20]. Post-processed [BKRS20]. posteriori [BV23, BM22, CDM^{+}20, CNV21, CFEV22, CF23, GK26, KBPS21, MV24b, WY22a]. postprocessing [AFGS23]. potential [EN24, EV20, GWWZ24]. potentials [MIJR21]. Power [BS24, GSW23, GCO22, KNRR22, KRRZ20]. powers [MS24]. practical [HH20, HKLP23]. precision [ERS21]. preconditioned [CSX23, LX22]. preconditioner [AHPS20, BNWX23]. preconditioners [AP24, SvV20, WZSP21]. preconditioning [FOS22]. Preparing [Chu25]. prescribed [FR22, HJ21]. presence [MKL21]. preserved [BCM^{+}24]. preserving [AHHT23, BL25b, BLM25, CEOR20, FW23, GGKO23, GWWZ24, HP21, HJQ23, LvBV25, LYZ25, LWW^{+}21, LNW25, SW25, WLF25]. pressure [BBDM26, LSL22, MLR22]. pressure-robustness [BBDM26]. primal [JCWH21, JW25, LD20, Wan25]. primal-dual [JCWH21, LD20, Wan25]. primality [BFW21]. prime [CHL22, CHL24, GSW23, KRRZ20, MS24, PT21]. Primes [CJdL21, FS21, AF20, BRS23, EHP22, LM26, LMP20, SW20]. primitive [GdLL25, JT22, MS25a, BGO24]. principle [BVA20, HKY22, HJQ23, LL21, LQXY24]. principle-preserving [HJQ23]. priors [HKS21]. probabilistic [SWZ21b]. probing [FRS25]. problem [ACS21, BBGS22, BBL20, BMP21, BHL24, CLY23, CFEV22, CF23, CLLZ25, CJS23, CMSV24, DD23, DLL21, Fla20, GH22, GSTZ22, GG24, GDST26, HHY20, HX24, KR22, KR25, KLM20, Kop21, LvBV25, Li22a, LX22, OSS25, PP20, RS23, SY23, TJW25, WGMZ25, ZC21, ZYYZ25, ZMZW23]. problems [ACEL21, AMU21, BEKU21, BM22, Bar21, BK23b, Ben22, BCFM25, BNPP20, CHW20, CSX23, CWS26, DSW22, FSTW25, GK26, GGS23, GKMV25, HP23, HW20, HKS21, HKLP23, HLLM25, HLZ23, JCWH21, JKK21, KR25, KK20, Ley21, lLW25, LM22, MNST20, SvV20, SJ24, WY22a, ZCL^{+}22]. procedure [CF23]. process [SSX26]. processed [BKRS20]. processes [COS25, FKRT24, LL20]. product [BHKS25, DS20b, PSY20]. products [AF20, BRS23, ERS21]. program [BCDF25]. programming [CJdL21, CT22c, HMXY23, ZWZ20, dCIDV24]. programs [LD20]. progression [AF20, LM26]. progressions [CJdL21, EHP22, GKW24]. projection [ABP25, DOS20, MR20]. projection-based [MR20]. projection-free [ABP25]. projections [DS21, NS23]. projective [DS20b, LPPV24]. Prolog [CSS24]. Prony [AGY21]. proof [FP25]. propagation [LRY20]. Properties [TFD25, CX25, MKL21]. property [LYZ24]. Prouhet [CMSV24]. proximal [JZCH24, YLLO24, ZWZ20]. Pseudoergodic [Col20]. pseudoprimes [FS20]. pseudorandom [MS20b, MS24]. pseudospectra [CGT25]. PSW [BFW21]. punctured [BBK^{+}24]. Pythagorean [KR20].

quadrant [LRR23]. Quadratic

[ABP25, Box21, MM21, NV23, AKMJ^{+}24, BNP24, BCM21, BV24, BK22, CIK^{+}25, FS24, GG24, JST22, MSSS23, OX20, Pag22]. quadratically [Zen23]. Quadrature [EMS26, BKPU24, EG24, Not22, RSM22, SH25, WW24]. quantile [HW21, THW25]. quantum [Chu25]. quartets [XB24]. quartic [HJ24]. Quasi [FMP21, HKS21, Li21, DH20, Fei22, HW21, IGMS23, KZ22, KPW21, MM21]. quasi-Cartan [MM21]. Quasi-Monte [HKS21, HW21, KPW21]. Quasi-optimal [FMP21, Li21, KZ22]. quasi-orthogonality [Fei22]. quasi-tight [DH20]. quasi-Trefftz [IGMS23]. quasilinear [CHEF24]. quaternion [QMSZ25]. Quinary [DPRT24]. quintic [Ula24]. quotient [BKS23].

radial [HR25]. radiative [DGM20, SH21]. Ramanujan [GM24, LR21]. random [BEL25, DK23, GG24, HRR25, KS20, LC22, WpL20]. Randomized [BRK25, KRRZ20, DGS22, LS20a, God26, PO23, PO24, Ste21]. Randomizing [GKS24]. range [LRR23]. rank [BEKU21, BDOV20, DGS22, FJRR21, GHS23, HMXY23, KNZ25, KMNN21, QMSZ25, YEHS25, Zim21]. rank- [KMNN21]. rank-1 [DGS22]. rank-one [Zim21]. rank-two [HMXY23]. Rate [GHPS21, AL20, DGS22, FMP21, GKS24, LZ25]. rates [BEL25, Car20, CL25, lLW25, SWZ21b]. ratio [SH22]. Rational [AL20, BGO24, BLH22, BDS20, CS23, DGJ20, FS24, KR20, Rak24, Ula24, WG20, ZYYZ25]. Rationality [HHY20]. Raviart [ST23, CS22, HM22]. reaction [CWS26, HLZZ21]. reactive [Zha22]. real [EPR21, LV22, Pag22, Sun21]. realizations [GEL24]. recognize [FP22]. reconstruct [HS25a]. reconstruction [BKPU24, BL21, DES22, EV20, KMNN21, WY22a]. Recovery [KNS22, LYZ25]. rectangular [PYAT21]. recursive [DMM24, MS24]. reduced [HP21, HY25b, FMZ23]. Reducing [Sun25]. reduction [Dem22, HPR24, JZCH24, Lap20, ZCL^{+}22]. reductions [Sgo23]. Refined [BBK^{+}24, EMS26, Li22a]. reflection [BKK24]. regime [BCY20, CRS22]. regular [DKS25, DMM24, FP22]. regularisation [KR25]. Regularity [BKKS24, BHL24, CL25, FMS24, Hua20, JLOS25, KK20, LS25, ORS22, SWZ21a, WZ22, WY22b]. regularization [EN24, LC22]. regularized [GT23, Wan25]. regulators [BM25]. relations [Bar21]. relative [DGS26, Hua20]. relaxation [HS21b, HS25b, HLZ23, LLZ23, LD20, YEHS25]. relaxed [HMXY23]. removal [VSB26]. repetitive [DOP23]. representation [HG26]. representations [BCM21, BK22, Dua21, Fre23, LC22, Mas20]. reproducing [KOG21]. resectioning [CDLT25]. resemblance [CC25]. residual [MSS24]. resolution [BCY20]. resolutions [DS20b]. respect [GHPS21]. restoration [BLM24]. restricted [GGS23, KNS22]. results [CS25, DKS25, PNR25, Sao22]. retrieval [XXX25]. reversion [GH22]. Revisiting [ZZY23]. revolution [OX20]. Reynolds [HHZ23]. Rham [CH24b]. Rice [FS22]. Richards [MV24b]. Richardson [LdCS^{+}21]. Riemann [EHP22, BPT21, LS20a]. Riemannian [CEOR20, HLR23]. right [BAFG20]. right-generators [BAFG20]. rigid [GEL24]. rings [GLNY20, KRRZ20]. Ritz [Nak20]. RK [HS25b]. RMSE [GKS24]. Robin [Sao22]. Robust [BV23, BNWX23, KBPS21, CFEV22, CF23, EV20, KZ22]. robustness [BBDM26]. Roch [LS20a]. root [Kho20, KRRZ20, MS25a, RSS23]. roots [PNR25]. rotating [HY25a]. rough [AV25, LS25, RS24]. rounding [MKL21]. rule [GKS24, God26, WGMZ25]. rules [DGS22, EMS26, NS23, WW24]. Runge [BRK25, FTY24, HS25b, LLZ23, lLW25, Sun25, WLSZ23]. Rupert’s [SY23].

saddle [HKLP23, JCWH21]. saddle-point

[HKLP23]. SAGE [KNT21]. Saint [BL21]. Salem [CW23]. sample [LS20b]. Sampling [DEG22, HLLM25, BHK^{+}25a, BEL25, CEHT23, Doe22, HW21, KNS22, Sun21, THW25]. Sampling-based [HLLM25]. sandpile [HS21a]. SAV [HHZ23, LSL22]. SAV-pressure [LSL22]. scalar [GN25, LNW25]. scale [HLZ23, LN22, MS25b]. Scaled [NS23, CGT25, SH21]. scattering [BY24, DLL21, MV22, Nic24, PV20, RSM22, SH21]. SCF [LL24]. scheme [BR21, BGH^{+}23, BBV21, BCHR20, BDF21, BL21, CHS22, EGH25, GNT24, GWWZ24, HHZ23, HJQ23, HS23, LQW21, lLZ21, LWW^{+}21, LNW25, MLR22, SW25]. schemes [AMU21, Ben22, BBDM26, Coe24, DGSW22, FM20, GK26, GM21, HLS21, HS25b, HPW20, KBL21, LL20, LMY21, MS20a, ORS22, Zha22, LL20]. Schinzel [LP25]. Schinzel-type [LP25]. Schrödinger [BCF23, BW24, DSS23, FMS24, PYAT21, WZZ22, WY22b]. Schubert [LdCS^{+}21]. Schur [CSX23, DSS23, Fla20, OSS25]. Schwarz [AHPS20, GGS23]. Schwarzian [JST22]. SDEs [LV20, LNW25]. search [BCM^{+}24, CSS24, dSKJV23]. Second [BBM21, EH20, BST21, CH24a, CFS21, HJQ23, KZ23, Kir20, Tra25, Wan25, WZZ22, XLL24, dSKJV23]. second-order [CH24a, CFS21, HJQ23, Wan25, WZZ22, dSKJV23]. sectional [DK21]. seeds [BA24]. Segel [WLF25]. Segrè [HH20]. self [BR21, KL23]. self-avoiding [BR21]. self-consistent [KL23]. selfdual [Dua21]. Selmer [BBK^{+}24]. Semi [AMU21, BW24, HS23, LQW21, SSX26, YCQ20, FM20]. Semi-explicit [AMU21]. semi-implicit [HS23, LQW21]. semi-Lagrangian [YCQ20, FM20]. semi-linear [SSX26]. semi-smooth [BW24]. semidefinite [CJdL21]. Semiexplicit [JO23]. semigroup [BAFG20, BA24]. semilinear [DL22]. Sensitivity [BLM23]. separable [HMXY23, JO23, LZ25]. sequence [BS23, BLN^{+}24]. sequences [AB20, BHKS25, DMM24, DOP23, FGN20, GLN22]. Series [GH22, LwaabAL21, LV20, PNR25]. Serre [Dua21]. set [AMV22, BDF21, GM25b, GG24, XXX25]. set-theoretic [AMV22]. sets [WDM20]. settings [CT22a]. several [WD24]. sextic [DGJ20]. Shape [CD23]. shaped [CDG22]. Sharp [HLZZ21, Nak20, Doe22]. Sharper [BKL^{+}21]. shifted [ACS21, CGT25]. shifts [GK26]. shocks [GK26]. short [BK24, MS20b]. shortening [BL25a]. shortfall [HW21]. Siegel [Fla20, OSS25]. sieve [GPS24, Hel20]. sign [GKPR25, Hua20, CLY23]. simple [BR21, CF23, CSS24, GG24, HJLZ20]. simplex [GI23, LYZ25]. simulated [BP24]. simulation [CKMS26, Chu25]. simulations [MS25b]. simultaneous [BMP21, Wal99, Wal21]. single [LRR23]. single-quadrant [LRR23]. singular [FP22, Hua20, Nak20, WX22, XB24]. singularity [KZ23]. singularly [CJS23]. Six [dCIDV24]. Six-dimensional [dCIDV24]. size [HMXY23, Ros22, XLL24]. Sketched [PSS25]. slice [BDM25]. sliced [TFD25]. slip [GS22]. small [BM25, BA22, HHY20, PS21, WWW21]. smallest [BBKP24]. Smolyak [WGMZ25]. smooth [BW24, BCFM25, FGN20, IGMS23, KK20]. Smoothed [EPR21]. smoothing [LYZ24]. smoothness [EMS26, GKS22, GHS23]. Smyth [Fla20, OSS25]. Sobolev [CLLZ25, DST23, Füh22, GKS22, GKS24, GHS23, HR25, KNS22, MSS24]. solution [BRT23, CHS22, DE25, GHM24, HX24]. solutions [AMV22, CMSV24, FLMSS22, FS22, KÇR23, Wal99, Wal21, Zen23]. solvability [JKK21]. solve [CT20]. solver [BY24, LZ25, MNST20]. solving [AGY21, BT22, COS25, MIJR21, SH21, Ste21]. Some [CCG20, BST21, DGSW22, ERS21, HLS21, KR25, KLM20, LR21, PT22, ZC21]. SONC [KNT21]. Sono [GPS24]. source [DGM20, EN24, QT20]. Space [MNST20, BK23a, CH25b, EV20, Hit21, HY25b, IGMS23, KHR23, KÇR23, YCQ20]. space-time [HY25b, IGMS23, KHR23, KÇR23]. spaces [BJT20, DST23, DK21, DOS20, EG24, Füh22, GKS22, GKS24, GSY23, GHS23, KOG21, KPW21, LS20a, LM22, NS23, Sun21]. Spanning [KNRR22]. Sparse [BB23, COS25, BT22, FRS25, HG26, LC22]. sparsity [EN24]. spatial [GHM24, OR23, TXS20]. specific [AC22]. spectra [DMM20]. Spectral [FW20, GGO20, BFS22, BGPW21, DH20, DSS23, GP24, HLZZ21, MS22]. spectral-null [BGPW21]. spectrum [HS21a]. speed [FM20]. speedup [HH22]. sphere [BKK24, CT22c, LC22, dCIDV24]. Spherical [KR20, BDH^{+}21]. spline [Sun21]. Splines [BHKS25, EMS26]. split [MS20c]. splits [CZ25, FGN20, GLN22]. Splitting [BCJ23, LZ25, BCY20, BFS22, BCF23, BW24, BBV21, CHS22, EH20, HMXY23, JLOS25, LYZ24, ORS22]. SPOD [CKNS21]. square [Gos20, Mar21]. square-free [Mar21]. Squares [BNPP20, Car20, HNP22, HX24, Lap20, MNST20, Ros20, Ros22]. Stability [BR21, Bar20, Ben22, BBDM26, DOS20, FM20, GS23, GS20, HS23, KK20, LQT22, PYAT21, WZZ22, XXX25, BCFM25, CS25, Coe24, Fei22, FP20, GCL23, HKLP23, HX24, LMY21, Li22a, LSL22, Sun25, ORZ21, ST23, WLSZ23, MLR22]. stabilization [CZ25]. stabilization-free [CZ25]. Stabilized [GMS21, AGdS22, BL25a, JKK21, LQW21]. Stable [CFEV22, EV20, CNV21, CS22, GM21, Li22b, LWW^{+}21, SSX26, WX22, WLF25, Zha22]. stage [Sun25]. staggered [HLMT21, HLS21]. standard [BJT20]. states [HY25a]. stationary [Wan25]. statistics [CT22b, ELM^{+}24]. Stein [CS25]. Steiner [BBL20]. step [HMXY23]. stepping [BBP24, GMS21, HJLW20, Li22b]. Stickelberger [BK24]. Stiefel [SLN23]. stiff [AGdS22, HS21b]. still [Ros20]. Stirling [CT22b]. Stochastic [GO20, JZCH24, JW25, KL23, LM22, Sog20, BV23, BCJ23, CHS22, FRS25, HJS24, HS23, KNZ25, Kir20, LMY21, LMS25, SWZ21b, dSKJV23, AK24]. Stokes [BLMM^{+}23, ASZ23, BGH^{+}23, CH24b, DOS20, GS22, GWWZ24, HHZ23, Han23, KHR23, KÇR23, LSL22, LMS25, Reu25, MLR22, ORS24, ST23]. Strang [BBV21]. strategies [VSB26]. strategy [QMSZ25]. Strengthening [BFW21]. stresses [CH25a]. Strictly [Zha22, GG24]. Strong [LMY21, LNW25, OR23]. strongly [BGO24].

Structure [HP21, AHHT23, BL25b, BNPP20, CZ25, GS23, JZCH24]. structure-adapted [JZCH24]. structure-aware [GS23]. Structure-preserving [HP21, AHHT23, BL25b]. structured [EPR21]. structures [KL23, WLL21]. study [LV20]. subdiffusion [BM22, LYZ24]. subdivision [CYB^{+}21, KBL21]. subgroups [BMR24, LMP20, MS20c]. Subresultants [CD23]. subsampled [BKPU24]. subsequences [MS20b]. Subspace [CHW20, GGO20, GHM24, GO20, PSS25, Zim21]. subspaces [HRR25]. sum [Lap20, MKL21]. sum-up [MKL21]. sums [BPT21, BKS25, dSKJV23]. sup [CS22, Fei22, ORZ21, ST23]. Super [BCY20, HP23, PO23, PO24]. Super-localization [HP23]. Super-polynomial [PO23, PO24]. Super-resolution [BCY20]. Supercloseness [CJS23]. Superconvergence [BBV21, HW22, XKLW23, YCQ20]. superconvergent [WLL21]. Superelliptic [OS21]. superpositions [LRY20]. surface [BL25b, Reu25, ORZ21, OX20, Reu22]. surfaces [BVA20, CCG20, Dem22, DS20b, FKRS21, ORS24, vLW23, SF25]. swarming [FKRT24]. Swinnerton [KMwaabAVS23]. Swinnerton-Dyer [KMwaabAVS23]. Sylow [LMP20]. Sylvester [PSS25]. symbols [Ros22]. symmetric [BDOT24, BCCET22, BMP21, CH25a, CH22b, CSX23, CH25b, FW20, FRS25, Kho20, LS23, LS24, MS25b, WZ22]. symmetric-conjugate [BCCET22]. symmetry [BKR20]. symplectic [HPR24, JO23]. system [AGY21, BLMM^{+}23, BL21, CS25, FLMSS22, GWWZ24, LS25, LWW^{+}21]. systems [BT22, BCJ23, CLT22, DGSW22, EPR21, FP22, FS22, GM21, HJLZ20, HP21, HS21b, HS25b, HLZ23, JO23, PS21, Ste21]. syzygy [GLNY20].

tables [MS25b]. tailored [BLM23]. Tamagawa [Rüd22]. Tamed [Coe24]. tangency [DGS26]. tangential [ORS24]. tangents [IL23]. Tarry [CMSV24]. Tate [vBHM20]. Taylor [GS23, Reu25, ORZ21]. tensor [BGV24, BHKS25, CH25b, DS20b, FW20, QMSZ25, SLN23, ZC21]. tensors [CH22b]. tent [DGSW22]. tent-based [DGSW22]. tents [GS23]. Terao [BK23a]. term [DGM20, EN24, GKW24, PT21]. terms [NS23, QT20, ZWZ20]. ternary [AC22, BK22]. test [BFW21]. tests [BB23]. tetrahedra [KR20, MR20]. th [CH20]. their [HSS21, PNR25]. theorem [GLNY20, PT21, BLH22]. theoretic [AMV22, LR21]. Theory [GK26, BHK^{+}25a, DKS25, KNT21, XLL24, vBHM20]. thermal [GGKO23]. Thermodynamically [GGKO23]. theta [MX26]. thin [GGKO23]. thin-film [GGKO23]. three [BGMS23, CH22b, CH22a, CH24b, Dua21, EV20, HL21, MV24a, Ros22, WZSP21]. three-dimensional [WZSP21]. threefolds [Ula24]. threshold [DOP23]. thrice [BBK^{+}24]. thrice-punctured [BBK^{+}24]. tight [DH20]. Time [CRS22, CDZ22, RSM22, BBP24, BM22, BCY20, BFS22, BCF23, BW24, BY24, CH24a, FMS24, FSTW25, GMS21, HW22, HKM25, Hit21, HJLW20, HY25b, IGMS23, KHR23, KÇR23, KRRZ20, LS21, LV20, LL20, Li22b, MNST20, WLSZ23]. time-dependent [FSTW25, KHR23]. time-fractional [BBP24, BM22]. time-space [Hit21]. time-splitting [BCY20, BFS22, BCF23, BW24]. time-stepping [BBP24, GMS21, Li22b]. tolerance [GO20]. tomography [BZ24, QESS24]. too [Zen23]. toolkit [HLRS21]. tools [DS20a]. topography [BL21]. topology [ZL20]. tori [HHY20, Rüd22, HKY22]. Toric [BT22, AMP22]. torsion [DGJ20, GJN20, HSS21, Mas20]. total [BRT23]. total-flux [BRT23]. Totally [WWW21, Fla21, LV22]. tournament [BBKP24]. trace [BB23, CW23, Fla20, FRS25, ORZ21, OSS25, WWW21, XB24]. tracking [CKMS26]. traction [MIJR21]. tradeoff [Hit21]. transfer [DGM20, SH21]. transform [GLFN20]. transformations [ABD20]. transforms [LPP21]. translational [BKS25]. transmission [GSTZ22, GGS23]. transport [ACEL21, BHK^{+}25a, Coe24, FP25, HLLM25, JKK21, Lav21]. transport-dominated [JKK21]. transportation [BZ24]. trapezoidal [GKS24]. Trefftz [DES22, IGMS23]. Tri [BGMS23]. Tri-linear [BGMS23]. triangles [EMS26]. triangular [CS22, HLR23]. Tribonacci [BLN^{+}24]. triples [KR20]. Tropical [IL23, PNR25, AAP22]. Trotter [JLOS25]. truncated [Har24, PSS25]. Two [LN22, SW20, BBK^{+}24, CWS26, DE25, EGH25, FP25, GT23, GJL^{+}24, Gos20, HMXY23, HLZ23, Hul22, LX22, Not20, Wal99, Wal21, ZL20]. two-dimensional [CWS26, GT23, Gos20, ZL20]. two-grid [Not20]. two-layer [DE25]. two-level [LX22]. two-marginal [FP25]. two-point [EGH25]. Two-scale [LN22, HLZ23]. type [AK24, BL25a, CL25, GK26, Hua20, Kop21, LS25, LP25, Sun25, MV22, MM21, MT22, Not22, QT20, SLN23, WZZ22, ZC21].

ultra [CX24]. ultra-weak [CX24]. ultraweak [FSTW25, TXS20]. ultraweak-local [FSTW25, TXS20]. unadjusted [BRK25]. unanchored [GKS22]. Unconditional [Lav21, GCL23]. unconditionally [WLF25]. unfitted [LvBV25]. Unified [XLL24, BL25b, CZ25, KNT21, LZ25]. Uniform [AP24, BFS22, HS25b, LPP21, SH21, SvV20, SW25, WLSZ23, BCF23, HS21b]. uniformly [AB20, CLT22, EMS26, Tra25]. unifying [DKS25]. Uniqueness [HX24, AC22]. unit [HJ21]. unitarily [LL24]. units [BGO24]. unstructured [BLM25, SX21]. updates [Zim21]. upon [GJN20]. Upper [Fla21]. upwind [CWS26, LL20, SH21]. UQ [WGMZ25]. Using [BCM^{+}24, BS22, BBV21, CT22b, DE25, HRR25, JST22, Kho20, KMNN21, MIJR21, PO23, PO24]. usual [Fla21].

value [Ben22, KR25]. valued [MR23, OW20]. values [BKPU24, BRS24, Hua20]. vanishing [MKL21]. variable [BGH^{+}23, FM20, GS20, XB24]. variables [BEL25]. variance [JZCH24]. variational [GNT24, GKMV25, Sog20, SJ24]. varieties [BLH22, HLRS21, Mar21]. vector [MR23, Reu22, WH24, XXX25]. vector-Laplace [Reu22]. vector-valued [MR23]. vectors [Nak20]. velocity [BBDM26]. Venant [BL21]. version [CDG22, HMXY23, MR20]. versus [CYB^{+}21]. vertex [GJL^{+}24]. via [AAP22, Bar21, COS25, CT22c, CKMS26, DEG22, DGS26, DS20b, GdLL25, LdCS^{+}21, Chu25, QMSZ25, QH22]. video [QMSZ25]. Virtual [CH20, KMRB25, DS20b, ZMZW23]. viscosity [BGH^{+}23, FLMSS22]. Vlasov [HPR24]. Vologodsky [Kay22]. Volterra [AK24]. volume [BLM25, BLMM^{+}23, BAL21, BCHR20, EGH25, HLMT21, KR20, LYZ25, WLL21]. vortex [CKMS26]. vorticity [BLM25].

Wasserstein [CDZ22, GNT24, TFD25]. wave [BKRS20, BY24, BFO20, CRS22, CH24a, DL22, GMS21, IGMS23, SX21]. Wavenumber [BCFM25]. Wavenumber-explicit [BCFM25]. ways [BDM25]. Weak [LL21, LQXY24, BFS22, CX24, HLMT21, KÇR23]. weakly [AMU21, LS24]. Weierstrass [RSS23]. Weierstrass-Durand [RSS23]. weigh [LS20b]. weight [PSY20]. weighted [DOS20, EN24, HNP22, KPW21, Ste21]. weights [CKNS21]. Weil [Box21]. well [AV25, BBGS22, Gos20]. well-balanced [Gos20]. well-conditioned [BBGS22]. well-posedness [AV25]. Westervelt [BBP24]. Where [LS20b]. whose [BKM20]. width [BA22]. Wiener [LX25]. without [BKS25, God26, WY22b]. witnesses [Sao22]. words [Ros22]. Worsey [GLN22].

Yang [AMV22].

Zakharov [LS25]. zeros [BMOR21, BLN^{+}24, BPT21, GdLL25, HJLZ20, HJ21, HSW22, MS25b]. zeta [BPT21, HSW22]. zeta-function [BPT21]. Zygmund [EG24].

References

Abreu:2022:AMN

[AAP22] Alex Abreu, Sally Andria, and Marco Pacini. Abel maps for nodal curves via tropical geometry. Mathematics of Computation, 91(336):1971– 2025, July 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                                             URL  https://www.ams.org/
                                                             journals/mcom/2022-91-336/

S0025-5718-2022-03717-0;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03717-0/S0025-5718-2022-
03717-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Andria%
2C%20Sally;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1270337;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
782652.

Almansi:2020:CUD

[AB20] Emilio Almansi and Verónica Becher. Completely uniformly distributed sequences based on de Bruijn sequences. Mathematics of Computation, 89(325): 2537–2551, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03534-0;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03534-
0/S0025-5718-2020-03534-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Almansi%2C%
20Emilio;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=368040.

Anguas:2020:CBE

[ABD20] Luis Miguel Anguas, Maria Isabel Bueno, and Froilán M. Dopico. Conditioning and backward errors of eigenvalues of homogeneous matrix polynomials under Möbius transformations. Mathematics of Computation, 89(322):

767–805, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-322/S0025-5718-2019-
                                              03472-5;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              322/S0025-5718-2019-03472-
                                              5/S0025-5718-2019-03472-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1298899;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=664010;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              708591.

Arul:2024:DIG

[ABG^{+}24] Vishal Arul, Jeremy Booher, Steven R. Groen, Everett W. Howe, Wanlin Li, Vlad Matei, Rachel Pries, and Caleb Springer. Doubly isogenous genus-2 curves with D_{4}-action. Mathematics of Computation, 93(345): 347–381, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03891-1.

Akrivis:2025:QCC

[ABP25] Georgios Akrivis, Sören Bartels, and Christian Palus. Quadratic constraint consistency in the projection-free approximation of harmonic maps and bending isometries. Mathematics of Computation, 94(355):2251– 2269, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
355/S0025-5718-2024-04035-8.

Angelini:2022:MUD

[AC22] Elena Angelini and Luca Chiantini. Minimality and uniqueness for decompositions of specific ternary forms. Mathematics of Computation, 91(334): 973–1006, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2022-91-334/S0025-5718-2021-
03681-9;   https://www.ams.
org/journals/mcom/2022-91-
334/S0025-5718-2021-03681-
9/S0025-5718-2021-03681-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=194958;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=944551.

Alfonsi:2021:AOT

[ACEL21] Aurélien Alfonsi, Rafaël Coyaud, Virginie Ehrlacher, and Damiano Lombardi. Approximation of optimal transport problems with marginal moments constraints. Mathematics of Computation, 90 (328):689–737, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03568-6;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03568-
6/S0025-5718-2020-03568-6.
pdf.

Atallah:2021:ASB

[ACS21] Nabil M. Atallah, Claudio Canuto, and Guglielmo Scovazzi. Analysis of the shifted boundary method for the Poisson problem in domains with corners. Mathematics of Computation, 90(331): 2041–2069, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-331/S0025-5718-2021-
                                              03641-8;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              331/S0025-5718-2021-03641-
                                              8/S0025-5718-2021-03641-8.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1343085;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=44965;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              780205.

Alzobydi:2025:CAD

[AE25] Khaled Alzobydi and Graham Ellis. Cellular approximations to the diagonal map. Mathematics of Computation, 94(352): 953–1002, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              352/S0025-5718-2024-03981-9.

Akbary:2020:EFP

[AF20] Amir Akbary and Forrest J. Francis. Euler’s function on products of primes in a fixed arithmetic progression. Mathematics of Computation, 89(322):

993–1026, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03463-4;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03463-
4/S0025-5718-2019-03463-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Francis%2C%
20Forrest%20J.;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=650700.

Awanou:2023:HPF

[AFGS23] Gerard Awanou, Maurice Fabien, Johnny Guzmán, and Ari Stern. Hybridization and postprocessing in finite element exterior calculus. Mathematics of Computation, 92 (339):79–115, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03743-
1/S0025-5718-2022-03743-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1290444;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=700956;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
775211;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=880437.

Akrivis:2021:HOL

[AFKL21] Georgios Akrivis, Michael Feischl, Balázs Kovács, and Christian Lubich. Higher-order linearly implicit full discretization of the Landau–Lifshitz– Gilbert equation. Mathematics of Computation, 90(329): 995–1038, May 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-329/
S0025-5718-2021-03597-8;
https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03597-8/S0025-5718-2021-
03597-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Akrivis%
2C%20Georgios;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Feischl%
2C%20Michael;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Kovacs%
2C%20Balazs;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Lubich%
2C%20Christian.

Asif:2022:CPP

[AFPwaabAVS22] Sualeh Asif, Francesc Fité, Dylan Pentland, and with an appendix by A. V. Sutherland. Computing L-polynomials of Picard curves from Cartier– Manin matrices. Mathematics of Computation, 91(334): 943–971, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-334/
S0025-5718-2021-03675-3;
https://www.ams.org/journals/
mcom/2022-91-334/S0025-5718-
2021-03675-3/S0025-5718-2021-
03675-3.pdf;   https://www.
ams.org/mathscinet/search/

authors.html?authorName=Asif%
2C%20Sualeh;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Pentland%
2C%20Dylan;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=852273;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
995332.

Abdulle:2022:ESM

[AGdS22] Assyr Abdulle, Marcus J. Grote, and Giacomo Rosilho de Souza. Explicit stabilized multirate method for stiff differential equations. Mathematics of Computation, 91(338):2681–2714, November 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-338/
S0025-5718-2022-03753-4/S0025-
5718-2022-03753-4.pdf;   https:
//www.ams.org/mathscinet/search/
authors.html?authorName=Abdulle%
2C%20Assyr;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1330077;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
360720.

Akinshin:2021:GEA

[AGY21] Andrey Akinshin, Gil Goldman, and Yosef Yomdin. Geometry of error amplification in solving the Prony system with near-colliding nodes. Mathematics of Computation, 90(327): 267–302, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-327/

S0025-5718-2020-03571-6;
https://www.ams.org/journals/
mcom/2021-90-327/S0025-5718-
2020-03571-6/S0025-5718-2020-
03571-6.pdf.

Ayi:2023:SPN

[AHHT23] Nathalie Ayi, Maxime Herda, Hélène Hivert, and Isabelle Tristani. On a structure-preserving numerical method for fractional Fokker–Planck equations. Mathematics of Computation, 92(340): 635–693, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03789-
3/S0025-5718-2022-03789-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Hivert%2C%
20Helene;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=1080256;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1154786;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1198167.

Antonietti:2020:ABM

[AHPS20] Paola F. Antonietti, Paul Houston, Giorgio Pennesi, and Endre Süli. An agglomerationbased massively parallel nonoverlapping additive Schwarz preconditioner for high-order discontinuous Galerkin methods on polytopic grids. Mathematics of Computation, 89(325):2047– 2083, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03510-8;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03510-
8/S0025-5718-2020-03510-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Antonietti%
2C%20Paola%20F.;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Suli%
2C%20Endre;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1298983;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
635107.

Alfonsi:2024:ASV

[AK24] Aurélien Alfonsi and Ahmed Kebaier. Approximation of Stochastic Volterra Equations with kernels of completely monotone type. Mathematics of Computation, 93 (346):643–677, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03911-4.

Adzaga:2024:CQP

[AKMJ^{+}24] Nikola Adzaga, Timo Keller, Philippe Michaud-Jacobs, Filip Najman, Ekin Ozman, and Borna Vukorepa. Computing quadratic points on modular curves X_{0}(N). Mathematics of Computation, 93 (347):1371–1397, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03902-3.

Auel:2025:CCF

[AKPW25] Asher Auel, Avinash Kulkarni, Jack Petok, and Jonah Weinbaum. A census of cubic fourfolds over 𝔽_{2}. Mathematics of Computation, 94(354): 2089–2112, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04010-3.

Adell:2020:RAE

[AL20] José A. Adell and Alberto Lekuona. Rational approximation to Euler’s constant at a geometric rate of convergence. Mathematics of Computation, 89(325): 2553–2561, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-325/S0025-5718-2020-
                                              03528-5;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              325/S0025-5718-2020-03528-
                                              5/S0025-5718-2020-03528-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=340766;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=663604.

Abrams:2025:EGI

[ALPS25] Aaron Abrams, Zeph Landau, Jamie Pommersheim, and Nikhil Srivastava. On eigenvalue gaps of integer matrices. Mathematics of Computation, 94(352): 853–862, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03905-4.

Amzallag:2022:DBT

[AMP22] Eli Amzallag, Andrei Minchenko, and Gleb Pogudin. Degree bound for toric envelope of a linear algebraic group. Mathematics of Computation, 91(335): 1501–1519, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03695-9;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03695-9/S0025-5718-2021-
03695-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1319513;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
807995;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=948033.

Altmann:2021:SED

[AMU21] R. Altmann, R. Maier, and B. Unger. Semi-explicit discretization schemes for weakly coupled elliptic-parabolic problems. Mathematics of Computation, 90(329):1089–1118, May 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2021-
90-329/S0025-5718-2021-03608-
X;   https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03608-X/S0025-5718-2021-
03608-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Altmann%

                                              2C%20R.;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?authorName=Maier%2C%20R.
                                              ;   https://www.ams.org/mathscinet/
                                              search/authors.html?authorName=
                                              Unger%2C%20B..

Akgun:2022:EST

[AMV22] Ö. Akgün, M. Mereb, and L. Vendramin. Enumeration of settheoretic solutions to the Yang– Baxter equation. Mathematics of Computation, 91(335): 1469–1481, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-335/
                                              S0025-5718-2022-03696-6;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-335/S0025-5718-
                                              2022-03696-6/S0025-5718-2022-
                                              03696-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Akgun%
                                              2C%20O.;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=829575;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=870581.

Ainsworth:2024:UPH

[AP24] Mark Ainsworth and Charles Parker. Uniform preconditioners for high order finite element approximations of planar linear elasticity. Mathematics of Computation, 93(349):2067– 2102, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              349/S0025-5718-2023-03926-6.

Ambrose:2023:CBI

[ASZ23] David M. Ambrose, Michael Siegel, and Keyang Zhang. Convergence of the boundary integral method for interfacial Stokes flow. Mathematics of Computation, 92 (340):695–748, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03787-
X/S0025-5718-2022-03787-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Zhang%2C%20Keyang;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
294740;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=720777.

Aggarwal:2025:CNA

[AV25] Aekta Aggarwal and Ganesh Vaidya. Convergence of the numerical approximations and wellposedness: Nonlocal conservation laws with rough flux. Mathematics of Computation, 94(352): 585–610, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03976-5.

Avdispahic:2021:EBH

[Avd21] Muharem Avdispahić. Effective bounds for Huber’s constant and Faltings’s delta function. Mathematics of Computation, 90(331): 2381–2414, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/

journals/mcom/2021-90-331/
S0025-5718-2021-03631-5;
https://www.ams.org/journals/
mcom/2021-90-331/S0025-5718-
2021-03631-5/S0025-5718-2021-
03631-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=28365.

Bingane:2022:ESO

[BA22] Christian Bingane and Charles Audet. The equilateral small octagon of maximal width. Mathematics of Computation, 91(336): 2027–2040, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03733-9;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03733-9/S0025-5718-2022-
03733-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Bingane%
2C%20Christian;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=619525.

Bras-Amoros:2024:SGG

[BA24] Maria Bras-Amorós. On the seeds and the great-grandchildren of a numerical semigroup. Mathematics of Computation, 93(345): 411–441, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
345/S0025-5718-2023-03881-9.

Bras-Amoros:2020:RGD

[BAFG20] Maria Bras-Amorós and Julio Fernández-González. The rightgenerators descendant of a numerical semigroup. Mathematics of Computation, 89(324): 2017–2030, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03502-9;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03502-
9/S0025-5718-2020-03502-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Bras-Amoros%
2C%20Maria;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=704910.

Ben-Artzi:2021:CFV

[BAL21] Matania Ben-Artzi and Jiequan Li. Consistency of finite volume approximations to nonlinear hyperbolic balance laws. Mathematics of Computation, 90(327): 141–169, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03569-8;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03569-
8/S0025-5718-2020-03569-8.
pdf.

Barvinok:2020:SCM

[Bar20]  Alexander Barvinok.  Stabil-

ity and complexity of mixed discriminants. Mathematics of Computation, 89(322):717–735, March 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

                                              /www.ams.org/AMSMathViewer;
                                               https://www.ams.org/journals/
                                              mcom/2020-89-322/S0025-5718-
                                              2019-03468-3;   https://www.
                                              ams.org/journals/mcom/2020-
                                              89-322/S0025-5718-2019-03468-
                                              3/S0025-5718-2019-03468-3.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=237145.

Bartels:2021:EEC

[Bar21] Sören Bartels. Error estimates for a class of discontinuous Galerkin methods for nonsmooth problems via convex duality relations. Mathematics of Computation, 90 (332):2579–2602, November 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-332/S0025-5718-2021-
                                              03656-X;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              332/S0025-5718-2021-03656-
                                              X/S0025-5718-2021-03656-X.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Bartels%2C%
                                              20Soren.

Brysiewicz:2023:STT

[BB23] Taylor Brysiewicz and Michael Burr. Sparse trace tests. Mathematics of Computation, 92(344): 2893–2922, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              344/S0025-5718-2023-03849-2.

Botti:2026:SCP

[BBDM26] Lorenzo Botti, Michele Botti, Daniele A. Di Pietro, and Francesco Carlo Massa. Stability, convergence, and pressurerobustness of numerical schemes for incompressible flows with hybrid velocity and pressure. Mathematics of Computation, 95 (357):1–28, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04049-3.

Ballester-Bolinches:2024:ELB

[BBERPC24] A. Ballester-Bolinches, R. Esteban-Romero, and V. Pérez-Calabuig. Enumeration of left braces with additive group C_{4} × C_{4} × C_{4}. Mathematics of Computation, 93 (346):911–919, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03871-6.

Beltran:2022:HWC

[BBGS22] Carlos Beltrán, Laurent Bétermin, Peter Grabner, and Stefan Steinerberger. How well-conditioned can the eigenvector problem be? Mathematics of Computation, 91 (335):1237–1245, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03706-0;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03706-0/S0025-5718-2021-
03706-0.pdf;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Betermin%
                                              2C%20Laurent;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=293266;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              764504;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=869041.

Best:2024:RSE

[BBK^{+}24] Alex J. Best, L. Alexander Betts, Theresa Kumpitsch, Martin Lüdtke, Angus W. McAndrew, Lie Qian, Elie Studnia, and Yujie Xu. Refined Selmer equations for the thrice-punctured line in depth two. Mathematics of Computation, 93(347): 1497–1527, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              347/S0025-5718-2023-03898-4.

Bellitto:2024:SCT

[BBKP24] Thomas Bellitto, Nicolas Bousquet, Adam Kabela, and Théo Pierron. The smallest 5chromatic tournament. Mathematics of Computation, 93(345): 443–458, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03887-X.

Bonnivard:2020:NAS

[BBL20] Matthieu Bonnivard, Elie Bretin, and Antoine Lemenant. Numerical approximation of the Steiner problem in dimension 2 and 3. Mathematics of Computation, 89 (321):1–43, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03442-7;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03442-
7/S0025-5718-2019-03442-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=889777;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=932768;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
937006.

Bonnans:2021:SOM

[BBM21] J. Frédéric Bonnans, Guillaume Bonnet, and Jean-Marie Mirebeau. Second order monotone finite differences discretization of linear anisotropic differential operators. Mathematics of Computation, 90(332):2671– 2703, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-332/
S0025-5718-2021-03671-6;
https://www.ams.org/journals/
mcom/2021-90-332/S0025-5718-
2021-03671-6/S0025-5718-2021-
03671-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Bonnans%
2C%20J.%20Frederic;   https:
//www.ams.org/mathscinet/search/
authors.html?authorName=Bonnet%
2C%20Guillaume;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=908991.

Baker:2024:NAT

[BBP24] Katherine Baker, Lehel Banjai, and Mariya Ptashnyk. Numerical analysis of a timestepping method for the Westervelt equation with time-fractional damping. Mathematics of Computation, 93(350):2711–2743, November 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2024-93-350/
                                              S0025-5718-2024-03945-5.

Bertoli:2021:SSS

[BBV21] Guillaume Bertoli, Christophe Besse, and Gilles Vilmart. Superconvergence of the Strang splitting when using the Crank– Nicolson scheme for parabolic PDEs with Dirichlet and oblique boundary conditions. Mathematics of Computation, 90(332): 2705–2729, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-332/S0025-5718-2021-
                                              03664-9;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              332/S0025-5718-2021-03664-
                                              9/S0025-5718-2021-03664-9.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1384999;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=639642;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              798890.

Blanes:2022:SCC

[BCCET22] S. Blanes, F. Casas, P. Chartier, and A. Escorihuela-Tomàs. On symmetric-conjugate composition methods in the numerical integration of differential equations. Mathematics of Computation, 91 (336):1739–1761, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2021-03715-1;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2021-03715-1/S0025-5718-2021-
03715-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Escorihuela-
Tomas%2C%20A.;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=314104;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
335517;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=633218.

Billerey:2025:DPG

[BCDF25] Nicolas Billerey, Imin Chen, Luis Dieulefait, and Nuno Freitas. On Darmon’s program for the generalized Fermat equation, II. Mathematics of Computation, 94 (354):1977–2003, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
354/S0025-5718-2024-04012-7.

Bao:2023:IUE

[BCF23] Weizhu Bao, Yongyong Cai, and Yue Feng. Improved uniform error bounds of the time-splitting methods for the long-time (nonlinear) Schrödinger equation. Mathematics of Computation, 92 (341):1109–1139, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2022-03801-
1/S0025-5718-2022-03801-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Feng%2C%20Yue;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
354327;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=819002.

Bernkopf:2025:WES

[BCFM25] M. Bernkopf, T. Chaumont-Frelet, and J. M. Melenk. Wavenumber-explicit stability and convergence analysis of hp finite element discretizations of Helmholtz problems in piecewise smooth media. Mathematics of Computation, 94(351): 73–122, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03958-3.

Bessemoulin-Chatard:2020:HDL

[BCHR20] Marianne Bessemoulin-Chatard, Maxime Herda, and Thomas Rey. Hypocoercivity and diffusion limit of a finite volume scheme for linear kinetic equations. Mathematics of Computation, 89(323): 1093–1133, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03490-7;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03490-
7/S0025-5718-2019-03490-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Rey%2C%20Thomas;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1154786;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=960946.

Brehier:2023:SIS

[BCJ23] Charles-Edouard Bréhier, David Cohen, and Tobias Jahnke. Splitting integrators for stochastic Lie–Poisson systems. Mathematics of Computation, 92(343): 2167–2216, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03829-7.

Barbero:2021:PRQ

[BCM21] Stefano Barbero, Umberto Cerruti, and Nadir Murru. Periodic representations for quadratic irrationals in the field of padic numbers. Mathematics of Computation, 90(331):2267– 2280, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03640-6;   https://www.ams.
org/journals/mcom/2021-90-
331/S0025-5718-2021-03640-

                                              6/S0025-5718-2021-03640-6.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=46835;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=891098;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              905269.

Bogfjellmo:2024:UAS

[BCM^{+}24] Geir Bogfjellmo, Elena Celledoni, Robert I. McLachlan, Brynjulf Owren, and G. R. W. Quispel. Using aromas to search for preserved measures and integrals in Kahan’s method. Mathematics of Computation, 93(348): 1633–1653, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              348/S0025-5718-2023-03921-7.

Bao:2020:SRT

[BCY20] Weizhu Bao, Yongyong Cai, and Jia Yin. Super-resolution of time-splitting methods for the Dirac equation in the nonrelativistic regime. Mathematics of Computation, 89(325):2141– 2173, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-325/S0025-5718-2020-
                                              03536-4;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              325/S0025-5718-2020-03536-
                                              4/S0025-5718-2020-03536-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1169777;   https:

//www.ams.org/mathscinet/search/
authors.html?mrauthid=354327;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
819002.

Bialecki:2021:CAF

[BDF21] Bernard Bialecki, Maksymilian Dryja, and Ryan I. Fernandes. Convergence analysis of the finite difference ADI scheme for the heat equation on a convex set. Mathematics of Computation, 90(332):2757– 2784, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-332/
S0025-5718-2021-03653-4;
https://www.ams.org/journals/
mcom/2021-90-332/S0025-5718-
2021-03653-4/S0025-5718-2021-
03653-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Dryja%
2C%20Maksymilian;   https:/
/www.ams.org/mathscinet/search/
authors.html?mrauthid=252450;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
312570.

Boyvalenkov:2021:BSC

[BDH^{+}21] P. G. Boyvalenkov, P. D. Dragnev, D. P. Hardin, E. B. Saff, and M. M. Stoyanova. Bounds for spherical codes: the Levenshtein framework lifted. Mathematics of Computation, 90(329): 1323–1356, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-329/
S0025-5718-2021-03621-2;

https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03621-2/S0025-5718-2021-
03621-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Boyvalenkov%
2C%20P.%20G.;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Dragnev%
2C%20P.%20D.;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Hardin%
2C%20D.%20P.;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Saff%
2C%20E.%20B.;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Stoyanova%
2C%20M.%20M..

Bonaldi:2024:NAM

[BDM24] Francesco Bonaldi, Jérôme Droniou, and Roland Masson. Numerical analysis of a mixeddimensional poromechanical model with frictionless contact at matrix-fracture interfaces. Mathematics of Computation, 93(349): 2103–2134, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
349/S0025-5718-2024-03949-2.

Brandenburg:2025:BWS

[BDM25] Marie-Charlotte Brandenburg, Jesús A. De Loera, and Chiara Meroni. The best ways to slice a polytope. Mathematics of Computation, 94(352): 1003–1042, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

org/journals/mcom/2025-94-
352/S0025-5718-2024-04006-1.

Bachmayr:2024:PAS

[BDOT24] Markus Bachmayr, Geneviève Dusson, Christoph Ortner, and Jack Thomas. Polynomial approximation of symmetric functions. Mathematics of Computation, 93(346):811–839, March 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2024-
93-346/S0025-5718-2023-03868-
6.

Bik:2020:MR

[BDOV20] Arthur Bik, Jan Draisma, Alessandro Oneto, and Emanuele Ventura. The monic rank. Mathematics of Computation, 89(325): 2481–2505, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03512-1;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03512-
1/S0025-5718-2020-03512-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Ventura%2C%
20Emanuele;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1087088;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1297289;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=683807.

Buse:2020:FIR

[BDS20] Laurent Busé, Alexandru Dimca, and Gabriel Sticlaru. Freeness and invariants of rational plane curves. Mathematics of Computation, 89(323): 1525–1546, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03495-6;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03495-
6/S0025-5718-2019-03495-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Buse%2C%20Laurent;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
58125;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=997310.

Bachmayr:2021:EDL

[BEKU21] Markus Bachmayr, Henrik Eisenmann, Emil Kieri, and André Uschmajew. Existence of dynamical low-rank approximations to parabolic problems. Mathematics of Computation, 90(330): 1799–1830, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-330/S0025-5718-2021-
03626-1;   https://www.ams.
org/journals/mcom/2021-90-
330/S0025-5718-2021-03626-
1/S0025-5718-2021-03626-1.
pdf;   https://www.ams.org/

mathscinet/search/authors.
html?authorName=Eisenmann%
2C%20Henrik;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Uschmajew%
2C%20Andre;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1078345;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
881952.

Blel:2025:ISC

[BEL25] Mohamed-Raed Blel, Virginie Ehrlacher, and Tony Lelièvre. Influence of sampling on the convergence rates of greedy algorithms for parameter-dependent random variables. Mathematics of Computation, 94(352): 863–908, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03979-0.

Benoit:2022:SFD

[Ben22] Antoine Benoit. Stability of finite difference schemes approximation for hyperbolic boundary value problems in an interval. Mathematics of Computation, 91 (335):1171–1212, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03698-4;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03698-4/S0025-5718-2021-
03698-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1060440.

Blaustein:2024:DFH

[BF24] Alain Blaustein and Francis Filbet. On a discrete framework of hypocoercivity for kinetic equations. Mathematics of Computation, 93(345): 163–202, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03862-5.

Burman:2020:FED

[BFO20] Erik Burman, Ali Feizmohammadi, and Lauri Oksanen. A finite element data assimilation method for the wave equation. Mathematics of Computation, 89 (324):1681–1709, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-324/S0025-5718-2020-
                                              03508-X;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              324/S0025-5718-2020-03508-
                                              X/S0025-5718-2020-03508-X.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Feizmohammadi%
                                              2C%20Ali;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=602430;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=906909.

Bao:2022:UEB

[BFS22] Weizhu Bao, Yue Feng, and Chunmei Su. Uniform error bounds of time-splitting spectral methods for the long-time dynamics of the nonlinear Klein–Gordon equation with weak nonlinearity. Mathematics of Computation, 91 (334):811–842, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-334/
S0025-5718-2021-03694-7;
https://www.ams.org/journals/
mcom/2022-91-334/S0025-5718-
2021-03694-7/S0025-5718-2021-
03694-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Feng%
2C%20Yue;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Su%2C%20Chunmei;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
354327.

Baillie:2021:SBP

[BFW21] Robert Baillie, Andrew Fiori, and Samuel S. Wagstaff, Jr. Strengthening the Baillie–PSW primality test. Mathematics of Computation, 90(330): 1931–1955, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-330/S0025-5718-2021-
03616-9;   https://www.ams.
org/journals/mcom/2021-90-
330/S0025-5718-2021-03616-
9/S0025-5718-2021-03616-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=179915;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=407930;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
997971.

Batteux:2023:CMS

[BGH^{+}23] L. Batteux, T. Gallouët, R. Herbin, J. C. Latché, and P. Poullet. Convergence of the MAC scheme for the incompressible Navier– Stokes equations with variable density and viscosity. Mathematics of Computation, 92(342): 1595–1631, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03803-0.

Brinkmann:2022:MGK

[BGM22] Gunnar Brinkmann, Jan Goedgebeur, and Brendan D. McKay. The minimality of the Georges– Kelmans graph. Mathematics of Computation, 91(335): 1483–1500, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-335/
                                              S0025-5718-2021-03701-1;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-335/S0025-5718-
                                              2021-03701-1/S0025-5718-2021-
                                              03701-1.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=122480;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              252206;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=945364.

Buse:2023:TLB

[BGMS23] Laurent Busé, Pablo González-Mazón, and Josef Schicho. Trilinear birational maps in dimension three. Mathematics of Computation, 92(342):

1837–1866, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03804-2.

Bakshi:2024:RGA

[BGO24] Gurmeet K. Bakshi, Jyoti Garg, and Gabriela Olteanu. Rational group algebras of generalized strongly monomial groups: Primitive idempotents and units. Mathematics of Computation, 93 (350):3027–3058, November 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03937-6.

Buhler:2021:LPS

[BGPW21] Joe Buhler, Shahar Golan, Rob Pratt, and Stan Wagon. Littlewood polynomials, spectralnull codes, and equipowerful partitions. Mathematics of Computation, 90(329):1435– 1453, May 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-329/
S0025-5718-2021-03612-1;
https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03612-1/S0025-5718-2021-
03612-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Buhler%
2C%20Joe;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Golan%2C%20Shahar;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Pratt%2C%20Rob;   https://

                                              www.ams.org/mathscinet/search/
                                              authors.html?authorName=Wagon%
                                              2C%20Stan.

Beorchia:2024:ETE

[BGV24] Valentina Beorchia, Francesco Galuppi, and Lorenzo Venturello. Equations of tensor eigenschemes. Mathematics of Computation, 93 (346):589–602, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              346/S0025-5718-2023-03882-0.

Baptista:2025:ATF

[BHK^{+}25a] Ricardo Baptista, Bamdad Hosseini, Nikola B. Kovachki, Youssef Marzouk, and Amir Sagiv. An approximation theory framework for measure-transport sampling algorithms. Mathematics of Computation, 94(354): 1863–1909, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04013-9.

Bourdon:2025:TCI

[BHK^{+}25b] Abbey Bourdon, Sachi Hashimoto, Timo Keller, Zev Klagsbrun, David Lowry-Duda, Travis Morrison, Filip Najman, and Himanshu Shukla. Towards a classification of isolated j-invariants. Mathematics of Computation, 94(351): 447–473, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-03956-X.

With an appendix by Maarten Derickx and Mark van Hoeij.

Bonizzoni:2025:DTP

[BHKS25] Francesca Bonizzoni, Kaibo Hu, Guido Kanschat, and Duygu Sap. Discrete tensor product BGG sequences: Splines and finite elements. Mathematics of Computation, 94(352): 517–549, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03969-8.

Burman:2024:LRE

[BHL24] Erik Burman, Peter Hansbo, and Mats G. Larson. Low regularity estimates for CutFEM approximations of an elliptic problem with mixed boundary conditions. Mathematics of Computation, 93 (345):35–54, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
345/S0025-5718-2023-03875-3.

Bochenski:2020:NSC

[BJT20] Maciej Bocheński, Piotr Jastrzebski, and Aleksy Tralle. Nonexistence of standard compact Clifford–Klein forms of homogeneous spaces of exceptional Lie groups. Mathematics of Computation, 89(323): 1487–1499, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/

2020-89-323/S0025-5718-2019-
03493-2;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03493-
2/S0025-5718-2019-03493-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Jastrzebski%
2C%20Piotr;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1069305;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
199440.

Bringmann:2022:CNR

[BK22] Kathrin Bringmann and Ben Kane. Class numbers and representations by ternary quadratic forms with congruence conditions. Mathematics of Computation, 91(333):295–329, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
333/S0025-5718-2021-03648-
0;   https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03648-0/S0025-5718-2021-
03648-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=774752;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
789505.

Barakat:2023:CNL

[BK23a] Mohamed Barakat and Lukas Kühne. Computing the nonfree locus of the moduli space of arrangements and Terao’s freeness conjecture. Mathematics of Computation, 92(341): 1431–1452, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2023-03812-
1/S0025-5718-2023-03812-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Kuhne%2C%20Lukas;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
706483.

Bartels:2023:EEE

[BK23b] Sören Bartels and Alex Kaltenbach.

Explicit and efficient error estimation for convex minimization problems. Mathematics of Computation, 92(343):2247– 2279, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03821-2.

Bernard:2024:SBS

[BK24] Olivier Bernard and Radan Kucera. A short basis of the Stickelberger ideal of a cyclotomic field. Mathematics of Computation, 93(346): 887–909, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03863-7.

Bogachev:2024:KSP

[BKK24] Nikolay Bogachev, Alexander Kolpakov, and Alex Kontorovich. Kleinian sphere packings, reflection groups, and arithmeticity. Mathematics of Computation, 93(345):505–521, January 2024.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
345/S0025-5718-2023-03858-3.

Bolin:2024:RNA

[BKKS24] David Bolin, Mihály Kovács, Vivek Kumar, and Alexandre B. Simas. Regularity and numerical approximation of fractional elliptic differential equations on compact metric graphs. Mathematics of Computation, 93(349): 2439–2472, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
349/S0025-5718-2023-03929-1.

Broadbent:2021:SBC

[BKL^{+}21] Samuel Broadbent, Habiba Kadiri, Allysa Lumley, Nathan Ng, and Kirsten Wilk. Sharper bounds for the Chebyshev function θ(x). Mathematics of Computation, 90(331):2281–2315, September 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-331/
S0025-5718-2021-03643-1;
https://www.ams.org/journals/
mcom/2021-90-331/S0025-5718-
2021-03643-1/S0025-5718-2021-
03643-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Broadbent%
2C%20Samuel;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1084240;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1403478;   https://www.ams.

org/mathscinet/search/authors.
html?mrauthid=721483;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=760548.

Birgin:2020:IEC

[BKM20] E. G. Birgin, N. Krejić, and J. M.

Martínez. Iteration and evaluation complexity for the minimization of functions whose computation is intrinsically inexact. Mathematics of Computation, 89(321): 253–278, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03445-2;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03445-
2/S0025-5718-2019-03445-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Krejic%2C%
20N.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=120570;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=662583.

Bartel:2024:RFV

[BKPU24] Felix Bartel, Lutz Kämmerer, Daniel Potts, and Tino Ullrich. On the reconstruction of functions from values at subsampled quadrature points. Mathematics of Computation, 93(346): 785–809, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03896-0.

Bohm:2020:CGF

[BKR20] Janko Böhm, Simon Keicher, and Yue Ren. Computing GIT-fans with symmetry and the Mori chamber decomposition of \overline{M}_{0,6}. Mathematics of Computation, 89(326):3003– 3021, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2020-89-326/
                                              S0025-5718-2020-03546-7;
                                              https://www.ams.org/journals/
                                              mcom/2020-89-326/S0025-5718-
                                              2020-03546-7/S0025-5718-2020-
                                              03546-7.pdf.

Bause:2020:PPG

[BKRS20] M. Bause, U. Köcher, F. A. Radu, and F. Schieweck. Post-processed Galerkin approximation of improved order for wave equations. Mathematics of Computation, 89 (322):595–627, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-322/S0025-5718-2019-
                                              03464-6;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              322/S0025-5718-2019-03464-
                                              6/S0025-5718-2019-03464-6.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Kocher%2C%
                                              20U.;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=155960;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=667155;
                                               https://www.ams.org/mathscinet/

search/authors.html?mrauthid=
740659.

Bars:2023:BQM

[BKS23] Francesc Bars, Mohamed Kamel, and Andreas Schweizer. Bielliptic quotient modular curves of X_{0}(N). Mathematics of Computation, 92(340):895–929, March 2023. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2023-
92-340/S0025-5718-2022-03800-
X/S0025-5718-2022-03800-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1217615;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=362896;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
647724.

Buchheit:2025:CLS

[BKS25] Andreas A. Buchheit, Torsten Keßler, and Kirill Serkh. On the computation of lattice sums without translational invariance. Mathematics of Computation, 94(355):2533–2574, September 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2025-
94-355/S0025-5718-2024-04024-
3.

Bouchut:2021:CKH

[BL21] François Bouchut and Xavier Lhébrard. Convergence of the kinetic hydrostatic reconstruction scheme for the Saint Venant system with topography. Mathematics of Computation, 90(329): 1119–1153, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-329/S0025-5718-2020-
03600-X;   https://www.ams.
org/journals/mcom/2021-90-
329/S0025-5718-2020-03600-
X/S0025-5718-2020-03600-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Bouchut%2C%
20Francois;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Lhebrard%
2C%20Xavier.

Bai:2025:CSP

[BL25a] Genming Bai and Buyang Li. Convergence of a stabilized parametric finite element method of the Barrett–Garcke–Nürnberg type for curve shortening flow. Mathematics of Computation, 94(355):2151–2220, September 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2025-
94-355/S0025-5718-2024-04019-
X.

Bao:2025:USP

[BL25b] Weizhu Bao and Yifei Li. A unified structure-preserving parametric finite element method for anisotropic surface diffusion. Mathematics of Computation, 94(355):2113–2149, September 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2025-
94-355/S0025-5718-2024-04022-
X.

Blanco:2022:AHI

[Bla22] Guillem Blanco. An algorithm for Hodge ideals. Mathematics of Computation, 91(338):2955– 2967, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03764-
9/S0025-5718-2022-03764-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1308706.

Barakat:2022:AAC

[BLH22] Mohamed Barakat and Markus Lange-Hegermann. An algorithmic approach to Chevalley’s Theorem on images of rational morphisms between affine varieties. Mathematics of Computation, 91(333):451– 490, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03632-7;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03632-7/S0025-5718-2021-
03632-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=706483;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
937506.

Birgin:2023:SAT

[BLM23] E. G. Birgin, A. Laurain, and T. C. Menezes. Sensitivity analysis and tailored design of minimization diagrams. Mathematics of Computation, 92(344):2715– 2768, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              344/S0025-5718-2023-03839-X.

Bueno:2024:IRM

[BLM24] L. F. Bueno, F. Larreal, and J. M. Martínez. Inexact restoration for minimization with inexact evaluation both of the objective function and the constraints. Mathematics of Computation, 93(345): 293–326, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03855-8.

Barsukow:2025:NCV

[BLM25] Wasilij Barsukow, Raphaël Loubère, and Pierre-Henri Maire. A node-conservative vorticity preserving finite volume method for linear acoustics on unstructured grids. Mathematics of Computation, 94(355):2299– 2343, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              355/S0025-5718-2024-04020-6.

Basaric:2023:EEF

[BLMM^{+}23] Danica Basarić, Mária Lukácová-Medvidová, Hana Mizerová, Bangwei She, and Yuhuan Yuan. Error estimates of a finite volume method for the compressible Navier–Stokes–Fourier system. Mathematics of Computation, 92 (344):2543–2574, November 2023.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03852-2.

Bilu:2024:AZT

[BLN^{+}24] Yuri Bilu, Florian Luca, Joris Nieuwveld, Joël Ouaknine, and James Worrell. On the p-adic zeros of the Tribonacci sequence. Mathematics of Computation, 93 (347):1333–1353, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03893-5.

Banjai:2022:PEA

[BM22] Lehel Banjai and Charalambos G. Makridakis. A posteriori error analysis for approximations of time-fractional subdiffusion problems. Mathematics of Computation, 91(336): 1711–1737, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03723-6;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03723-6/S0025-5718-2022-
03723-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=289627;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
702930.

Battistoni:2025:GPI

[BM25] Francesco Battistoni and Giuseppe Molteni. Generalized Pohst inequality and small regulators.

Mathematics of Computation, 94(351):475–504, January 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03954-6.

Bennett:2021:CZD

[BMOR21] Michael A. Bennett, Greg Martin, Kevin O’Bryant, and Andrew Rechnitzer. Counting zeros of Dirichlet L-functions. Mathematics of Computation, 90(329): 1455–1482, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-329/S0025-5718-2021-
03599-1;   https://www.ams.
org/journals/mcom/2021-90-
329/S0025-5718-2021-03599-
1/S0025-5718-2021-03599-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Bennett%2C%
20Michael%20A.;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Martin%
2C%20Greg;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=O%27Bryant%
2C%20Kevin;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Rechnitzer%
2C%20Andrew.

Brodnik:2021:SCP

[BMP21] Andrej Brodnik, Aleksander Malnic, and Rok Pozar. The simultaneous conjugacy problem in the symmetric group. Mathematics of Computation, 90(332): 2977–2995, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03637-6;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03637-
6/S0025-5718-2021-03637-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Malnic%2C%
20Aleksander;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Pozar%
2C%20Rok;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=623283.

Berghaus:2024:CMF

[BMR24] David Berghaus, Hartmut Monien, and Danylo Radchenko. On the computation of modular forms on noncongruence subgroups. Mathematics of Computation, 93(347): 1399–1425, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03903-5.

Banwait:2024:CIE

[BNP24] Barinder S. Banwait, Filip Najman, and Oana Padurariu. Cyclic isogenies of elliptic curves over fixed quadratic fields. Mathematics of Computation, 93(346): 841–862, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03894-7.

Bonizzoni:2020:FLS

[BNPP20] Francesca Bonizzoni, Fabio Nobile, Ilaria Perugia, and Davide Pradovera. Fast Least-Squares Padeé approximation of problems with normal operators and meromorphic structure. Mathematics of Computation, 89(323): 1229–1257, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-323/S0025-5718-2020-
                                              03511-X;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              323/S0025-5718-2020-03511-
                                              X/S0025-5718-2020-03511-X.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Pradovera%
                                              2C%20Davide;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1081834;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              366660;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=650310.

Borthagaray:2023:RBP

[BNWX23] Juan Pablo Borthagaray, Ricardo H. Nochetto, Shuonan Wu, and Jinchao Xu. Robust BPX preconditioner for fractional Laplacians on bounded Lipschitz domains. Mathematics of Computation, 92(344):2439– 2473, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              344/S0025-5718-2023-03857-1.

MC-06670010

[Bör25] Steffen Börm. Adaptive fast multiplication of H^{2}-matrices. Mathematics of Computation, 94(352): 825–852, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03978-9.

Box:2021:QPM

[Box21] Josha Box. Quadratic points on modular curves with infinite Mordell–Weil group. Mathematics of Computation, 90(327): 321–343, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03547-9;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03547-
9/S0025-5718-2020-03547-9.
pdf.

Bras:2024:CLS

[BP24] Pierre Bras and Gilles Pagès.

Convergence of Langevin-simulated annealing algorithms with multiplicative noise. Mathematics of Computation, 93(348): 1761–1803, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2024-03899-1.

Brent:2021:AES

[BPT21] Richard P. Brent, David J. Platt, and Timothy S. Trudgian. Accurate estimation of sums over zeros of the Riemann zeta-function. Mathematics of Computation, 90 (332):2923–2935, November 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03652-2;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03652-
2/S0025-5718-2021-03652-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Brent%2C%20Richard%
20P.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1045993;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=909247.

Bartels:2021:SSS

[BR21] Sören Bartels and Philipp Reiter. Stability of a simple scheme for the approximation of elastic knots and self-avoiding inextensible curves. Mathematics of Computation, 90(330): 1499–1526, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-330/
S0025-5718-2021-03633-9;
https://www.ams.org/journals/
mcom/2021-90-330/S0025-5718-
2021-03633-9/S0025-5718-2021-
03633-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Bartels%
2C%20Soren;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=854824.

Bou-Rabee:2025:RRK

[BRK25] Nawaf Bou-Rabee and Tore Selland Kleppe. Randomized Runge– Kutta–Nyström methods for unadjusted Hamiltonian and kinetic Langevin Monte Carlo. Mathematics of Computation, 94(356): 2839–2865, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2025-04061-4.

Balasubramanian:2023:EBP

[BRS23] Ramachandran Balasubramanian, Olivier Ramaré, and Priyamvad Srivastav. Explicit bounds for products of primes in AP. Mathematics of Computation, 92(343):2381– 2411, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03853-4.

Bostan:2024:MDE

[BRS24] Alin Bostan, Tanguy Rivoal, and Bruno Salvy. Minimization of differential equations and algebraic values of E-functions. Mathematics of Computation, 93(347): 1427–1472, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03912-6.

Barucq:2023:CAH

[BRT23] Hélène Barucq, Nathan Rouxelin, and Sébastien Tordeux. Construction and analysis of a HDG solution for the totalflux formulation of the convected Helmholtz equation. Mathematics of Computation, 92(343):2097– 2131, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03850-9.

Barbulescu:2022:CEF

[BS22] Razvan Barbulescu and Sudarshan Shinde. A classification of ECM-friendly families of elliptic curves using modular curves. Mathematics of Computation, 91 (335):1405–1436, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03697-2;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03697-2/S0025-5718-2021-
03697-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Shinde%
2C%20Sudarshan;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=1003662.

Becher:2023:CPG

[BS23] Verónica Becher and Gabriel Sac Himelfarb. A construction of a λ-Poisson generic sequence. Mathematics of Computation, 92 (341):1453–1466, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2023-03806-
6/S0025-5718-2023-03806-6.
pdf;   https://www.ams.org/

mathscinet/search/authors.
html?authorName=Sac%20Himelfarb%
2C%20Gabriel;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=368040.

Belabas:2024:PDN

[BS24] Karim Belabas and Denis Simon. Power detection over number fields. Mathematics of Computation, 93(348): 1953–1961, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2024-03913-3.

Berthe:2021:SLE

[BST21] Valérie Berthé, Wolfgang Steiner, and Jörg M. Thuswaldner. On the second Lyapunov exponent of some multidimensional continued fraction algorithms. Mathematics of Computation, 90(328): 883–905, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03592-3;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03592-
3/S0025-5718-2020-03592-3.
pdf.

Burr:2023:NHK

[BSW23] M. Burr, F. Sottile, and E. Walker. Numerical homotopies from Khovanskii bases. Mathematics of Computation, 92(343): 2333–2353, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              343/S0025-5718-2023-03689-4.

Balakrishnan:2020:ECI

[BT20] Jennifer S. Balakrishnan and Jan Tuitman. Explicit Coleman integration for curves. Mathematics of Computation, 89(326): 2965–2984, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-326/S0025-5718-2020-
                                              03542-X;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              326/S0025-5718-2020-03542-
                                              X/S0025-5718-2020-03542-X.
                                              pdf.

Bender:2022:TEM

[BT22] Matías R. Bender and Simon Telen. Toric eigenvalue methods for solving sparse polynomial systems. Mathematics of Computation, 91(337):2397– 2429, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-337/
                                              S0025-5718-2022-03744-3;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-337/S0025-5718-
                                              2022-03744-3/S0025-5718-2022-
                                              03744-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Bender%
                                              2C%20Matias%20R.;   https:/
                                              /www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=1277543.

Banas:2023:RPE

[BV23] Lubomír Banas and Christian Vieth. Robust a posteriori estimates for the stochastic Cahn– Hilliard equation. Mathematics of Computation, 92(343):2025– 2063, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03836-4.

Brandhorst:2024:HLA

[BV24] Simon Brandhorst and Davide Cesare Veniani. Hensel lifting algorithms for quadratic forms. Mathematics of Computation, 93(348):1963–1991, July 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2024-
93-348/S0025-5718-2023-03909-
6.

Berg:2020:OOO

[BVA20] Jennifer Berg and Anthony Várilly-Alvarado. Odd order obstructions to the Hasse principle on general K3 surfaces. Mathematics of Computation, 89(323): 1395–1416, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03485-3;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03485-
3/S0025-5718-2019-03485-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Varilly-Alvarado%
2C%20Anthony;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1061301.

Bao:2024:EET

[BW24] Weizhu Bao and Chushan Wang. Error estimates of the timesplitting methods for the nonlinear Schrödinger equation with semi-smooth nonlinearity. Mathematics of Computation, 93(348): 1599–1631, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              348/S0025-5718-2023-03900-X.

Bruno:2024:MSF

[BY24] Oscar P. Bruno and Tao Yin. Multiple-scattering frequencytime hybrid solver for the wave equation in interior domains. Mathematics of Computation, 93 (346):551–587, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              346/S0025-5718-2023-03872-8.

Bao:2024:OTE

[BZ24] Gang Bao and Yixuan Zhang. Optimal transportation for electrical impedance tomography. Mathematics of Computation, 93(349):2361–2389, September 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

                                              ams.org/journals/mcom/2024-
                                              93-349/S0025-5718-2023-03919-
                                              9.

Carstensen:2020:CMA

[Car20]  Carsten Carstensen.  Collec-

tive marking for adaptive leastsquares finite element methods with optimal rates. Mathematics of Computation, 89(321): 89–103, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03474-9;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03474-
9/S0025-5718-2019-03474-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Carstensen%
2C%20Carsten.

Chen:2025:ACP

[CC25] Li-An Chen and Robert S. Coulter. Algorithms for computing the permutation resemblance of functions over finite groups. Mathematics of Computation, 94 (353):1543–1570, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03996-0.

Calegari:2020:SMA

[CCG20] Frank Calegari, Shiva Chidambaram, and Alexandru Ghitza. Some modular abelian surfaces. Mathematics of Computation, 89(321):387–394, January 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03434-8;   https://www.ams.
org/journals/mcom/2020-89-

                                              321/S0025-5718-2019-03434-
                                              8/S0025-5718-2019-03434-8.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Chidambaram%
                                              2C%20Shiva;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=678536;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              713726.

Cox:2023:SSL

[CD23] David A. Cox and Carlos D’Andrea. Subresultants and the Shape Lemma. Mathematics of Computation, 92(343):2355– 2379, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              343/S0025-5718-2023-03840-6.

Cangiani:2022:VDG

[CDG22] Andrea Cangiani, Zhaonan Dong, and Emmanuil H. Georgoulis. hpversion discontinuous Galerkin methods on essentially arbitrarilyshaped elements. Mathematics of Computation, 91(333):1– 35, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-333/
                                              S0025-5718-2021-03667-4;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-333/S0025-5718-
                                              2021-03667-4/S0025-5718-2021-
                                              03667-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1163337;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              750860;   https://www.ams.

org/mathscinet/search/authors.
html?mrauthid=757643.

Connelly:2025:AGC

[CDLT25] Erin Connelly, Timothy Duff, and Jessie Loucks-Tavitas. Algebra and geometry of camera resectioning. Mathematics of Computation, 94(355):2613– 2643, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
355/S0025-5718-2024-04030-9.

Cances:2020:GPB

[CDM^{+}20] Eric Cancès, Geneviève Dusson, Yvon Maday, Benjamin Stamm, and Martin Vohralík. Guaranteed a posteriori bounds for eigenvalues and eigenvectors: Multiplicities and clusters. Mathematics of Computation, 89(326): 2563–2611, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2020-89-326/
S0025-5718-2020-03549-2;
https://www.ams.org/journals/
mcom/2020-89-326/S0025-5718-
2020-03549-2/S0025-5718-2020-
03549-2.pdf.

Cui:2022:TDW

[CDZ22] Jianbo Cui, Luca Dieci, and Haomin Zhou. Time discretizations of Wasserstein–Hamiltonian flows. Mathematics of Computation, 91(335):1019–1075, May 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2022-
91-335/S0025-5718-2022-03726-

1;   https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2022-03726-1/S0025-5718-2022-
03726-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Cui%
2C%20Jianbo;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=247339;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
643741.

Curry:2020:MEP

[CEFO20] Charles Curry, Kurusch Ebrahimi-Fard, and Brynjulf Owren. The Magnus expansion and post-Lie algebras. Mathematics of Computation, 89(326):2785– 2799, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03541-8;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03541-
8/S0025-5718-2020-03541-8.
pdf.

Craig:2023:BMI

[CEHT23] Katy Craig, Karthik Elamvazhuthi, Matt Haberland, and Olga Turanova. A blob method for inhomogeneous diffusion with applications to multi-agent control and sampling. Mathematics of Computation, 92(344):2575– 2654, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03841-8.

Celledoni:2020:EPM

[CEOR20] Elena Celledoni, Sølve Eidnes, Brynjulf Owren, and Torbjørn Ringholm. Energypreserving methods on Riemannian manifolds. Mathematics of Computation, 89(322): 699–716, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03470-1;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03470-
1/S0025-5718-2019-03470-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Eidnes%2C%
20Solve;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Owren%2C%20Brynjulf;
https://www.ams.org/mathscinet/
search/authors.html?authorName=
Ringholm%2C%20Torbjorn;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=623033.

Chaumont-Frelet:2023:SEP

[CF23] T. Chaumont-Frelet. A simple equilibration procedure leading to polynomial-degree-robust a posteriori error estimators for the curl–curl problem. Mathematics of Computation, 92(344): 2413–2437, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03817-0.

Chaumont-Frelet:2022:SBP

[CFEV22] T. Chaumont-Frelet, A. Ern, and M. Vohralík. Stable broken H(curl) polynomial extensions and p-robust a posteriori error estimates by broken patchwise equilibration for the curl–curl problem. Mathematics of Computation, 91(333):37– 74, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03673-X;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03673-X/S0025-5718-2021-
03673-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Vohralik%
2C%20M.;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=349433;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=999028.

Carrillo:2021:SON

[CFS21] José A. Carrillo, Ulrik S. Fjordholm, and Susanne Solem. A second-order numerical method for the aggregation equations. Mathematics of Computation, 90(327):103–139, January 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03563-7;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03563-
7/S0025-5718-2020-03563-7.
pdf.

Chen:2025:GPC

[CGT25] Tyler Chen, Anne Greenbaum, and Thomas Trogdon. GMRES, pseudospectra, and Crouzeix’s conjecture for shifted and scaled Ginibre matrices. Mathematics of Computation, 94(351): 241–261, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03963-7.

Chen:2020:NVE

[CH20] Long Chen and Xuehai Huang.

Nonconforming virtual element method for 2m th order partial differential equations in ℝ^{n}. Mathematics of Computation, 89 (324):1711–1744, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2019-
03498-1;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2019-03498-
1/S0025-5718-2019-03498-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=735779;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=854280.

Chen:2022:FEE

[CH22a] Long Chen and Xuehai Huang.

A finite element elasticity complex in three dimensions. Mathematics of Computation, 91(337): 2095–2127, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03739-X;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03739-X/S0025-5718-2022-
03739-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=735779;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
854280.

Chen:2022:FED

[CH22b] Long Chen and Xuehai Huang. Finite elements for div div conforming symmetric tensors in three dimensions. Mathematics of Computation, 91(335): 1107–1142, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03700-X;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03700-X/S0025-5718-2021-
03700-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=735779;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
854280.

Carle:2024:EAS

[CH24a] Constantin Carle and Marlis Hochbruck. Error analysis of second-order local time integration methods for discontinuous Galerkin discretizations of linear wave equations. Mathematics of Computation, 93(350): 2611–2641, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03952-2.

Chen:2024:FER

[CH24b] Long Chen and Xuehai Huang.

Finite element de Rham and Stokes complexes in three dimensions. Mathematics of Computation, 93(345):55–110, January 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2024-
93-345/S0025-5718-2023-03859-
5.

Carstensen:2025:NNC

[CH25a] Carsten Carstensen and Norbert Heuer. Normal-normal continuous symmetric stresses in mixed finite element elasticity. Mathematics of Computation, 94(354): 1571–1602, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
354/S0025-5718-2024-04017-6.

Chen:2025:NDD

[CH25b] Long Chen and Xuehai Huang.

A new div–div-conforming symmetric tensor finite element space with applications to the biharmonic equation. Mathematics of Computation, 94(351): 33–72, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03957-1.

Chainais-Hillairet:2024:MNF

[CHEF24] C. Chainais-Hillairet, R. Eymard, and J. Fuhrmann. A monotone numerical flux for quasilinear convection diffusion equation. Mathematics of Computation, 93(345): 203–231, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
345/S0025-5718-2023-03870-4.

Cully-Hugill:2022:EIE

[CHL22] Michaela Cully-Hugill and Ethan S. Lee. Explicit interval estimates for prime numbers. Mathematics of Computation, 91(336): 1955–1970, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03719-4;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03719-4/S0025-5718-2022-
03719-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1432725;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=

1460044. See corrigendum [CHL24].

Cully-Hugill:2024:CEI

[CHL24] Michaela Cully-Hugill and Ethan S. Lee. Corrigendum to ‘Explicit interval estimates for prime numbers’. Mathematics of Computation, 93(346): 1019–1025, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

org/journals/mcom/2024-93-
346/S0025-5718-2023-03884-4.

Cui:2022:DFN

[CHS22] Jianbo Cui, Jialin Hong, and Derui Sheng. Density function of numerical solution of splitting AVF scheme for stochastic Langevin equation. Mathematics of Computation, 91(337): 2283–2333, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03752-2;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03752-2/S0025-5718-2022-
03752-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1159032;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1491841;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=258322.

MC-50080007

[Chu25]  Moody T. Chu.  Preparing

Hamiltonians for quantum simulation: a computational framework for Cartan decomposition via Lax dynamics. Mathematics of Computation, 94(356):2867– 2893, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2025-04056-0.

Cances:2024:MOM

[CHV24] Eric Cancès, Muhammad Hassan, and Laurent Vidal. Modifiedoperator method for the calculation of band diagrams of crystalline materials. Mathematics of Computation, 93(347): 1203–1245, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03897-2.

Chen:2020:CAF

[CHW20] Long Chen, Xiaozhe Hu, and Steven M. Wise. Convergence analysis of the Fast Subspace Descent method for convex optimization problems. Mathematics of Computation, 89(325):2249– 2282, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03526-1;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03526-
1/S0025-5718-2020-03526-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=615795;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=735779;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
793307.

Cho:2025:EFE

[CIK^{+}25] Sungmun Cho, Tamotsu Ikeda, Hidenori Katsurada, Chul hee Lee, and Takuya Yamauchi. An explicit formula for the extended Gross–Keating datum of a quadratic form. Mathematics of Computation, 94(353): 1503–1541, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2025-04050-X.

Chirre:2021:PAP

[CJdL21] Andrés Chirre, Valdir José Pereira Júnior, and David de Laat. Primes in arithmetic progressions and semidefinite programming. Mathematics of Computation, 90(331):2235–2246, September 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

/www.ams.org/AMSMathViewer;
 https://www.ams.org/journals/
mcom/2021-90-331/S0025-5718-
2021-03638-8;   https://www.
ams.org/journals/mcom/2021-
90-331/S0025-5718-2021-03638-
8/S0025-5718-2021-03638-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Chirre%2C%
20Andres;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Pereira%20Junior%
2C%20Valdir%20Jose;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1067611.

Cheng:2023:SLD

[CJS23] Yao Cheng, Shan Jiang, and Martin Stynes. Supercloseness of the local discontinuous Galerkin method for a singularly perturbed convection– diffusion problem. Mathematics of Computation, 92(343):2065– 2095, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

                                              org/journals/mcom/2023-92-
                                              343/S0025-5718-2023-03844-3.

Corso:2026:NSG

[CKMS26] Thiago Carvalho Corso, Gaspard Kemlin, Christof Melcher, and Benjamin Stamm. Numerical simulation of the Gross–Pitaevskii equation via vortex tracking. Mathematics of Computation, 95(357):227–262, January 2026. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2026-95-
                                              357/S0025-5718-2025-04071-7.

Cools:2021:FCC

[CKNS21] Ronald Cools, Frances Y. Kuo, Dirk Nuyens, and Ian H. Sloan. Fast component-by-component construction of lattice algorithms for multivariate approximation with POD and SPOD weights. Mathematics of Computation, 90 (328):787–812, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-328/S0025-5718-2020-
                                              03586-8;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              328/S0025-5718-2020-03586-
                                              8/S0025-5718-2020-03586-8.
                                              pdf.

Cheval:2025:LRA

[CL25] Horatiu Cheval and Laurentiu Leustean. Linear rates of asymptotic regularity for Halperntype iterations. Mathematics of Computation, 94(353): 1323–1333, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03991-1.

Caruso:2026:ACN

[CL26] Xavier Caruso and Antoine Leudière. Algorithms for computing norms and characteristic polynomials on general Drinfeld modules. Mathematics of Computation, 95(357): 415–455, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04052-3.

Chen:2025:FDS

[CLLZ25] Ziang Chen, Jianfeng Lu, Yulong Lu, and Xiangxiong Zhang. Fully discretized Sobolev gradient flow for the Gross–Pitaevskii eigenvalue problem. Mathematics of Computation, 94(356):2723– 2760, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2024-04032-2.

Chartier:2022:UAN

[CLT22] Philippe Chartier, Mohammed Lemou, and Léopold Trémant. A uniformly accurate numerical method for a class of dissipative systems. Mathematics of Computation, 91(334): 843–869, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-334/

S0025-5718-2021-03688-1;
https://www.ams.org/journals/
mcom/2022-91-334/S0025-5718-
2021-03688-1/S0025-5718-2021-
03688-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Tremant%
2C%20Leopold;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=335517;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
355223.

Cai:2023:NAI

[CLY23] Zhenning Cai, Jianfeng Lu, and Siyao Yang. Numerical analysis for inchworm Monte Carlo method: Sign problem and error growth. Mathematics of Computation, 92(341): 1141–1209, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2022-03785-
6/S0025-5718-2022-03785-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1264228;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=822782;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
892483.

Cohen:2023:NOA

[CM23] Albert Cohen and Giovanni Migliorati. Near-optimal approximation methods for elliptic PDEs with lognormal coefficients. Mathematics of Computation, 92 (342):1665–1691, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03825-X.

Carneiro:2025:FOM

[CMR25] Emanuel Carneiro, Micah B.

Milinovich, and Antonio Pedro Ramos. Fourier optimization and Montgomery’s pair correlation conjecture. Mathematics of Computation, 94(351): 409–424, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03990-X.

Coppersmith:2024:ISP

[CMSV24] Don Coppersmith, Michael J.

Mossinghoff, Danny Scheinerman, and Jeffrey M. VanderKam. Ideal solutions in the Prouhet–Tarry– Escott problem. Mathematics of Computation, 93(349):2473– 2501, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
349/S0025-5718-2023-03917-5.

Cances:2021:CPE

[CNV21] Clément Cancès, Flore Nabet, and Martin Vohralík. Convergence and a posteriori error analysis for energy-stable finite element approximations of degenerate parabolic equations. Mathematics of Computation, 90(328): 517–563, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-328/
S0025-5718-2020-03577-7;

                                              https://www.ams.org/journals/
                                              mcom/2021-90-328/S0025-5718-
                                              2020-03577-7/S0025-5718-2020-
                                              03577-7.pdf.

Coeuret:2024:TSF

[Coe24] Lucas Coeuret. Tamed stability of finite difference schemes for the transport equation on the half-line. Mathematics of Computation, 93(347):1097– 1151, May 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2024-93-347/
                                              S0025-5718-2023-03901-1.

Colbrook:2020:POP

[Col20] Matthew J. Colbrook. Pseudoergodic operators and periodic boundary conditions. Mathematics of Computation, 89(322): 737–766, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-322/S0025-5718-2019-
                                              03475-0;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              322/S0025-5718-2019-03475-
                                              0/S0025-5718-2019-03475-0.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1251387.

Chen:2021:CEB

[COS21] Yifan Chen, Houman Owhadi, and Andrew M. Stuart. Consistency of empirical Bayes and kernel flow for hierarchical parameter estimation. Mathematics of Computation, 90(332):2527– 2578, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03649-2;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03649-
2/S0025-5718-2021-03649-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1291574;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=168395;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
695100.

Chen:2025:SCF

[COS25] Yifan Chen, Houman Owhadi, and Florian Schäfer. Sparse Cholesky factorization for solving nonlinear PDEs via Gaussian processes. Mathematics of Computation, 94(353): 1235–1280, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03992-3.

Camano:2026:CFE

[COSS26] Jessika Camaño, Ricardo Oyarzúa, Miguel Serón, and Manuel Solano. A conforming finite element method for a nonisothermal fluid– membrane interaction. Mathematics of Computation, 95(357): 29–72, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04045-6.

Creutz:2020:GJE

[Cre20] Brendan Creutz. Generalized Jacobians and explicit descents. Mathematics of Computation, 89 (323):1365–1394, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-323/S0025-5718-2019-
                                              03491-9;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              323/S0025-5718-2019-03491-
                                              9/S0025-5718-2019-03491-9.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=949383.

Calvo:2022:TID

[CRS22] María Cabrera Calvo, Frédéric Rousset, and Katharina Schratz. Time integrators for dispersive equations in the long wave regime. Mathematics of Computation, 91(337):2197– 2214, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-337/
                                              S0025-5718-2022-03745-5;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-337/S0025-5718-
                                              2022-03745-5/S0025-5718-2022-
                                              03745-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Cabrera%
                                              20Calvo%2C%20Maria;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?authorName=Rousset%
                                              2C%20Frederic;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=990639.

Carstensen:2022:CRT

[CS22] Carsten Carstensen and Stefan A. Sauter. Crouzeix– Raviart triangular elements are inf-sup stable. Mathematics of Computation, 91(337):2041– 2057, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03742-X;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03742-X/S0025-5718-2022-
03742-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=263782;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
313335.

Casulli:2023:CGM

[CS23] Angelo A. Casulli and Igor Simunec. Computation of generalized matrix functions with rational Krylov methods. Mathematics of Computation, 92(340): 749–777, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03788-
1/S0025-5718-2022-03788-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1399012;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1446320.

Carrillo:2025:CSR

[CS25] José A. Carrillo and Jakub Skrzeczkowski. Convergence and stability results for the particle system in the Stein gradient descent method. Mathematics of Computation, 94(354): 1793–1814, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
354/S0025-5718-2025-04000-6.

Cushing:2024:PAS

[CSS24] David Cushing, George W. Stagg, and David I. Stewart. A Prolog assisted search for new simple Lie algebras. Mathematics of Computation, 93(347): 1473–1495, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03889-3.

Chen:2023:LOP

[CSX23] Wenbin Chen, Nian Shao, and Xuejun Xu. A locally optimal preconditioned Newton– Schur method for symmetric elliptic eigenvalue problems. Mathematics of Computation, 92(344): 2655–2684, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03860-1.

Chassagneux:2020:CMS

[CT20] Jean-Francois Chassagneux and Camilo A. Garcia Trillos. Cubature method to solve BS-DEs: Error expansion and complexity control. Mathematics of Computation, 89(324): 1895–1932, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03522-4;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03522-
4/S0025-5718-2020-03522-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1096828;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=810835.

Chada:2022:CAE

[CT22a] Neil K. Chada and Xin T.

Tong. Convergence acceleration of ensemble Kalman inversion in nonlinear settings. Mathematics of Computation, 91(335): 1247–1280, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03709-6;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03709-6/S0025-5718-2021-
03709-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Tong%
2C%20Xin%20T.;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1268846.

Chen:2022:DFP

[CT22b] Swaine L. Chen and Nico M.

Temme. A distribution function from population genetics statistics using Stirling numbers of the first kind: Asymptotics, inversion and numerical evaluation. Mathematics of Computation, 91 (334):871–885, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03711-4;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03711-4/S0025-5718-2021-
                                              03711-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Chen%
                                              2C%20Swaine%20L.;   https:/
                                              /www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=171490.

Cohn:2022:DLP

[CT22c] Henry Cohn and Nicholas Triantafillou. Dual linear programming bounds for sphere packing via modular forms. Mathematics of Computation, 91(333): 491–508, January 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-333/
                                              S0025-5718-2021-03662-5;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-333/S0025-5718-
                                              2021-03662-5/S0025-5718-2021-
                                              03662-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=606578;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              978350.

Chen:2023:SNM

[CW23] Qiong Chen and Qiang Wu. Salem numbers with minimal trace. Mathematics of Computation, 92(342):1779–1790, July 2023. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

                                              ams.org/journals/mcom/2023-

92-342/S0025-5718-2023-03833-
9.

Cheng:2026:OBN

[CWS26] Yao Cheng, Xuesong Wang, and Martin Stynes. Optimal balanced-norm error estimate of the LDG method for reaction–diffusion problems II: The two-dimensional case with layer-upwind flux. Mathematics of Computation, 95(357): 73–103, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04048-1.

Chen:2024:OEE

[CX24] Yuan Chen and Yulong Xing. Optimal error estimates of ultraweak discontinuous Galerkin methods with generalized numerical fluxes for multi-dimensional convection-diffusion and biharmonic equations. Mathematics of Computation, 93(349):2135– 2183, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
349/S0025-5718-2023-03927-8.

Cui:2025:MLP

[CX25] Tengteng Cui and Chuanju Xu. Müntz Legendre polynomials: Approximation properties and applications. Mathematics of Computation, 94(353): 1377–1410, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

                                              org/journals/mcom/2025-94-
                                              353/S0025-5718-2024-03987-X.

Chen:2021:NMB

[CYB^{+}21] Jingmin Chen, Thomas Yu, Patrick Brogan, Robert Kusner, Yilin Yang, and Andrew Zigerelli. Numerical methods for biomembranes: Conforming subdivision methods versus nonconforming PL methods. Mathematics of Computation, 90(328): 471–516, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2021-90-328/
                                              S0025-5718-2020-03584-4;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-328/S0025-5718-
                                              2020-03584-4/S0025-5718-2020-
                                              03584-4.pdf.

Chung:2025:SFH

[CZ25] Eric Chung and Lina Zhao. The stabilization-free HDG method for fluid-structure interaction in a unified mixed formulation on Alfeld splits. Mathematics of Computation, 94(354): 1633–1665, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04009-7.

deCourcy-Ireland:2024:SDS

[dCIDV24] Matthew de Courcy-Ireland, Maria Dostert, and Maryna Viazovska. Six-dimensional sphere packing and linear programming. Mathematics of Computation, 93 (348):1993–2029, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2024-03959-5.

DiPietro:2023:FDP

[DD23] Daniele A. Di Pietro and Jérôme Droniou. A fully discrete plates complex on polygonal meshes with application to the Kirchhoff–Love problem. Mathematics of Computation, 92(339): 51–77, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03765-
0/S0025-5718-2022-03765-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=655312;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=790640.

DelCorso:2020:CDI

[DDDR20] Gianna M. Del Corso, Ilaria Del Corso, Roberto Dvornicich, and Francesco Romani. On computing the density of integers of the form 2^{n} + p. Mathematics of Computation, 89(325): 2365–2386, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03537-6;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03537-
6/S0025-5718-2020-03537-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=150045;   https:

                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=313164;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              600466;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=61055.

DeTeran:2020:BEC

[De 20] Fernando De Terán. Backward error and conditioning of Fiedler companion linearizations. Mathematics of Computation, 89(323): 1259–1300, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-323/S0025-5718-2019-
                                              03480-4;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              323/S0025-5718-2019-03480-
                                              4/S0025-5718-2019-03480-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=De%20Teran%
                                              2C%20Fernando.

Dus:2025:NSP

[DE25] Mathias Dus and Virginie Ehrlacher. Numerical solution of Poisson partial differential equation in high dimension using two-layer neural networks. Mathematics of Computation, 94(351): 159–208, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-03971-6.

DiRocco:2022:SHB

[DEG22] Sandra Di Rocco, David Eklund, and Oliver Gäfvert. Sampling and homology via bottlenecks. Mathematics of Computation, 91(338): 2969–2995, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03757-
1/S0025-5718-2022-03757-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Gafvert%2C%
20Oliver;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=606949;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=906385.

Dembele:2022:EAS

[Dem22] Lassina Dembélé. On the existence of abelian surfaces with everywhere good reduction. Mathematics of Computation, 91(335): 1381–1403, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03692-3;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03692-3/S0025-5718-2021-
03692-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Dembele%
2C%20Lassina.

Despres:2022:TMR

[DES22] Bruno Després, Maria El Ghaoui, and Toni Sayah. A Trefftz method with reconstruction of the normal derivative applied to elliptic equations. Mathematics of Computation, 91(338): 2645–2679, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2022-91-
                                              338/S0025-5718-2022-03756-
                                              X/S0025-5718-2022-03756-X.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Despres%2C%
                                              20Bruno;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?authorName=El%20Ghaoui%
                                              2C%20Maria;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=704612.

Daniels:2020:TRE

[DGJ20] Harris B. Daniels and Enrique González-Jiménez. On the torsion of rational elliptic curves over sextic fields. Mathematics of Computation, 89(321): 411–435, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-321/S0025-5718-2019-
                                              03440-3;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              321/S0025-5718-2019-03440-
                                              3/S0025-5718-2019-03440-3.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1105200;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=703386.

Dahmen:2020:ANS

[DGM20] Wolfgang Dahmen, Felix Gruber, and Olga Mula. An adaptive nested source term iteration for radiative transfer equations. Mathematics of Computation, 89 (324):1605–1646, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03505-4;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03505-
4/S0025-5718-2020-03505-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Gruber%2C%
20Felix;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=1027794;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=54100.

Dick:2022:CCC

[DGS22] Josef Dick, Takashi Goda, and Kosuke Suzuki. Componentby-component construction of randomized rank-1 lattice rules achieving almost the optimal randomized error rate. Mathematics of Computation, 91(338):2771– 2801, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03769-
8/S0025-5718-2022-03769-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1057946;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=735535;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
936012.

DiRocco:2026:CED

[DGS26] Sandra Di Rocco, Lukas Gustafsson, and Luca Sodomaco. Conditional Euclidean distance optimization via relative tangency. Mathematics of Computation, 95(357):477–524, January 2026. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2026-95-
                                              357/S0025-5718-2025-04047-X.

Drake:2022:CAS

[DGSW22] Dow Drake, Jay Gopalakrishnan, Joachim Schöberl, and Christoph Wintersteiger. Convergence analysis of some tent-based schemes for linear hyperbolic systems. Mathematics of Computation, 91 (334):699–733, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03686-8;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03686-8/S0025-5718-2021-
                                              03686-8.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Schoberl%
                                              2C%20Joachim;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1237024;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1241583;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=661361.

Diao:2020:GMS

[DH20] Chenzhe Diao and Bin Han. Generalized matrix spectral factorization and quasi-tight framelets with a minimum number of generators. Mathematics of Computation, 89(326):2867–2911, November 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

/www.ams.org/AMSMathViewer;
 https://www.ams.org/journals/
mcom/2020-89-326/S0025-5718-
2020-03523-6;   https://www.
ams.org/journals/mcom/2020-
89-326/S0025-5718-2020-03523-
6/S0025-5718-2020-03523-6.
pdf.

Dina:2023:IHN

[DIS23] Bogdan Dina, Sorina Ionica, and Jeroen Sijsling. Isogenous hyperelliptic and non-hyperelliptic Jacobians with maximal complex multiplication. Mathematics of Computation, 92(339): 349–383, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03776-
5/S0025-5718-2022-03776-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1476109;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=891818;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
974789.

Dostert:2021:KNN

[DK21] Maria Dostert and Alexander Kolpakov. Kissing number in non-Euclidean spaces of constant sectional curvature. Mathematics of Computation, 90(331):2507– 2525, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-

                                              03622-4;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              331/S0025-5718-2021-03622-
                                              4/S0025-5718-2021-03622-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1165370;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=774696.

Delbourgo:2023:IIA

[DK23] Daniel Delbourgo and Heiko Knospe. On Iwasawa λ-invariants for abelian number fields and random matrix heuristics. Mathematics of Computation, 92(342): 1817–1836, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03823-6.

Dajani:2025:UTM

[DKS25] Karma Dajani, Cor Kraaikamp, and Slade Sanderson. A unifying theory for metrical results on regular continued fraction convergents and mediants. Mathematics of Computation, 94(356): 3101–3144, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              356/S0025-5718-2025-04046-8.

Dorich:2022:FDE

[DL22] Benjamin Dörich and Jan Leibold. Full discretization error analysis of exponential integrators for semilinear wave equations. Mathematics of Computation, 91(336):1687–1709, July 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2022-
91-336/S0025-5718-2022-03736-
4;   https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03736-4/S0025-5718-2022-
03736-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Dorich%
2C%20Benjamin;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1385394.

Dong:2021:HAB

[DLL21] Heping Dong, Jun Lai, and Peijun Li. A highly accurate boundary integral method for the elastic obstacle scattering problem. Mathematics of Computation, 90 (332):2785–2814, November 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03660-1;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03660-
1/S0025-5718-2021-03660-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Dong%2C%20Heping;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Lai%2C%20Jun;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=682916.

DeRyck:2024:EAD

[DM24] T. De Ryck and S. Mishra. Error analysis for deep neural network approximations of parametric hyperbolic conservation laws. Mathematics of Computation, 93 (350):2643–2677, November 2024.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              350/S0025-5718-2023-03934-5.

Delecroix:2020:ALM

[DMM20] Vincent Delecroix, Carlos Matheus, and Carlos Gustavo Moreira. Approximations of the Lagrange and Markov spectra. Mathematics of Computation, 89(325):2521– 2536, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-325/S0025-5718-2020-
                                              03513-3;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              325/S0025-5718-2020-03513-
                                              3/S0025-5718-2020-03513-3.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1026380;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=366894;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              730493.

Dong:2024:CEB

[DMM24] Ruiwen Dong, Stephen Melczer, and Marc Mezzarobba. Computing error bounds for asymptotic expansions of regular P - recursive sequences. Mathematics of Computation, 93(346): 977–1017, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              346/S0025-5718-2023-03888-1.

Dietzel:2023:EFA

[DMV23] C. Dietzel, P. Menchón, and L. Vendramin. On the enumeration of finite L-algebras. Mathematics of Computation, 92(341): 1363–1381, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2023-03814-
5/S0025-5718-2023-03814-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Menchon%2C%
20P.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1305522;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=829575.

Doerr:2022:SDB

[Doe22] Benjamin Doerr. A sharp discrepancy bound for jittered sampling. Mathematics of Computation, 91 (336):1871–1892, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03727-3;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03727-3/S0025-5718-2022-
03727-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=667218.

Dvorakova:2023:ART

[DOP23] L’ubomíra Dvoráková, Daniela Opocenská, and Edita Pelantová. Asymptotic repetitive threshold of balanced sequences. Mathematics of Computation, 92(341): 1403–1429, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2023-03816-
9/S0025-5718-2023-03816-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Dvorakova%
2C%20L%27ubomira;   https:/
/www.ams.org/mathscinet/search/
authors.html?authorName=Opocenska%
2C%20Daniela;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Pelantova%
2C%20Edita.

Duran:2020:SSP

[DOS20] Ricardo G. Durán, Enrique Otárola, and Abner J. Salgado. Stability of the Stokes projection on weighted spaces and applications. Mathematics of Computation, 89(324): 1581–1603, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03509-1;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03509-
1/S0025-5718-2020-03509-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Duran%2C%20Ricardo%
20G.;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Otarola%2C%
20Enrique;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=847180.

Dummigan:2024:QFP

[DPRT24] N. Dummigan, A. Pacetti, G. Rama, and G. Tornaría. Quinary forms and paramodular forms. Mathematics of Computation, 93(348):1805–1858, July 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2024-
93-348/S0025-5718-2024-03815-
2.

Du:2020:NAT

[DS20a] Shukai Du and Francisco-Javier Sayas. New analytical tools for HDG in elasticity, with applications to elastodynamics. Mathematics of Computation, 89(324): 1745–1782, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2019-
03499-3;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2019-03499-
3/S0025-5718-2019-03499-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1303285;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=621885.

Duarte:2020:ITP

[DS20b] Eliana Duarte and Alexandra Seceleanu. Implicitization of tensor product surfaces via virtual projective resolutions. Mathematics of Computation, 89(326): 3023–3056, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03548-0;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03548-
0/S0025-5718-2020-03548-0.
pdf.

Du:2021:NDH

[DS21] Shukai Du and Francisco-Javier Sayas. A note on devising HDG+ projections on polyhedral elements. Mathematics of Computation, 90(327): 65–79, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03573-X;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03573-
X/S0025-5718-2020-03573-X.
pdf.

diSerafino:2023:LLS

[dSKJV23] Daniela di Serafino, Natasa Krejić, Natasa Krklec Jerinkić, and Marco Viola. LSOS: Linesearch second-order stochastic optimization methods for nonconvex finite sums. Mathematics of Computation, 92(341): 1273–1299, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2022-03802-
3/S0025-5718-2022-03802-3.
pdf;   https://www.ams.org/

mathscinet/search/authors.
html?authorName=Krejic%2C%
20Natasa;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Krklec%20Jerinkic%
2C%20Natasa;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1197355;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
811010.

Dusson:2023:AFS

[DSS23] Geneviève Dusson, Israel Michael Sigal, and Benjamin Stamm. Analysis of the Feshbach–Schur method for the Fourier spectral discretizations of Schrödinger operators. Mathematics of Computation, 92(339):217–249, January 2023. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2023-
92-339/S0025-5718-2022-03774-
1/S0025-5718-2022-03774-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Dusson%2C%
20Genevieve;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=161895;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
824171.

Diening:2023:ION

[DST23] Lars Diening, Johannes Storn, and Tabea Tscherpel. Interpolation operator on negative Sobolev spaces. Mathematics of Computation, 92(342): 1511–1541, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03824-8.

Dahmen:2022:ACD

[DSW22] Wolfgang Dahmen, Rob Stevenson, and Jan Westerdiep. Accuracy controlled data assimilation for parabolic problems. Mathematics of Computation, 91 (334):557–595, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03680-7;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03680-7/S0025-5718-2021-
                                              03680-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1356589;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              310898;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=54100.

Duan:2021:FSM

[Dua21]  Lian Duan.  Faltings–Serre

method on three dimensional selfdual representations. Mathematics of Computation, 90(328): 931–951, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-328/S0025-5718-2020-
                                              03591-1;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              328/S0025-5718-2020-03591-
                                              1/S0025-5718-2020-03591-1.
                                              pdf.

Ehler:2024:AAM

[EG24] Martin Ehler and Karlheinz Gröchenig. An abstract approach to Marcinkiewicz–Zygmund inequalities for approximation and quadrature in modulation spaces. Mathematics of Computation, 93 (350):2885–2919, November 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2023-03930-8.

Eymard:2025:OEB

[EGH25] Robert Eymard, Thierry Gallouët, and Raphaèle Herbin. Optimal error bounds for the twopoint flux approximation finite volume scheme. Mathematics of Computation, 94(355):2271– 2298, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
355/S0025-5718-2024-04033-4.

Esedoglu:2024:MFM

[EGL24] Selim Esedoğlu, Jiajia Guo, and David Li. On median filters for motion by mean curvature. Mathematics of Computation, 93(350): 2679–2710, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03940-6.

Elliott:2020:SOS

[EH20] Charles M. Elliott and Philip J.

Herbert. Second order splitting of a class of fourth order PDEs with point constraints. Mathematics of Computation, 89(326): 2613–2648, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03556-X;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03556-
X/S0025-5718-2020-03556-X.
pdf.

Ernvall-Hytonen:2022:EBN

[EHP22] Anne-Maria Ernvall-Hytönen and Neea Palojärvi. Explicit bound for the number of primes in arithmetic progressions assuming the Generalized Riemann Hypothesis. Mathematics of Computation, 91 (335):1317–1365, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2022-03691-7;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2022-03691-7/S0025-5718-2022-
03691-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Ernvall-
Hytonen%2C%20Anne-Maria;
https://www.ams.org/mathscinet/
search/authors.html?authorName=
Palojarvi%2C%20Neea.

Elder:2024:HPA

[ELM^{+}24] Jennifer Elder, Nadia Lafrenière, Erin McNicholas, Jessica Striker, and Amanda Welch. Homomesies on permutations: an analysis of maps and statistics in the FindStat database. Mathematics of Computation, 93(346): 921–976, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03866-2.

Eddargani:2026:QRS

[EMS26] Salah Eddargani, Carla Manni, and Hendrik Speleers. Quadrature rules for splines of high smoothness on uniformly refined triangles. Mathematics of Computation, 95(357): 321–338, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04058-4.

Elvetun:2024:IST

[EN24] Ole Løseth Elvetun and Bjørn Fredrik Nielsen. Identifying the source term in the potential equation with weighted sparsity regularization. Mathematics of Computation, 93(350):2811– 2836, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03941-8.

Ergur:2021:SAC

[EPR21] Alperen A. Ergür, Grigoris Paouris, and J. Maurice Rojas. Smoothed analysis for the condition number of structured real polynomial systems. Mathematics of Computation, 90(331): 2161–2184, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03647-9;   https://www.ams.
org/journals/mcom/2021-90-
331/S0025-5718-2021-03647-
9/S0025-5718-2021-03647-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Ergur%2C%20Alperen%
20A.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=354192;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=671202.

Ettahri:2021:FMP

[ERS21] S. Ettahri, O. Ramaré, and L. Surel. Fast multi-precision computation of some Euler products. Mathematics of Computation, 90(331):2247–2265, September 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-331/
S0025-5718-2021-03630-3;
https://www.ams.org/journals/
mcom/2021-90-331/S0025-5718-
2021-03630-3/S0025-5718-2021-
03630-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Ettahri%
2C%20S.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Ramare%2C%
20O.;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Surel%2C%20L.

.

Elsenhans:2024:MH

[ES24] Andreas-Stephan Elsenhans and Michael Stoll. Minimization of hypersurfaces. Mathematics of Computation, 93(349):2513– 2555, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
349/S0025-5718-2024-03924-8.

Ern:2020:SBP

[EV20] Alexandre Ern and Martin Vohralík. Stable broken H^{1} and H(div) polynomial extensions for polynomial-degree-robust potential and flux reconstruction in three space dimensions. Mathematics of Computation, 89(322): 551–594, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03482-8;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03482-
8/S0025-5718-2019-03482-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Vohralik%2C%
20Martin;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=349433.

Feischl:2022:ISS

[Fei22] Michael Feischl. Inf-sup stability implies quasi-orthogonality. Mathematics of Computation, 91(337):2059–2094, September 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2022-
91-337/S0025-5718-2022-03748-

                                              0;   https://www.ams.org/journals/
                                              mcom/2022-91-337/S0025-5718-
                                              2022-03748-0/S0025-5718-2022-
                                              03748-0.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=965785.

Fu:2020:ESP

[FGN20] Guosheng Fu, Johnny Guzmán, and Michael Neilan. Exact smooth piecewise polynomial sequences on Alfeld splits. Mathematics of Computation, 89(323): 1059–1091, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-323/S0025-5718-2020-
                                              03520-0;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              323/S0025-5718-2020-03520-
                                              0/S0025-5718-2020-03520-0.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1061680;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=775211;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              824091.

Fuhrer:2025:MFE

[FH25] Thomas Führer and Norbert Heuer. Mixed finite elements for Kirchhoff–Love plate bending. Mathematics of Computation, 94 (353):1065–1099, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              353/S0025-5718-2024-03995-9.

Friedman:2021:MFA

[FJRR21] Eduardo Friedman, Fredrik Johansson, and Gabriel Ramirez-Raposo. The minimal Fried average entropy for higher-rank Cartan actions. Mathematics of Computation, 90(328): 973–978, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03582-0;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03582-
0/S0025-5718-2020-03582-0.
pdf.

Fruhbis-Kruger:2021:EDA

[FKRS21] Anne Frühbis-Krüger, Lukas Ristau, and Bernd Schober. Embedded desingularization for arithmetic surfaces — toward a parallel implementation. Mathematics of Computation, 90(330): 1957–1997, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-330/S0025-5718-2021-
03624-8;   https://www.ams.
org/journals/mcom/2021-90-
330/S0025-5718-2021-03624-
8/S0025-5718-2021-03624-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Fruhbis-Kruger%
2C%20Anne;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Ristau%2C%
20Lukas;   https://www.ams.

                                              org/mathscinet/search/authors.
                                              html?mrauthid=1218416.

Feng:2024:LPS

[FKRT24] Jinchao Feng, Charles Kulick, Yunxiang Ren, and Sui Tang. Learning particle swarming models from data with Gaussian processes. Mathematics of Computation, 93(349):2391– 2437, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              349/S0025-5718-2023-03915-1.

Flammang:2020:ASS

[Fla20]  V. Flammang.  An ana-

log to the Schur–Siegel–Smyth trace problem. Mathematics of Computation, 89(325):2387– 2398, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-325/S0025-5718-2020-
                                              03518-2;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              325/S0025-5718-2020-03518-
                                              2/S0025-5718-2020-03518-2.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=360354.

Flammang:2021:UBU

[Fla21] V. Flammang. Upper bounds for the usual measures of totally positive algebraic integers with house less than 5.8. Mathematics of Computation, 90(327): 379–388, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03580-7;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03580-
7/S0025-5718-2020-03580-7.
pdf.

Feireisl:2022:AVS

[FLMSS22] Eduard Feireisl, Mária Lukácová-Medvid’ová, Simon Schneider, and Bangwei She. Approximating viscosity solutions of the Euler system. Mathematics of Computation, 91(337):2129– 2164, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03738-8;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03738-8/S0025-5718-2022-
03738-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Lukacova-
Medvid%27ova%2C%20Maria;
https://www.ams.org/mathscinet/
search/authors.html?authorName=
Schneider%2C%20Simon;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1165111;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
65780.

Ferretti:2020:SSL

[FM20] Roberto Ferretti and Michel Mehrenberger. Stability of Semi-Lagrangian schemes of arbitrary odd degree under constant and variable advection speed. Mathematics of Computation, 89(324):

1783–1805, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-324/S0025-5718-2019-
                                              03494-4;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              324/S0025-5718-2019-03494-
                                              4/S0025-5718-2019-03494-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=272089;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=736186.

Faustmann:2021:QOC

[FMP21] Markus Faustmann, Jens Markus Melenk, and Dirk Praetorius. Quasi-optimal convergence rate for an adaptive method for the integral fractional Laplacian. Mathematics of Computation, 90(330): 1557–1587, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2021-90-330/
                                              S0025-5718-2021-03603-0;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-330/S0025-5718-
                                              2021-03603-0/S0025-5718-2021-
                                              03603-0.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1123286;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              613978;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=702616.

Feng:2024:LTE

[FMS24] Yue Feng, Georg Maierhofer, and Katharina Schratz. Longtime error bounds of lowregularity integrators for nonlinear Schrödinger equations. Mathematics of Computation, 93(348): 1569–1598, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2023-03922-9.

Franco:2023:DLA

[FMZ23] Nicola R. Franco, Andrea Manzoni, and Paolo Zunino. A deep learning approach to Reduced Order Modelling of parameter dependent partial differential equations. Mathematics of Computation, 92(340): 483–524, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03781-
9/S0025-5718-2022-03781-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Franco%2C%
20Nicola%20R.;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=690612;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
948813.

Farrell:2022:FEA

[FOS22] Patrick Farrell, Pablo Alexei Gazca Orozco, and Endre Süli. Finite element approximation and preconditioning for anisothermal flow of implicitly-constituted non-Newtonian fluids. Mathematics of Computation, 91(334): 659–697, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03703-5;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03703-5/S0025-5718-2021-
                                              03703-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Gazca%
                                              20Orozco%2C%20Pablo%20Alexei;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              168795;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=873997.

Fiordilino:2020:DHI

[FP20] Joseph A. Fiordilino and Ali Pakzad. A discrete Hopf interpolant and stability of the finite element method for natural convection. Mathematics of Computation, 89(322): 629–643, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-322/S0025-5718-2019-
                                              03489-0;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              322/S0025-5718-2019-03489-
                                              0/S0025-5718-2019-03489-0.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1072925;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=1236501.

Faverjon:2022:ARR

[FP22] Colin Faverjon and Marina Poulet. An algorithm to recognize regular singular Mahler systems. Mathematics of Computation, 91 (338):2905–2928, November 2022.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03758-
3/S0025-5718-2022-03758-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Poulet%2C%
20Marina;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=969708.

Friesecke:2025:CPG

[FP25] Gero Friesecke and Maximilian Penka. Convergence proof for the GenCol algorithm in the case of two-marginal optimal transport. Mathematics of Computation, 94(351):263–275, January 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03968-6.

Frei:2022:CAE

[FR22] Christopher Frei and Rodolphe Richard. Constructing abelian extensions with prescribed norms. Mathematics of Computation, 91(333):381–399, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
333/S0025-5718-2021-03663-
7;   https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03663-7/S0025-5718-2021-
03663-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=875874;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
938397.

Frengley:2023:CEC

[Fre23]  Sam Frengley.  Congruences

of elliptic curves arising from nonsurjective mod N Galois representations. Mathematics of Computation, 92(339): 409–450, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              339/S0025-5718-2022-03770-
                                              4/S0025-5718-2022-03770-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Frengley%2C%
                                              20Sam.

Frommer:2025:ASP

[FRS25] Andreas Frommer, Michele Rinelli, and Marcel Schweitzer. Analysis of stochastic probing methods for estimating the trace of functions of sparse symmetric matrices. Mathematics of Computation, 94 (352):801–823, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              352/S0025-5718-2024-03984-4.

Fiori:2020:ALC

[FS20] Andrew Fiori and Andrew Shallue. Average liar count for degree-2 Frobenius pseudoprimes. Mathematics of Computation, 89(321): 493–514, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-321/S0025-5718-2019-
                                              03452-X;   https://www.ams.

org/journals/mcom/2020-89-
321/S0025-5718-2019-03452-
X/S0025-5718-2019-03452-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=805175;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=997971.

Filaseta:2021:PBC

[FS21] Michael Filaseta and Jeremiah Southwick. Primes that become composite after changing an arbitrary digit. Mathematics of Computation, 90(328): 979–993, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03593-5;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03593-
5/S0025-5718-2020-03593-5.
pdf.

Feliu:2022:KRF

[FS22] Elisenda Feliu and AmirHosein Sadeghimanesh. Kac–Rice formulas and the number of solutions of parametrized systems of polynomial equations. Mathematics of Computation, 91(338): 2739–2769, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03760-
1/S0025-5718-2022-03760-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1311158;   https:

                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=865998.

Fu:2024:DQP

[FS24] Hang Fu and Michael Stoll. Dynamics of quadratic polynomials and rational points on a curve of genus 4. Mathematics of Computation, 93(345): 397–410, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03883-2.

Fu:2025:ABC

[FSTW25] Fengyu Fu, Chi-Wang Shu, Qi Tao, and Boying Wu. Analysis of the boundary conditions for the ultraweak-local discontinuous Galerkin method of time-dependent linear fourthorder problems. Mathematics of Computation, 94(351): 123–158, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-03955-8.

Fu:2024:EDI

[FTY24] Zhaohui Fu, Tao Tang, and Jiang Yang. Energy diminishing implicit–explicit Runge– Kutta methods for gradient flows. Mathematics of Computation, 93 (350):2745–2767, November 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              350/S0025-5718-2024-03950-9.

Fuhrer:2022:MDN

[Füh22] Thomas Führer. Multilevel decompositions and norms for negative order Sobolev spaces. Mathematics of Computation, 91(333): 183–218, January 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03674-1;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03674-1/S0025-5718-2021-
03674-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1017746.

Friedland:2020:SNS

[FW20] Shmuel Friedland and Li Wang.

Spectral norm of a symmetric tensor and its computation. Mathematics of Computation, 89(325): 2175–2215, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03525-X;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03525-
X/S0025-5718-2020-03525-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Wang%2C%20Li;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
69405.

Falk:2023:CPP

[FW23] Richard S. Falk and Ragnar Winther. Construction of polynomial preserving cochain extensions by blending. Mathematics of Computation, 92(342): 1575–1594, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03819-4.

Giesl:2021:MDI

[GAHW21] Peter Giesl, Carlos Argáez, Sigurdur Hafstein, and Holger Wendland. Minimization with differential inequality constraints applied to complete Lyapunov functions. Mathematics of Computation, 90(331):2137– 2160, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03629-7;   https://www.ams.
org/journals/mcom/2021-90-
331/S0025-5718-2021-03629-
7/S0025-5718-2021-03629-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Argaez%2C%
20Carlos;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=602098;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=697182;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
725648.

Gallistl:2023:MML

[Gal23] Dietmar Gallistl. Mixed methods and lower eigenvalue bounds. Mathematics of Computation, 92 (342):1491–1509, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03820-0.

Gu:2023:IDA

[GCL23] Huipeng Gu, Mingchao Cai, and Jingzhi Li. An iterative decoupled algorithm with unconditional stability for Biot model. Mathematics of Computation, 92 (341):1087–1108, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2023-03809-
1/S0025-5718-2023-03809-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Gu%2C%20Huipeng;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Li%2C%20Jingzhi;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=724403.

Gutleb:2022:CEM

[GCO22] Timon S. Gutleb, José A. Carrillo, and Sheehan Olver. Computing equilibrium measures with power law kernels. Mathematics of Computation, 91(337):2247– 2281, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03740-6;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03740-6/S0025-5718-2022-
03740-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Carrillo%
2C%20Jose%20A.;   https://

                                              www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=1384319;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              783322.

Goncalves:2025:MNZ

[GdLL25] Felipe Gonçalves, David de Laat, and Nando Leijenhorst. Multiplicity of nontrivial zeros of primitive L-functions via higherlevel correlations. Mathematics of Computation, 94(354): 2041–2058, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04005-X.

Guglielmi:2026:CNO

[GDST26] Nicola Guglielmi, Arturo De Marinis, Anton Savostianov, and Francesco Tudisco. Contractivity of neural ODEs: an eigenvalue optimization problem. Mathematics of Computation, 95(357): 293–319, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2026-95-
                                              357/S0025-5718-2025-04059-6.

Grasegger:2024:CCM

[GEL24] Georg Grasegger, Boulos El Hilany, and Niels Lubbes. Coupler curves of moving graphs and counting realizations of rigid graphs. Mathematics of Computation, 93(345):459– 504, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.

org/journals/mcom/2024-93-
345/S0025-5718-2023-03886-8.

Gu:2024:RAS

[GG24] Ran Gu and Bing Gao. A random active set method for strictly convex quadratic problem with simple bounds. Mathematics of Computation, 93(350): 2837–2860, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03982-0.

Gess:2023:TCP

[GGKO23] Benjamin Gess, Rishabh S.

Gvalani, Florian Kunick, and Felix Otto. Thermodynamically consistent and positivitypreserving discretization of the thin-film equation with thermal noise. Mathematics of Computation, 92(343):1931–1976, September 2023. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2023-92-343/
S0025-5718-2023-03830-3.

Gopalakrishnan:2020:SDE

[GGO20] Jay Gopalakrishnan, Luka Grubišić, and Jeffrey Ovall. Spectral discretization errors in filtered subspace iteration. Mathematics of Computation, 89(321): 203–228, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03483-X;   https://www.ams.

org/journals/mcom/2020-89-
321/S0025-5718-2019-03483-
X/S0025-5718-2019-03483-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Grubisic%2C%
20Luka;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=661361;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=728623.

Gong:2023:CRA

[GGS23] Shihua Gong, Ivan G. Graham, and Euan A. Spence. Convergence of restricted additive Schwarz with impedance transmission conditions for discretised Helmholtz problems. Mathematics of Computation, 92(339): 175–215, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03772-
8/S0025-5718-2022-03772-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1240731;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=76020;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
858295.

Garde:2022:SRC

[GH22] Henrik Garde and Nuutti Hyvönen. Series reversion in Calderón’s problem. Mathematics of Computation, 91(336):1925–1953, July 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2022-
91-336/S0025-5718-2022-03729-

7;   https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03729-7/S0025-5718-2022-
03729-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Hyvonen%
2C%20Nuutti;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1077217.

Gortz:2024:ISS

[GHM24] Morgan Görtz, Fredrik Hellman, and Axel Målqvist. Iterative solution of spatial network models by subspace decomposition. Mathematics of Computation, 93(345): 233–258, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
345/S0025-5718-2023-03861-3.

Gantner:2021:ROA

[GHPS21] Gregor Gantner, Alexander Haberl, Dirk Praetorius, and Stefan Schimanko. Rate optimality of adaptive finite element methods with respect to overall computational costs. Mathematics of Computation, 90(331): 2011–2040, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03654-6;   https://www.ams.
org/journals/mcom/2021-90-
331/S0025-5718-2021-03654-
6/S0025-5718-2021-03654-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1107233;   https:
//www.ams.org/mathscinet/search/

                                              authors.html?mrauthid=1147415;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1311878;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=702616.

Griebel:2023:LRA

[GHS23] Michael Griebel, Helmut Harbrecht, and Reinhold Schneider. Low-rank approximation of continuous functions in Sobolev spaces with dominating mixed smoothness. Mathematics of Computation, 92(342): 1729–1746, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03813-3.

Gao:2023:EFM

[GI23] Jing Gao and Arieh Iserles. On an extended Filon method for highly oscillatory integrals over a simplex. Mathematics of Computation, 92(340): 867–893, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              340/S0025-5718-2022-03797-
                                              2/S0025-5718-2022-03797-2.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Gao%2C%20Jing;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              91725.

Goedgebeur:2024:FHC

[GJL^{+}24] Jan Goedgebeur, Jorik Jooken, On-Hei Solomon Lo, Ben Seamone, and Carol T. Zamfirescu.

Few Hamiltonian cycles in graphs with one or two vertex degrees. Mathematics of Computation, 93 (350):3059–3082, November 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03943-1.

Gonzalez–Jimenez:2020:GTG

[GJN20] Enrique González-Jiménez and Filip Najman. Growth of torsion groups of elliptic curves upon base change. Mathematics of Computation, 89(323): 1457–1485, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03478-6;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03478-
6/S0025-5718-2019-03478-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=703386;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=886852.

Giesselmann:2026:TSS

[GK26] Jan Giesselmann and Sam G.

Krupa. Theory of shifts, shocks, and the intimate connections to L^{2}-type A posteriori error analysis of numerical schemes for hyperbolic problems. Mathematics of Computation, 95(357): 147–225, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

                                              org/journals/mcom/2026-95-
                                              357/S0025-5718-2025-04043-2.

Granger:2021:CBB

[GKL^{+}21] Robert Granger, Thorsten Kleinjung, Arjen K. Lenstra, Benjamin Wesolowski, and Jens Zumbrägel. Computation of a 30750-bit binary field discrete logarithm. Mathematics of Computation, 90(332):2997– 3022, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2021-90-332/
                                              S0025-5718-2021-03669-8;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-332/S0025-5718-
                                              2021-03669-8/S0025-5718-2021-
                                              03669-8.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=112545;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1163085;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=704259;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=744248;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              843678.

Grekas:2025:CDG

[GKMV25] G. Grekas, K. Koumatos, C. Makridakis, and A. Vikelis. A class of discontinuous Galerkin methods for nonlinear variational problems. Mathematics of Computation, 94(355):2221– 2250, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.

org/journals/mcom/2025-94-
355/S0025-5718-2025-04040-7.

Grzeskowiak:2025:SC

[GKPR25] Maciej Grześkowiak, Jerzy Kaczorowski, Lukasz Pańkowski, and Maciej Radziejewski. On the sign changes of ψ(x) − x. Mathematics of Computation, 94(356): 3083–3100, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2025-04053-5.

Gilbert:2022:EBS

[GKS22] Alexander D. Gilbert, Frances Y.

Kuo, and Ian H. Sloan. Equivalence between Sobolev spaces of first-order dominating mixed smoothness and unanchored ANOVA spaces on ℝ^{d}. Mathematics of Computation, 91(336): 1837–1869, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03718-2;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03718-2/S0025-5718-2022-
03718-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1225284;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
163675;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=703418.

Goda:2024:RTR

[GKS24] Takashi Goda, Yoshihito Kazashi, and Yuya Suzuki. Randomizing the trapezoidal rule gives the optimal RMSE rate in Gaussian Sobolev spaces. Mathematics of Computation, 93(348): 1655–1676, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2023-03910-2.

Greenwood:2024:DCM

[GKW24] Torin Greenwood, Jonathan Kariv, and Noah Williams. From discrete to continuous: Monochromatic 3-term arithmetic progressions. Mathematics of Computation, 93(350):2959– 2983, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03970-4.

Guglielmi:2020:NIL

[GLFN20] Nicola Guglielmi, María López-Fernández, and Giancarlo Nino. Numerical inverse Laplace transform for convection-diffusion equations. Mathematics of Computation, 89(323):1161–1191, May 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

/www.ams.org/AMSMathViewer;
 https://www.ams.org/journals/
mcom/2020-89-323/S0025-5718-
2020-03497-8;   https://www.
ams.org/journals/mcom/2020-
89-323/S0025-5718-2020-03497-
8/S0025-5718-2020-03497-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Lopez-Fernandez%
2C%20Maria;   https://www.
ams.org/mathscinet/search/

authors.html?authorName=Nino%
2C%20Giancarlo;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=603494.

Guzman:2022:ESW

[GLN22] Johnny Guzmán, Anna Lischke, and Michael Neilan. Exact sequences on Worsey–Farin splits. Mathematics of Computation, 91 (338):2571–2608, November 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03746-
7/S0025-5718-2022-03746-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1214513;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=775211;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
824091.

Gamanda:2020:STB

[GLNY20] Maroua Gamanda, Henri Lombardi, Stefan Neuwirth, and Ihsen Yengui. The syzygy theorem for Bézout rings. Mathematics of Computation, 89(322): 941–964, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03466-X;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03466-
X/S0025-5718-2019-03466-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1235530;   https:

                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=277650;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              635157;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=657905.

Gkanis:2021:NCE

[GM21] Ioannis Gkanis and Charalambos G. Makridakis. A new class of entropy stable schemes for hyperbolic systems: Finite element methods. Mathematics of Computation, 90(330): 1663–1699, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2021-90-330/
                                              S0025-5718-2021-03617-0;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-330/S0025-5718-
                                              2021-03617-0/S0025-5718-2021-
                                              03617-0.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Gkanis%
                                              2C%20Ioannis;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=289627.

Gautschi:2024:RID

[GM24] Walter Gautschi and Gradimir V. Milovanović. A Ramanujan integral and its derivatives: computation and analysis. Mathematics of Computation, 93(347): 1297–1308, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              347/S0025-5718-2023-03892-3.

Gajovic:2025:CAH

[GM25a] Stevan Gajović and J. Steffen Müller. Computing p-adic heights on hyperelliptic curves. Mathematics of Computation, 94(354): 2059–2088, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
354/S0025-5718-2024-04028-0.

Grenie:2025:BBB

[GM25b] Loïc Grenié and Giuseppe Molteni.

Breaking the 4 barrier for the bound of a generating set of the class group. Mathematics of Computation, 94(356):3145– 3163, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2025-04114-0.

Grote:2021:SLB

[GMS21] Marcus J. Grote, Simon Michel, and Stefan A. Sauter. Stabilized leapfrog based local time-stepping method for the wave equation. Mathematics of Computation, 90 (332):2603–2643, November 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03650-9;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03650-
9/S0025-5718-2021-03650-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Michel%2C%
20Simon;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=313335;   https:

                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=360720.

Goedgebeur:2020:GFH

[GMZ20] Jan Goedgebeur, Barbara Meersman, and Carol T. Zamfirescu. Graphs with few Hamiltonian cycles. Mathematics of Computation, 89(322): 965–991, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-322/S0025-5718-2019-
                                              03465-8;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              322/S0025-5718-2019-03465-
                                              8/S0025-5718-2019-03465-8.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Meersman%2C%
                                              20Barbara;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=756214;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=945364.

Gawlik:2025:FEA

[GN25] Evan S. Gawlik and Michael Neunteufel. Finite element approximation of scalar curvature in arbitrary dimension. Mathematics of Computation, 94(356): 2685–2722, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              356/S0025-5718-2024-04038-3.

Gallouet:2024:GEV

[GNT24] Thomas O. Gallouët, Andrea Natale, and Gabriele Todeschi. From geodesic extrapolation to a variational BDF2 scheme for Wasserstein gradient flows. Mathematics of Computation, 93(350): 2769–2810, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03951-0.

Griebel:2020:SSC

[GO20] Michael Griebel and Peter Oswald. Stochastic subspace correction methods and fault tolerance. Mathematics of Computation, 89(321):279–312, January 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03459-2;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03459-
2/S0025-5718-2019-03459-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Oswald%2C%
20Peter;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=270664.

MC-68360010

[God26] Takashi Goda. A randomized lattice rule without component-by-component construction. Mathematics of Computation, 95(357): 339–361, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04051-1.

Gosse:2020:ATD

[Gos20] Laurent Gosse. Aliasing and two-dimensional well-balanced for drift-diffusion equations on square grids. Mathematics of Computation, 89(321):139– 168, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-321/S0025-5718-2019-
                                              03451-8;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              321/S0025-5718-2019-03451-
                                              8/S0025-5718-2019-03451-8.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=611045.

Guglielmi:2024:CSG

[GP24] Nicola Guglielmi and Vladimir Yu. Protasov. Computing the spectral gap of a family of matrices. Mathematics of Computation, 93(345): 259–291, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03856-X.

Goldston:2024:ECS

[GPS24] Daniel A. Goldston, Apoorva Panidapu, and Jordan Schettler. Explicit calculations for Sono’s multidimensional sieve of E_{2}-numbers. Mathematics of Computation, 93(350):2943– 2958, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.

org/journals/mcom/2024-93-
350/S0025-5718-2024-03938-8.

Gnewuch:2021:GFI

[GPW21] Michael Gnewuch, Hendrik Pasing, and Christian Weiß. A generalized Faulhaber inequality, improved bracketing covers, and applications to discrepancy. Mathematics of Computation, 90(332): 2873–2898, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03666-2;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03666-
2/S0025-5718-2021-03666-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Weiss%2C%20Christian;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1392287;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=728124.

Graham:2020:SFE

[GS20] I. G. Graham and S. A. Sauter.

Stability and finite element error analysis for the Helmholtz equation with variable coefficients. Mathematics of Computation, 89(321):105–138, January 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03457-9;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03457-

                                              9/S0025-5718-2019-03457-9.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Sauter%2C%
                                              20S.%20A.;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=76020.

Gjerde:2022:NMN

[GS22] Ingeborg G. Gjerde and L. Ridgway Scott. Nitsche’s method for Navier–Stokes equations with slip boundary conditions. Mathematics of Computation, 91(334): 597–622, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03682-0;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03682-0/S0025-5718-2021-
                                              03682-0.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1339212;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              157720.

Gopalakrishnan:2023:SSA

[GS23] Jay Gopalakrishnan and Zheng Sun. Stability of structure-aware Taylor methods for tents. Mathematics of Computation, 92(341): 1061–1086, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              341/S0025-5718-2023-03811-
                                              X/S0025-5718-2023-03811-X.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Sun%2C%20Zheng;
                                               https://www.ams.org/mathscinet/

search/authors.html?mrauthid=
661361.

Gong:2022:FEH

[GSTZ22] Bo Gong, Jiguang Sun, Tiara Turner, and Chunxiong Zheng. Finite element/holomorphic operator function method for the transmission eigenvalue problem. Mathematics of Computation, 91 (338):2517–2537, November 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03767-
4/S0025-5718-2022-03767-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Gong%2C%20Bo;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Zheng%2C%20Chunxiong;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=738684;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
995099.

Guralnick:2023:IGA

[GSW23] Robert M. Guralnick, John Shareshian, and Russ Woodroofe. On invariable generation of alternating groups by elements of prime and prime power order. Mathematics of Computation, 92 (341):1349–1361, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2023-03807-
8/S0025-5718-2023-03807-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=618746;   https:

                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=656572;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              78455.

Greaves:2023:ELE

[GSY23] Gary R. W. Greaves, Jeven Syatriadi, and Pavlo Yatsyna. Equiangular lines in Euclidean spaces: Dimensions 17 and 18. Mathematics of Computation, 92 (342):1867–1903, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03832-7.

Gallistl:2023:CRF

[GT23] Dietmar Gallistl and Ngoc Tien Tran. Convergence of a regularized finite element discretization of the two-dimensional Monge– Ampère equation. Mathematics of Computation, 92(342): 1467–1490, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03794-2.

Guo:2024:CAP

[GWWZ24] Yunzhuo Guo, Cheng Wang, Steven M. Wise, and Zhengru Zhang. Convergence analysis of a positivity-preserving numerical scheme for the Cahn–Hilliard– Stokes system with Flory– Huggins energy potential. Mathematics of Computation, 93(349): 2185–2214, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.

org/journals/mcom/2024-93-
349/S0025-5718-2023-03916-3.

Hanot:2023:AOF

[Han23]  Marien-Lorenzo Hanot.  An

arbitrary-order fully discrete Stokes complex on general polyhedral meshes. Mathematics of Computation, 92(343):1977– 2023, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03837-6.

Harvey:2021:EOF

[Har21] David Harvey. An exponent onefifth algorithm for deterministic integer factorisation. Mathematics of Computation, 90(332): 2937–2950, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-332/
S0025-5718-2021-03658-3;
https://www.ams.org/journals/
mcom/2021-90-332/S0025-5718-
2021-03658-3/S0025-5718-2021-
03658-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=734771.

Harvey:2024:FTI

[Har24] David Harvey. Faster truncated integer multiplication. Mathematics of Computation, 93(347): 1265–1296, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2024-03939-X.

Hubert:2022:AFI

[HB22] Evelyne Hubert and Erick Rodriguez Bazan. Algorithms for fundamental invariants and equivariants of finite groups. Mathematics of Computation, 91(337):2459–2488, September 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2022-
91-337/S0025-5718-2022-03749-
2;   https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03749-2/S0025-5718-2022-
03749-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1219623;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
626362.

Helfgott:2020:ISE

[Hel20] Harald Andrés Helfgott. An improved sieve of Eratosthenes. Mathematics of Computation, 89(321):333–350, January 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03438-5;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03438-
5/S0025-5718-2019-03438-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=644718.

Huang:2026:BCP

[HG26] Qiao-Long Huang and Xiao-Shan Gao. Bit complexity of polynomial GCD on sparse representation. Mathematics of Computation, 95(357):389– 413, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04054-7.

Harris:2020:SCC

[HH20] Corey Harris and Martin Helmer.

Segrè class computation and practical applications. Mathematics of Computation, 89(321): 465–491, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03448-8;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03448-
8/S0025-5718-2019-03448-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Harris%2C%
20Corey;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=884722.

Harvey:2022:LLS

[HH22] David Harvey and Markus Hittmeir. A log-log speedup for exponent one-fifth deterministic integer factorisation. Mathematics of Computation, 91(335): 1367–1379, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03708-4;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-

                                              2021-03708-4/S0025-5718-2021-
                                              03708-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1220564;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              734771.

Hasegawa:2020:RPN

[HHY20] Sumito Hasegawa, Akinari Hoshi, and Aiichi Yamasaki. Rationality problem for norm one tori in small dimensions. Mathematics of Computation, 89(322): 923–940, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-322/S0025-5718-2019-
                                              03469-5;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              322/S0025-5718-2019-03469-
                                              5/S0025-5718-2019-03469-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Hasegawa%2C%
                                              20Sumito;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=602892;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=714371.

Han:2023:ADF

[HHZ23] Yongbin Han, Yanren Hou, and Min Zhang. Analysis of divergence-free H^{1} conforming FEM with IMEX-SAV scheme for the Navier–Stokes equations at high Reynolds number. Mathematics of Computation, 92(340): 557–582, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.

org/journals/mcom/2023-92-
340/S0025-5718-2022-03790-
X/S0025-5718-2022-03790-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Hou%2C%20Yanren;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Zhang%2C%20Min;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=1418549.

Hittmeir:2021:TST

[Hit21] Markus Hittmeir. A time-space tradeoff for Lehman’s deterministic integer factorization method. Mathematics of Computation, 90 (330):1999–2010, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-330/
S0025-5718-2021-03623-6;
https://www.ams.org/journals/
mcom/2021-90-330/S0025-5718-
2021-03623-6/S0025-5718-2021-
03623-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1220564.

Hofmann:2020:CIB

[HJ20] Tommy Hofmann and Henri Johnston. Computing isomorphisms between lattices. Mathematics of Computation, 89(326): 2931–2963, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03543-1;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03543-

                                              1/S0025-5718-2020-03543-1.
                                              pdf.

Hare:2021:NLP

[HJ21] Kevin G. Hare and Jonas Jankauskas. On Newman and Littlewood polynomials with a prescribed number of zeros inside the unit disk. Mathematics of Computation, 90(328): 831–870, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2021-90-328/
                                              S0025-5718-2020-03570-4;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-328/S0025-5718-
                                              2020-03570-4/S0025-5718-2020-
                                              03570-4.pdf.

Hwang:2024:MCI

[HJ24] WonTae Hwang and Daeyeol Jeon. Modular curves with infinitely many quartic points. Mathematics of Computation, 93(345):383–395, January 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03864-9.

Hoang:2020:HOE

[HJLW20] Thi-Thao-Phuong Hoang, Lili Ju, Wei Leng, and Zhu Wang. High order explicit local time stepping methods for hyperbolic conservation laws. Mathematics of Computation, 89(324): 1807–1842, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:

//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03507-8;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03507-
8/S0025-5718-2020-03507-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Leng%2C%20Wei;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1047473;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=645968;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=762949.

Hao:2020:ISM

[HJLZ20] Zhiwei Hao, Wenrong Jiang, Nan Li, and Lihong Zhi. On isolation of simple multiple zeros and clusters of zeros of polynomial systems. Mathematics of Computation, 89(322): 879–909, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03479-8;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03479-
8/S0025-5718-2019-03479-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Jiang%2C%20Wenrong;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1203148;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=615840;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=968435.

Hou:2023:LSO

[HJQ23] Dianming Hou, Lili Ju, and Zhonghua Qiao. A linear secondorder maximum bound principlepreserving BDF scheme for the Allen–Cahn equation with a general mobility. Mathematics of Computation, 92(344):2515– 2542, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              344/S0025-5718-2023-03843-1.

Hong:2024:DCF

[HJS24] Jialin Hong, Diancong Jin, and Derui Sheng. Density convergence of a fully discrete finite difference method for stochastic Cahn–Hilliard equation. Mathematics of Computation, 93(349): 2215–2264, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              349/S0025-5718-2023-03928-X.

Hong:2023:NPF

[HKLP23] Qingguo Hong, Johannes Kraus, Maria Lymbery, and Fadi Philo. A new practical framework for the stability analysis of perturbed saddle-point problems and applications. Mathematics of Computation, 92(340): 607–634, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              340/S0025-5718-2022-03795-
                                              9/S0025-5718-2022-03795-9.
                                              pdf;   https://www.ams.org/

mathscinet/search/authors.
html?authorName=Lymbery%2C%
20Maria;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=1007990;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1330732;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
672969.

Hirsch:2025:CEA

[HKM25] Dean Hirsch, Ido Kessler, and Uri Mendlovic. Computing π(N): an elementary approach in Õ(\sqrt{N}) time. Mathematics of Computation, 94(356):3043– 3081, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2024-04039-5.

Herrmann:2021:QMC

[HKS21] Lukas Herrmann, Magdalena Keller, and Christoph Schwab. Quasi-Monte Carlo Bayesian estimation under Besov priors in elliptic inverse problems. Mathematics of Computation, 90(330): 1831–1860, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-330/
S0025-5718-2021-03615-7;
https://www.ams.org/journals/
mcom/2021-90-330/S0025-5718-
2021-03615-7/S0025-5718-2021-
03615-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Keller%
2C%20Magdalena;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=1247312;

                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              305221.

Hoshi:2022:NOT

[HKY22] Akinari Hoshi, Kazuki Kanai, and Aiichi Yamasaki. Norm one Tori and Hasse norm principle. Mathematics of Computation, 91(337): 2431–2458, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-337/
                                              S0025-5718-2022-03735-2;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-337/S0025-5718-
                                              2022-03735-2/S0025-5718-2022-
                                              03735-2.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Kanai%
                                              2C%20Kazuki;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=602892;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              714371.

Hu:2021:CDG

[HL21] Jun Hu and Yizhou Liang. Conforming discrete Gradgradcomplexes in three dimensions. Mathematics of Computation, 90 (330):1637–1662, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-330/S0025-5718-2021-
                                              03628-5;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              330/S0025-5718-2021-03628-
                                              5/S0025-5718-2021-03628-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.

html?authorName=Liang%2C%20Yizhou;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
714525.

Hu:2025:SBM

[HLLM25] Yukuan Hu, Mengyu Li, Xin Liu, and Cheng Meng. Samplingbased methods for multi-block optimization problems over transport polytopes. Mathematics of Computation, 94(353): 1281–1322, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03989-3.

Herbin:2021:CWC

[HLMT21] R. Herbin, J.-C. Latché, S. Minjeaud, and N. Therme. Conservativity and weak consistency of a class of staggered finite volume methods for the Euler equations. Mathematics of Computation, 90 (329):1155–1177, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-329/S0025-5718-2020-
03575-3;   https://www.ams.
org/journals/mcom/2021-90-
329/S0025-5718-2020-03575-
3/S0025-5718-2020-03575-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Herbin%2C%
20R.;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Latche%2C%
20J.-C.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Minjeaud%2C%

                                              20S.;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Therme%2C%
                                              20N..

Herzog:2023:MPT

[HLR23] Roland Herzog and Estefanía Loayza-Romero. A manifold of planar triangular meshes with complete Riemannian metric. Mathematics of Computation, 92 (339):1–50, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              339/S0025-5718-2022-03775-
                                              3/S0025-5718-2022-03775-3.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Loayza-Romero%
                                              2C%20Estefania;   https://
                                              www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=726929.

Hauenstein:2021:NTM

[HLRS21] Jonathan D. Hauenstein, Anton Leykin, Jose Israel Rodriguez, and Frank Sottile. A numerical toolkit for multiprojective varieties. Mathematics of Computation, 90(327):413– 440, January 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2021-90-327/
                                              S0025-5718-2020-03566-2;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-327/S0025-5718-
                                              2020-03566-2/S0025-5718-2020-
                                              03566-2.pdf.

Herbin:2021:LMN

[HLS21] R. Herbin, J.-C. Latché, and K. Saleh. Low Mach number limit of some staggered schemes for compressible barotropic flows. Mathematics of Computation, 90 (329):1039–1087, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-329/
S0025-5718-2021-03604-2;
https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03604-2/S0025-5718-2021-
03604-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Herbin%
2C%20R.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Latche%2C%
20J.-C.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Saleh%2C%20K.

.

Hong:2022:EGA

[HLX22] Qingguo Hong, Yuwen Li, and Jinchao Xu. An extended Galerkin analysis in finite element exterior calculus. Mathematics of Computation, 91(335): 1077–1106, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03707-2;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03707-2/S0025-5718-2021-
03707-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1007990;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1129421;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=228866.

Huang:2023:CCL

[HLZ23] Juntao Huang, Ruo Li, and Yizhou Zhou. Coupling conditions for linear hyperbolic relaxation systems in two-scale problems. Mathematics of Computation, 92(343):2133–2165, September 2023. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2023-92-343/
                                              S0025-5718-2023-03845-5.

Hao:2021:SEE

[HLZZ21] Zhaopeng Hao, Huiyuan Li, Zhimin Zhang, and Zhongqiang Zhang. Sharp error estimates of a spectral Galerkin method for a diffusion-reaction equation with integral fractional Laplacian on a disk. Mathematics of Computation, 90(331):2107– 2135, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-331/S0025-5718-2021-
                                              03645-5;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              331/S0025-5718-2021-03645-
                                              5/S0025-5718-2021-03645-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Zhang%2C%20Zhimin;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1089669;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=708582;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=875642.

Hu:2022:AEE

[HM22] Jun Hu and Limin Ma. Asymptotic expansions of eigenvalues by both the Crouzeix– Raviart and enriched Crouzeix– Raviart elements. Mathematics of Computation, 91(333):75– 109, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03635-2;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03635-2/S0025-5718-2021-
03635-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1261964;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
714525.

He:2023:RTR

[HMXY23] Bingsheng He, Feng Ma, Shengjie Xu, and Xiaoming Yuan. A ranktwo relaxed parallel splitting version of the augmented Lagrangian method with step size in (0, 2) for separable convex programming. Mathematics of Computation, 92 (342):1633–1663, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03822-4.

Haberstich:2022:BOW

[HNP22] Cécile Haberstich, Anthony Nouy, and Guillaume Perrin. Boosted optimal weighted least-squares. Mathematics of Computation, 91 (335):1281–1315, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2022-03710-8;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2022-03710-8/S0025-5718-2022-
03710-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Haberstich%
2C%20Cecile;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1009710;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
702952.

Hesthaven:2021:SPR

[HP21] Jan S. Hesthaven and Cecilia Pagliantini. Structurepreserving reduced basis methods for Poisson systems. Mathematics of Computation, 90(330): 1701–1740, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-330/
S0025-5718-2021-03618-2;
https://www.ams.org/journals/
mcom/2021-90-330/S0025-5718-
2021-03618-2/S0025-5718-2021-
03618-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1150107;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
350602.

Hauck:2023:SLE

[HP23] Moritz Hauck and Daniel Peterseim. Super-localization of elliptic multiscale problems. Mathematics of Computation, 92(341): 981–1003, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2022-03798-
4/S0025-5718-2022-03798-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1416098;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=848711.

Hesthaven:2024:ASM

[HPR24] Jan S. Hesthaven, Cecilia Pagliantini, and Nicolò Ripamonti. Adaptive symplectic model order reduction of parametric particlebased Vlasov–Poisson equation. Mathematics of Computation, 93 (347):1153–1202, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03885-6.

Hure:2020:DBS

[HPW20] Côme Huré, Huyên Pham, and Xavier Warin. Deep backward schemes for high-dimensional nonlinear PDEs. Mathematics of Computation, 89(324): 1547–1579, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03514-5;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03514-
5/S0025-5718-2020-03514-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Hure%2C%20Come;
 https://www.ams.org/mathscinet/

                                              search/authors.html?authorName=
                                              Pham%2C%20Huyen;   https://
                                              www.ams.org/mathscinet/search/
                                              authors.html?authorName=Warin%
                                              2C%20Xavier.

Hangelbroek:2025:EEB

[HR25] T. Hangelbroek and C. Rieger. Extending error bounds for radial basis function interpolation to measuring the error in higher order Sobolev norms. Mathematics of Computation, 94(351): 381–407, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-03960-1.

Hare:2025:EDD

[HRR25] Warren Hare, Lindon Roberts, and Clément W. Royer. Expected decrease for derivative-free algorithms using random subspaces. Mathematics of Computation, 94(351):277–304, January 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-04011-5.

Hough:2021:SAS

[HS21a] Robert Hough and Hyojeong Son. The spectrum of the abelian sandpile model. Mathematics of Computation, 90(327): 441–469, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-327/S0025-5718-2020-
                                              03565-0;   https://www.ams.

org/journals/mcom/2021-90-
327/S0025-5718-2020-03565-
0/S0025-5718-2020-03565-0.
pdf.

Hu:2021:UAI

[HS21b] Jingwei Hu and Ruiwen Shu. On the uniform accuracy of implicitexplicit backward differentiation formulas (IMEX-BDF) for stiff hyperbolic relaxation systems and kinetic equations. Mathematics of Computation, 90(328): 641–670, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03602-3;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03602-
3/S0025-5718-2020-03602-3.
pdf.

Huang:2023:SCA

[HS23] Can Huang and Jie Shen. Stability and convergence analysis of a fully discrete semiimplicit scheme for stochastic Allen–Cahn equations with multiplicative noise. Mathematics of Computation, 92(344):2685– 2713, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03846-7.

Hoxhaj:2025:HRP

[HS25a] Eriola Hoxhaj and Josef Schicho. How to reconstruct a planar map from its branching curve.

Mathematics of Computation, 94 (352):935–952, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03988-1.

Hu:2025:UAI

[HS25b] Jingwei Hu and Ruiwen Shu. Uniform accuracy of implicitexplicit Runge–Kutta (IMEX-RK) schemes for hyperbolic systems with relaxation. Mathematics of Computation, 94(351): 209–240, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03967-4.

Hanselman:2021:GCG

[HSS21] Jeroen Hanselman, Sam Schiavone, and Jeroen Sijsling. Gluing curves of genus 1 and 2 along their 2-torsion. Mathematics of Computation, 90(331): 2333–2379, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03627-3;   https://www.ams.
org/journals/mcom/2021-90-
331/S0025-5718-2021-03627-
3/S0025-5718-2021-03627-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1112256;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1322117;
 https://www.ams.org/mathscinet/

search/authors.html?mrauthid=
974789.

Hasanalizade:2022:CZD

[HSW22] Elchin Hasanalizade, Quanli Shen, and Peng-Jie Wong. Counting zeros of Dedekind zeta functions. Mathematics of Computation, 91(333):277– 293, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03665-0;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03665-0/S0025-5718-2021-
03665-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Hasanalizade%
2C%20Elchin;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1166842;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1211484.

Heintze:2023:DE

[HTVZ23] Sebastian Heintze, Robert F.

Tichy, Ingrid Vukusic, and Volker Ziegler. On the Diophantine equation U_{n} − b^{m} = c. Mathematics of Computation, 92(344): 2825–2859, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03854-6.

Huang:2020:TMC

[Hua20] Rong Huang. A qd-type method for computing generalized singular values of BF matrix pairs with sign regularity to high relative accuracy. Mathematics of Computation, 89(321): 229–252, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03444-0;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03444-
0/S0025-5718-2019-03444-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=787036.

Hulpke:2022:PGO

[Hul22] Alexander Hulpke. The perfect groups of order up to two million. Mathematics of Computation, 91(334):1007– 1017, March 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-334/
S0025-5718-2021-03684-4;
https://www.ams.org/journals/
mcom/2022-91-334/S0025-5718-
2021-03684-4/S0025-5718-2021-
03684-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=600556.

Heid:2020:AIL

[HW20] Pascal Heid and Thomas P. Wihler. Adaptive iterative linearization Galerkin methods for nonlinear problems. Mathematics of Computation, 89(326):2707– 2734, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:

//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03545-5;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03545-
5/S0025-5718-2020-03545-5.
pdf.

He:2021:CAQ

[HW21] Zhijian He and Xiaoqun Wang.

Convergence analysis of quasi-Monte Carlo sampling for quantile and expected shortfall. Mathematics of Computation, 90(327): 303–319, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03555-8;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03555-
8/S0025-5718-2020-03555-8.
pdf.

Henning:2022:STI

[HW22] Patrick Henning and Johan Wärnegård. Superconvergence of time invariants for the Gross– Pitaevskii equation. Mathematics of Computation, 91(334): 509–555, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-334/
S0025-5718-2021-03693-5;
https://www.ams.org/journals/
mcom/2022-91-334/S0025-5718-
2021-03693-5/S0025-5718-2021-
03693-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Warnegard%

                                              2C%20Johan;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=881716.

Huang:2024:USS

[HX24] Meng Huang and Zhiqiang Xu. Uniqueness and stability for the solution of a nonlinear least squares problem. Mathematics of Computation, 93(347): 1247–1264, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              347/S0025-5718-2023-03918-7.

He:2024:CAH

[HXDM24] Songnian He, Hong-Kun Xu, Qiao-Li Dong, and Na Mei. Convergence analysis of the Halpern iteration with adaptive anchoring parameters. Mathematics of Computation, 93(345): 327–345, January 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03851-0.

Henning:2025:DGS

[HY25a] Patrick Henning and Mahima Yadav. On discrete ground states of rotating Bose–Einstein condensates. Mathematics of Computation, 94(351):1–32, January 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-03962-5.

Hoang:2025:STM

[HY25b] Thi-Thao-Phuong Hoang and Ivan Yotov. A space-time mixed finite element method for reduced fracture flow models on nonmatching grids. Mathematics of Computation, 94(356):2645– 2684, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2024-04031-0.

Imbert-Gerard:2023:STQ

[IGMS23] Lise-Marie Imbert-Gérard, Andrea Moiola, and Paul Stocker. A space-time quasi-Trefftz DG method for the wave equation with piecewise-smooth coefficients. Mathematics of Computation, 92(341):1211–1249, May 2023. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2023-
92-341/S0025-5718-2022-03786-
8/S0025-5718-2022-03786-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Imbert-Gerard%
2C%20Lise-Marie;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=1377412;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
931770.

Ilten:2023:TTC

[IL23] Nathan Ilten and Yoav Len. Tropical tangents for complete intersection curves. Mathematics of Computation, 92(340): 931–979, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03782-

0/S0025-5718-2022-03782-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1080284;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=864815.

Jiang:2021:AFO

[JCWH21] Fan Jiang, Xingju Cai, Zhongming Wu, and Deren Han. Approximate first-order primal-dual algorithms for saddle point problems. Mathematics of Computation, 90(329):1227–1262, May 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2021-
90-329/S0025-5718-2021-03610-
8;   https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03610-8/S0025-5718-2021-
03610-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Cai%
2C%20Xingju;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Han%
2C%20Deren;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Jiang%
2C%20Fan;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Wu%2C%20Zhongming.

John:2021:SNP

[JKK21] Volker John, Petr Knobloch, and Paul Korsmeier. On the solvability of the nonlinear problems in an algebraically stabilized finite element method for evolutionary transport-dominated equations. Mathematics of Computation, 90 (328):595–611, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03576-5;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03576-
5/S0025-5718-2020-03576-5.
pdf.

Ji:2025:FLT

[JLOS25] Lun Ji, Hang Li, Alexander Ostermann, and Chunmei Su. Filtered Lie–Trotter splitting for the “good” Boussinesq equation: Low regularity error estimates. Mathematics of Computation, 94(355): 2345–2365, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
355/S0025-5718-2024-04023-1.

Jayawardana:2023:SSI

[JO23] Buddhika Jayawardana and Tomoki Ohsawa. Semiexplicit symplectic integrators for non-separable Hamiltonian systems. Mathematics of Computation, 92(339): 251–281, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03778-
9/S0025-5718-2022-03778-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Jayawardana%
2C%20Buddhika;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=892043.

Jorgenson:2022:ACG

[JST22] Jay Jorgenson, Lejla Smajlović, and Holger Then. An approach for computing generators of class fields of imaginary quadratic number fields using the Schwarzian derivative. Mathematics of Computation, 91(333): 331–379, January 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03619-4;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03619-4/S0025-5718-2021-
03619-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Smajlovic%
2C%20Lejla;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=292611;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
742378.

Jarso:2022:FCP

[JT22] Tamiru Jarso and Tim Trudgian. Four consecutive primitive elements in a finite field. Mathematics of Computation, 91 (335):1521–1532, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03716-3;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03716-3/S0025-5718-2021-
03716-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1305238;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
909247.

Jin:2025:SNP

[JW25] Lingzi Jin and Xiao Wang.

Stochastic nested primal–dual method for nonconvex constrained composition optimization. Mathematics of Computation, 94(351):305–358, January 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2025-
94-351/S0025-5718-2024-03965-
0.

Jia:2024:SAS

[JZCH24] Zehui Jia, Wenxing Zhang, Xingju Cai, and Deren Han. Stochastic alternating structureadapted proximal gradient descent method with variance reduction for nonconvex nonsmooth optimization. Mathematics of Computation, 93(348): 1677–1714, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2024-03867-X.

Kaya:2022:EVI

[Kay22]  Enis Kaya.  Explicit Volo-

godsky integration for hyperelliptic curves. Mathematics of Computation, 91(337):2367– 2396, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03720-0;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03720-0/S0025-5718-2022-
03720-0.pdf;   https://www.

ams.org/mathscinet/search/
authors.html?authorName=Kaya%
2C%20Enis.

Kui:2021:CMA

[KBL21] Zhiqing Kui, Jean Baccou, and Jacques Liandrat. On the construction of multiresolution analyses associated to general subdivision schemes. Mathematics of Computation, 90(331):2185– 2208, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-331/
S0025-5718-2021-03646-7;
https://www.ams.org/journals/
mcom/2021-90-331/S0025-5718-
2021-03646-7/S0025-5718-2021-
03646-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1246334;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
328783;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=774914.

Khan:2021:RPE

[KBPS21] Arbaz Khan, Alex Bespalov, Catherine E. Powell, and David J. Silvester. Robust a posteriori error estimation for parameter-dependent linear elasticity equations. Mathematics of Computation, 90(328): 613–636, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03572-8;   https://www.ams.
org/journals/mcom/2021-90-

328/S0025-5718-2020-03572-
8/S0025-5718-2020-03572-8.
pdf.

Kirk:2023:CWS

[KÇR23] Keegan L. A. Kirk, Ayçil Çesmelioglu, and Sander Rhebergen. Convergence to weak solutions of a space-time hybridized discontinuous Galerkin method for the incompressible Navier–Stokes equations. Mathematics of Computation, 92(339): 147–174, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03780-
7/S0025-5718-2022-03780-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1324700;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=844849;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
864513.

Kreuzer:2021:CCA

[KG21] Christian Kreuzer and Emmanuil H. Georgoulis. Corrigendum to “Convergence of adaptive, discontinuous Galerkin methods”. Mathematics of Computation, 90 (328):637–640, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03611-4;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03611-

                                              4/S0025-5718-2020-03611-4.
                                              pdf.

Khomovsky:2020:USP

[Kho20] Dmitry I. Khomovsky. On using symmetric polynomials for constructing root finding methods. Mathematics of Computation, 89(325):2321–2331, September 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

                                              /www.ams.org/AMSMathViewer;
                                               https://www.ams.org/journals/
                                              mcom/2020-89-325/S0025-5718-
                                              2020-03531-5;   https://www.
                                              ams.org/journals/mcom/2020-
                                              89-325/S0025-5718-2020-03531-
                                              5/S0025-5718-2020-03531-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1067846.

Kirk:2023:AEM

[KHR23] Keegan L. A. Kirk, Tamás L. Horváth, and Sander Rhebergen. Analysis of an exactly mass conserving space-time hybridized discontinuous Galerkin method for the time-dependent Navier–Stokes equations. Mathematics of Computation, 92(340): 525–556, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              340/S0025-5718-2022-03796-
                                              0/S0025-5718-2022-03796-0.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Horvath%2C%
                                              20Tamas%20L.;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1324700;
                                               https://www.ams.org/mathscinet/

search/authors.html?mrauthid=
844849.

Kirchner:2020:NMD

[Kir20]  Kristin Kirchner.  Numerical

methods for the deterministic second moment equation of parabolic stochastic PDEs. Mathematics of Computation, 89(326):2801– 2845, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03524-8;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03524-
8/S0025-5718-2020-03524-8.
pdf.

Kashiwabara:2020:SAM

[KK20] Takahito Kashiwabara and Tomoya Kemmochi. Stability, analyticity, and maximal regularity for parabolic finite element problems on smooth domains. Mathematics of Computation, 89 (324):1647–1679, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03500-5;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03500-
5/S0025-5718-2020-03500-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1179149;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=997963.

Kluners:2021:IMN

[KK21] Jürgen Klüners and Toru Komatsu. Imaginary multiquadratic number fields with class group of exponent 3 and 5. Mathematics of Computation, 90(329): 1483–1497, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2021-90-329/
                                              S0025-5718-2021-03609-1;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-329/S0025-5718-
                                              2021-03609-1/S0025-5718-2021-
                                              03609-1.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Kluners%
                                              2C%20Jurgen;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Komatsu%
                                              2C%20Toru.

Ko:2023:SAS

[KL23] Taehee Ko and Xiantao Li. Stochastic algorithms for selfconsistent calculations of electronic structures. Mathematics of Computation, 92(342): 1693–1728, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03826-1.

Kharlampovich:2020:DPS

[KLM20] Olga Kharlampovich, Laura López, and Alexei Myasnikov. The Diophantine problem in some metabelian groups. Mathematics of Computation, 89(325):2507– 2519, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03533-9;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03533-
9/S0025-5718-2020-03533-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Lopez%2C%20Laura;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
191704;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=670299.

Kuo:2021:FIR

[KMNN21] Frances Y. Kuo, Giovanni Migliorati, Fabio Nobile, and Dirk Nuyens. Function integration, reconstruction and approximation using rank-1 lattices. Mathematics of Computation, 90(330): 1861–1897, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-330/
S0025-5718-2021-03595-4;
https://www.ams.org/journals/
mcom/2021-90-330/S0025-5718-
2021-03595-4/S0025-5718-2021-
03595-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=650310;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
703418;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=777310;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=941196.

Khot:2025:VEM

[KMRB25] Rekha Khot, David Mora, and Ricardo Ruiz-Baier. Virtual element methods for Biot– Kirchhoff poroelasticity. Mathematics of Computation, 94(353): 1101–1146, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03983-2.

Kim:2023:BSD

[KMwaabAVS23] Seoyoung Kim, M. Ram Murty, and with an appendix by Andrew V. Sutherland. From the Birch and Swinnerton-Dyer conjecture to Nagao’s conjecture. Mathematics of Computation, 92(339): 385–408, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03773-
X/S0025-5718-2022-03773-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=128555;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1343309;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
852273.

Kirschmer:2022:SIC

[KNRR22] Markus Kirschmer, Fabien Narbonne, Christophe Ritzenthaler, and Damien Robert. Spanning the isogeny class of a power of an elliptic curve. Mathematics of Computation, 91(333): 401–449, January 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/

S0025-5718-2021-03672-8;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03672-8/S0025-5718-2021-
03672-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Narbonne%
2C%20Fabien;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=702917;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
892539;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=919800.

Krieg:2022:RSF

[KNS22] David Krieg, Erich Novak, and Mathias Sonnleitner. Recovery of Sobolev functions restricted to iid sampling. Mathematics of Computation, 91(338):2715– 2738, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03763-
7/S0025-5718-2022-03763-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1225170;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1245488;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
132370.

Katthan:2021:UFS

[KNT21] Lukas Katthän, Helen Naumann, and Thorsten Theobald. A unified framework of SAGE and SONC polynomials and its duality theory. Mathematics of Computation, 90(329):1297– 1322, May 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-329/
S0025-5718-2021-03607-8;
https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03607-8/S0025-5718-2021-
03607-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Katthan%
2C%20Lukas;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Naumann%
2C%20Helen;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Theobald%
2C%20Thorsten.

Kazashi:2025:DLR

[KNZ25] Yoshihito Kazashi, Fabio Nobile, and Fabio Zoccolan. Dynamical low-rank approximation for stochastic differential equations. Mathematics of Computation, 94 (353):1335–1375, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03999-6.

Karvonen:2021:IRK

[KOG21] Toni Karvonen, Chris J. Oates, and Mark Girolami. Integration in reproducing kernel Hilbert spaces of Gaussian kernels. Mathematics of Computation, 90(331): 2209–2233, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03659-5;   https://www.ams.

org/journals/mcom/2021-90-
331/S0025-5718-2021-03659-
5/S0025-5718-2021-03659-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1256975;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=665653;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
998570.

Kopteva:2021:EAL

[Kop21] Natalia Kopteva. Error analysis of an L2-type method on graded meshes for a fractionalorder parabolic problem. Mathematics of Computation, 90(327): 19–40, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03552-2;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03552-
2/S0025-5718-2020-03552-2.
pdf.

Kritzer:2021:QMC

[KPW21] P. Kritzer, F. Pillichshammer, and G. W. Wasilkowski. On quasi-Monte Carlo methods in weighted ANOVA spaces. Mathematics of Computation, 90(329): 1381–1406, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-329/S0025-5718-2021-
03598-X;   https://www.ams.
org/journals/mcom/2021-90-

329/S0025-5718-2021-03598-
X/S0025-5718-2021-03598-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Kritzer%2C%
20P.;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Pillichshammer%
2C%20F.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Wasilkowski%
2C%20G.%20W..

Kolpakov:2020:STR

[KR20] Alexander Kolpakov and Sinai Robins. Spherical tetrahedra with rational volume, and spherical Pythagorean triples. Mathematics of Computation, 89(324): 2031–2046, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2019-
03496-8;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2019-03496-
8/S0025-5718-2019-03496-8.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=342098;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=774696.

Kaltenbacher:2022:IPN

[KR22] Barbara Kaltenbacher and William Rundell. On an inverse problem of nonlinear imaging with fractional damping. Mathematics of Computation, 91(333): 245–276, January 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/

journals/mcom/2022-91-333/
S0025-5718-2021-03683-2;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03683-2/S0025-5718-2021-
03683-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=213241;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
616341.

Kaltenbacher:2025:FRC

[KR25] Barbara Kaltenbacher and William Rundell. Fractional regularisation of the Cauchy problem for Laplace’s equation and application in some free boundary value problems. Mathematics of Computation, 94(352): 647–679, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03974-1.

Kopp:2020:RPT

[KRRZ20] Leann Kopp, Natalie Randall, J. Maurice Rojas, and Yuyu Zhu. Randomized polynomialtime root counting in prime power rings. Mathematics of Computation, 89(321):373– 385, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03431-2;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03431-
2/S0025-5718-2019-03431-2.
pdf;   https://www.ams.org/

                                              mathscinet/search/authors.
                                              html?authorName=Kopp%2C%20Leann;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?authorName=
                                              Randall%2C%20Natalie;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?authorName=Zhu%
                                              2C%20Yuyu;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=354192.

Kahle:2020:CID

[KS20] Thomas Kahle and Christian Stump. Counting inversions and descents of random elements in finite Coxeter groups. Mathematics of Computation, 89(321): 437–464, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-321/S0025-5718-2019-
                                              03443-9;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              321/S0025-5718-2019-03443-
                                              9/S0025-5718-2019-03443-9.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=869155;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=904921.

Khan:2022:NAQ

[KZ22] Arbaz Khan and Pietro Zanotti. A nonsymmetric approach and a quasi-optimal and robust discretization for the Biot’s model. Mathematics of Computation, 91 (335):1143–1170, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-335/
                                              S0025-5718-2021-03699-6;

https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03699-6/S0025-5718-2021-
03699-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1108913;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1275475.

Karp:2023:CEB

[KZ23] Dmitrii Karp and Yi Zhang. Convergent expansions and bounds for the incomplete elliptic integral of the second kind near the logarithmic singularity. Mathematics of Computation, 92(344): 2769–2794, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03874-1.

Laplagne:2020:FRE

[Lap20] Santiago Laplagne. Facial reduction for exact polynomial sum of squares decomposition. Mathematics of Computation, 89(322): 859–877, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03476-2;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03476-
2/S0025-5718-2019-03476-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=805321.

Lavenant:2021:UCD

[Lav21] Hugo Lavenant. Unconditional convergence for discretizations of dynamical optimal transport. Mathematics of Computation, 90 (328):739–786, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03567-4;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03567-
4/S0025-5718-2020-03567-4.
pdf.

Li:2022:INL

[LC22] Chao Li and Xiaojun Chen. Isotropic non-Lipschitz regularization for sparse representations of random fields on the sphere. Mathematics of Computation, 91(333):219– 243, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03655-8;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03655-8/S0025-5718-2021-
03655-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Li%
2C%20Chao;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=196364.

Liu:2020:GCP

[LD20] Xin-Wei Liu and Yu-Hong Dai. A globally convergent primal-dual interior-point relaxation method for nonlinear programs. Mathematics of Computation, 89(323): 1301–1329, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03487-7;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03487-
7/S0025-5718-2019-03487-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=620453;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=667307.

Leykin:2021:NSC

[LdCS^{+}21] Anton Leykin, Abraham Martín del Campo, Frank Sottile, Ravi Vakil, and Jan Verschelde. Numerical Schubert calculus via the Littlewood–Richardson homotopy algorithm. Mathematics of Computation, 90(329): 1407–1433, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-329/
S0025-5718-2021-03579-6;
https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03579-6/S0025-5718-2021-
03579-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Leykin%
2C%20Anton;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Martin%
20del%20Campo%2C%20Abraham;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Sottile%2C%20Frank;   https:
//www.ams.org/mathscinet/search/
authors.html?authorName=Vakil%
2C%20Ravi;   https://www.ams.

                                              org/mathscinet/search/authors.
                                              html?authorName=Verschelde%
                                              2C%20Jan.

Liu:2023:NAL

[LDHS23] Xin-Wei Liu, Yu-Hong Dai, Ya-Kui Huang, and Jie Sun. A novel augmented Lagrangian method of multipliers for optimization with general inequality constraints. Mathematics of Computation, 92 (341):1301–1330, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              341/S0025-5718-2022-03799-
                                              6/S0025-5718-2022-03799-6.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Huang%2C%20Ya-
                                              Kui;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Sun%2C%20Jie;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              620453;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=667307.

Leykekhman:2021:PEE

[Ley21] D. Leykekhman. Pointwise error estimates for C^{0} interior penalty approximation of biharmonic problems. Mathematics of Computation, 90(327): 41–63, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-327/S0025-5718-2020-
                                              03596-0;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              327/S0025-5718-2020-03596-

0/S0025-5718-2020-03596-0.
pdf.

Li:2021:QOA

[Li21]  Yuwen Li.  Quasi-optimal

adaptive mixed finite element methods for controlling natural norm errors. Mathematics of Computation, 90(328): 565–593, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03590-X;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03590-
X/S0025-5718-2020-03590-X.
pdf.

Li:2022:MNS

[Li22a]  Buyang Li.  Maximum-norm

stability of the finite element method for the Neumann problem in nonconvex polygons with locally refined mesh. Mathematics of Computation, 91(336): 1533–1585, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03724-8;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03724-8/S0025-5718-2022-
03724-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=910552.

Li:2022:WLT

[Li22b] Dong Li. Why large time-stepping methods for the Cahn–Hilliard equation is stable. Mathematics of Computation, 91(338):2501– 2515, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03768-
6/S0025-5718-2022-03768-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=723577.

MC-50080008

[Li25] Weilin Li. Multiscale estimates for the condition number of nonharmonic Fourier matrices. Mathematics of Computation, 94(356): 2895–2929, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2024-04034-6.

Li:2020:LTB

[LL20] Lei Li and Jian-Guo Liu. Large time behaviors of upwind schemes and B-schemes for Fokker–Planck equations on ℝ by jump processes. Mathematics of Computation, 89(325): 2283–2320, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03516-9;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03516-
9/S0025-5718-2020-03516-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=232398;   https:

//www.ams.org/mathscinet/search/
authors.html?mrauthid=233036.

Leykekhman:2021:WDM

[LL21] Dmitriy Leykekhman and Buyang Li. Weak discrete maximum principle of finite element methods in convex polyhedra. Mathematics of Computation, 90(327): 1–18, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03560-1;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03560-
1/S0025-5718-2020-03560-1.
pdf.

Lu:2024:LUI

[LL24] Ding Lu and Ren-Cang Li. Locally unitarily invariantizable NEPv and convergence analysis of SCF. Mathematics of Computation, 93(349):2291– 2329, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
349/S0025-5718-2024-03925-X.

Liao:2025:AED

[lLW25] Hong lin Liao and Xuping Wang. Average energy dissipation rates of explicit exponential Runge–Kutta methods for gradient flow problems. Mathematics of Computation, 94(354): 1721–1759, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

org/journals/mcom/2025-94-
354/S0025-5718-2024-04015-2.

Liao:2021:AAB

[lLZ21] Hong lin Liao and Zhimin Zhang. Analysis of adaptive BDF2 scheme for diffusion equations. Mathematics of Computation, 90 (329):1207–1226, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-329/S0025-5718-2020-
03585-6;   https://www.ams.
org/journals/mcom/2021-90-
329/S0025-5718-2020-03585-
6/S0025-5718-2020-03585-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Liao%2C%20Hong-
lin;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Zhang%2C%20Zhimin.

Li:2023:IER

[LLZ23] Dongfang Li, Xiaoxi Li, and Zhimin Zhang. Implicit-explicit relaxation Runge–Kutta methods: construction, analysis and applications to PDEs. Mathematics of Computation, 92(339): 117–146, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03766-
2/S0025-5718-2022-03766-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Li%2C%20Dongfang;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Li%2C%20Xiaoxi;   https://

www.ams.org/mathscinet/search/
authors.html?authorName=Zhang%
2C%20Zhimin.

Lu:2022:SGD

[LM22] Shuai Lu and Peter Mathé.

Stochastic gradient descent for linear inverse problems in Hilbert spaces. Mathematics of Computation, 91(336):1763–1788, July 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2022-
91-336/S0025-5718-2021-03714-
X;   https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2021-03714-X/S0025-5718-2021-
03714-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Mathe%
2C%20Peter;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=759390.

Languasco:2026:ECP

[LM26] Alessandro Languasco and Pieter Moree. Euler constants from primes in arithmetic progression. Mathematics of Computation, 95(357):363–387, January 2026. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04057-2.

Le:2020:CSS

[LMP20] Tung Le, Kay Magaard, and Alessandro Paolini. On the characters of Sylow p-subgroups of finite Chevalley groups G(p^{f}) for arbitrary primes. Mathematics of Computation, 89(323): 1501–1524, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03488-9;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03488-
9/S0025-5718-2019-03488-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Le%2C%20Tung;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1178905;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=252279.

Li:2025:OAF

[LMS25] Buyang Li, Shu Ma, and Weiwei Sun. Optimal analysis of finite element methods for the stochastic Stokes equations. Mathematics of Computation, 94(352): 551–583, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03972-8.

Li:2021:SCA

[LMY21] Xiaoyue Li, Xuerong Mao, and Hongfu Yang. Strong convergence and asymptotic stability of explicit numerical schemes for nonlinear stochastic differential equations. Mathematics of Computation, 90(332):2827– 2872, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-

03661-3;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03661-
3/S0025-5718-2021-03661-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Li%2C%20Xiaoyue;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Yang%2C%20Hongfu;   https:/
/www.ams.org/mathscinet/search/
authors.html?mrauthid=199088.

Li:2022:TSM

[LN22] Wenbo Li and Ricardo H. Nochetto. Two-scale methods for convex envelopes. Mathematics of Computation, 91(333): 111–139, January 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03521-8;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03521-8/S0025-5718-2021-
03521-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Li%
2C%20Wenbo;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=131850.

Liu:2025:SOB

[LNW25] Ruishu Liu, Andreas Neuenkirch, and Xiaojie Wang. A strong order 1.5 boundary preserving discretization scheme for scalar SDEs defined in a domain. Mathematics of Computation, 94(354): 1815–1862, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04014-0.

Lampio:2020:OGB

[LÖS20] Pekka H. J. Lampio, Patric R. J. Östergård, and Ferenc Szöllősi. Orderly generation of Butson Hadamard matrices. Mathematics of Computation, 89(321): 313–331, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-321/S0025-5718-2019-
                                              03453-1;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              321/S0025-5718-2019-03453-
                                              1/S0025-5718-2019-03453-1.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Ostergard%
                                              2C%20Patric%20R.%20J.;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?authorName=Szollosi%
                                              2C%20Ferenc;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=919188.

Looney:2025:STB

[LP25] Ryan Looney and Igor Pritsker. Schinzel-type bounds for the Mahler measure on lemniscates. Mathematics of Computation, 94 (352):919–933, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              352/S0025-5718-2024-03985-6.

Lopez:2021:UCE

[LPP21] José L. López, Pablo Palacios, and Pedro J. Pagola. Uniform convergent expansions of integral transforms. Mathematics of Computation, 90(329): 1357–1380, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-329/S0025-5718-2021-
03601-7;   https://www.ams.
org/journals/mcom/2021-90-
329/S0025-5718-2021-03601-
7/S0025-5718-2021-03601-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Lopez%2C%20Jose%
20L.;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Pagola%2C%
20Pedro%20J.;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Palacios%
2C%20Pablo.

Lairez:2024:EHP

[LPPV24] Pierre Lairez, Eric Pichon-Pharabod, and Pierre Vanhove. Effective homology and periods of complex projective hypersurfaces. Mathematics of Computation, 93 (350):2985–3025, November 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03947-9.

Li:2022:SCA

[LQT22] Dong Li, Chaoyu Quan, and Tao Tang. Stability and convergence analysis for the implicitexplicit method to the Cahn– Hilliard equation. Mathematics of Computation, 91(334): 785–809, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03704-7;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03704-7/S0025-5718-2021-
                                              03704-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1178215;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              248423;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=723577.

Li:2021:CAS

[LQW21] Xiao Li, Zhonghua Qiao, and Cheng Wang. Convergence analysis for a stabilized linear semiimplicit numerical scheme for the nonlocal Cahn–Hilliard equation. Mathematics of Computation, 90(327):171–188, January 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-327/S0025-5718-2020-
                                              03578-9;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              327/S0025-5718-2020-03578-
                                              9/S0025-5718-2020-03578-9.
                                              pdf.

Li:2024:WDM

[LQXY24] Buyang Li, Weifeng Qiu, Yupei Xie, and Wenshan Yu. Weak discrete maximum principle of isoparametric finite element methods in curvilinear polyhedra. Mathematics of Computation, 93(345):1–34, January 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
345/S0025-5718-2023-03876-5.

Languasco:2021:FAC

[LR21] Alessandro Languasco and Luca Righi. A fast algorithm to compute the Ramanujan–Deninger gamma function and some number-theoretic applications. Mathematics of Computation, 90 (332):2899–2921, November 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03668-
6;   https://www.ams.org/journals/
mcom/2021-90-332/S0025-5718-
2021-03668-6/S0025-5718-2021-
03668-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Righi%
2C%20Luca;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=354780.

Li:2023:RCS

[LRR23] Weilin Li, Kui Ren, and Donsub Rim. A range characterization of the single-quadrant ADRT. Mathematics of Computation, 92(339): 283–306, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03750-
9/S0025-5718-2022-03750-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1052146;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=711179;
https://www.ams.org/mathscinet/

search/authors.html?mrauthid=
990462.

Liu:2020:GSG

[LRY20] Hailiang Liu, James Ralston, and Peimeng Yin. General superpositions of Gaussian beams and propagation errors. Mathematics of Computation, 89(322): 675–697, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03462-2;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03462-
2/S0025-5718-2019-03462-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Liu%2C%20Hailiang;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1067782;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=144300.

LeGluher:2020:FRG

[LS20a] Aude Le Gluher and Pierre-Jean Spaenlehauer. A fast randomized geometric algorithm for computing Riemann–Roch spaces. Mathematics of Computation, 89(325): 2399–2433, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03517-0;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03517-
0/S0025-5718-2020-03517-0.

pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Le%20Gluher%
2C%20Aude;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=916294.

Linderman:2020:NIG

[LS20b] George C. Linderman and Stefan Steinerberger. Numerical integration on graphs: Where to sample and how to weigh. Mathematics of Computation, 89 (324):1933–1952, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03515-7;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03515-
7/S0025-5718-2020-03515-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1203460;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=869041.

Lange:2021:CTL

[LS21] Theresa Lange and Wilhelm Stannat. On the continuous time limit of the ensemble Kalman filter. Mathematics of Computation, 90(327):233–265, January 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03588-1;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03588-

                                              1/S0025-5718-2020-03588-1.
                                              pdf.

Lederer:2023:ENA

[LS23] Philip L. Lederer and Rolf Stenberg. Energy norm analysis of exactly symmetric mixed finite elements for linear elasticity. Mathematics of Computation, 92 (340):583–605, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              340/S0025-5718-2022-03784-
                                              4/S0025-5718-2022-03784-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1215016;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=167000.

Lederer:2024:AWS

[LS24] Philip L. Lederer and Rolf Stenberg. Analysis of weakly symmetric mixed finite elements for elasticity. Mathematics of Computation, 93(346): 523–550, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              346/S0025-5718-2023-03865-0.

Li:2025:LRE

[LS25] Hang Li and Chunmei Su. Lowregularity exponential-type integrators for the Zakharov system with rough data in all dimensions. Mathematics of Computation, 94 (352):727–762, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.

org/journals/mcom/2025-94-
352/S0025-5718-2024-03973-X.

Li:2022:NSP

[LSL22] Xiaoli Li, Jie Shen, and Zhengguang Liu. New SAV-pressure correction methods for the Navier–Stokes equations: stability and error analysis. Mathematics of Computation, 91(333): 141–167, January 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-333/
S0025-5718-2021-03651-0;
https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03651-0/S0025-5718-2021-
03651-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Liu%
2C%20Zhengguang;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=1152951;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
257933.

Laurent:2020:EAB

[LV20] Adrien Laurent and Gilles Vilmart. Exotic aromatic B-series for the study of long time integrators for a class of ergodic SDEs. Mathematics of Computation, 89(321): 169–202, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03455-5;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03455-
5/S0025-5718-2019-03455-5.

                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Laurent%2C%
                                              20Adrien;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=798890.

Lauder:2022:CAF

[LV22] Alan Lauder and Jan Vonk. Computing p-adic L-functions of totally real fields. Mathematics of Computation, 91(334): 921–942, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2022-91-334/S0025-5718-2021-
                                              03678-9;   https://www.ams.
                                              org/journals/mcom/2022-91-
                                              334/S0025-5718-2021-03678-
                                              9/S0025-5718-2021-03678-9.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=639407;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=858428.

Lehrenfeld:2025:ADP

[LvBV25] Christoph Lehrenfeld, Tim van Beeck, and Igor Voulis. Analysis of divergence-preserving unfitted finite element methods for the mixed Poisson problem. Mathematics of Computation, 94(354): 1667–1699, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04027-9.

Larsen:2021:MSM

[LwaabAL21] Michael Larsen and with an appendix by Anne Larsen. Multiplicative series, modular forms, and Mandelbrot polynomials. Mathematics of Computation, 90(327):345–377, January 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03564-9;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03564-
9/S0025-5718-2020-03564-9.
pdf.

Liu:2021:PPE

[LWW^{+}21] Chun Liu, Cheng Wang, Steven M.

Wise, Xingye Yue, and Shenggao Zhou. A positivity-preserving, energy stable and convergent numerical scheme for the Poisson– Nernst–Planck system. Mathematics of Computation, 90(331): 2071–2106, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03642-X;   https://www.ams.
org/journals/mcom/2021-90-
331/S0025-5718-2021-03642-
X/S0025-5718-2021-03642-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Yue%2C%20Xingye;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Zhou%2C%20Shenggao;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=362496;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=

                                              615795;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=652762.

Liang:2022:TLP

[LX22] Qigang Liang and Xuejun Xu. A two-level preconditioned Helmholtz– Jacobi–Davidson method for the Maxwell eigenvalue problem. Mathematics of Computation, 91 (334):623–657, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03702-3;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03702-3/S0025-5718-2021-
                                              03702-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Liang%
                                              2C%20Qigang;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=365400.

Liu:2025:ECW

[LX25] Changli Liu and Jungong Xue. Efficient computation of Wiener– Hopf factorization of Markovmodulated Brownian motion. Mathematics of Computation, 94(355):2367–2408, September 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

                                              ams.org/journals/mcom/2025-
                                              94-355/S0025-5718-2025-04055-
                                              9.

Li:2024:HOS

[LYZ24] Buyang Li, Zongze Yang, and Zhi Zhou. High-order splitting finite element methods for the subdiffusion equation with limited smoothing property. Mathematics of Computation, 93(350): 2557–2586, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
350/S0025-5718-2024-03944-3.

Li:2025:PPR

[LYZ25] Yonghai Li, Peng Yang, and Zhimin Zhang. Polynomial preserving recovery for the finite volume element methods under simplex meshes. Mathematics of Computation, 94(352): 611–645, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03980-7.

Luo:2025:UDE

[LZ25] Hao Luo and Zihang Zhang. A unified differential equation solver approach for separable convex optimization: Splitting, acceleration and nonergodic rate. Mathematics of Computation, 94(356): 3009–3041, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2025-04063-8.

Li:2020:ALD

[LZMW20] Jia Li, Dazhi Zhang, Xiong Meng, and Boying Wu. Analysis of local discontinuous Galerkin methods with generalized numerical fluxes for linearized KdV equations. Mathematics of Computation, 89(325):2085– 2111, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03550-9;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03550-
9/S0025-5718-2020-03550-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Li%2C%20Jia;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
261930;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=891440;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=998988.

Marseglia:2021:CSF

[Mar21] Stefano Marseglia. Computing square-free polarized abelian varieties over finite fields. Mathematics of Computation, 90(328): 953–971, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03594-7;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03594-
7/S0025-5718-2020-03594-7.
pdf.

Mascot:2020:HLT

[Mas20] Nicolas Mascot. Hensel-lifting torsion points on Jacobians and Galois representations. Mathematics of Computation, 89(323): 1417–1455, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03484-1;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03484-
1/S0025-5718-2019-03484-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1040021.

Martinez:2021:NAM

[MIJR21] Jorge Albella Martínez, Sébastien Imperiale, Patrick Joly, and Jerónimo Rodríguez. Numerical analysis of a method for solving 2D linear isotropic elastodynamics with traction free boundary condition using potentials and finite elements. Mathematics of Computation, 90(330): 1589–1636, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-330/
S0025-5718-2021-03613-3;
https://www.ams.org/journals/
mcom/2021-90-330/S0025-5718-
2021-03613-3/S0025-5718-2021-
03613-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Albella%
20Martinez%2C%20Jorge;   https:
//www.ams.org/mathscinet/search/
authors.html?authorName=Imperiale%
2C%20Sebastien;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Rodriguez%
2C%20Jeronimo;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=234723.

Milio:2020:CIB

[Mil20] Enea Milio. Computing isogenies between Jacobians of curves of genus 2 and 3. Mathematics of Computation, 89(323): 1331–1364, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03486-5;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03486-
5/S0025-5718-2019-03486-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1129392.

Manns:2021:APS

[MKL21] Paul Manns, Christian Kirches, and Felix Lenders. Approximation properties of sum-up rounding in the presence of vanishing constraints. Mathematics of Computation, 90(329): 1263–1296, May 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-329/
S0025-5718-2021-03606-6;
https://www.ams.org/journals/
mcom/2021-90-329/S0025-5718-
2021-03606-6/S0025-5718-2021-
03606-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Kirches%
2C%20Christian;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Lenders%
2C%20Felix;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Manns%
2C%20Paul.

Masri:2022:DGP

[MLR22] Rami Masri, Chen Liu, and Beatrice Riviere. A discontinuous Galerkin pressure correction scheme for the incompressible Navier–Stokes equations: Stability and convergence. Mathematics of Computation, 91(336): 1625–1654, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03731-5;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03731-5/S0025-5718-2022-
03731-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Liu%
2C%20Chen;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Masri%2C%20Rami;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
661345.

Makuracki:2021:QAC

[MM21] Bartosz Makuracki and Andrzej Mróz. Quadratic algorithm to compute the Dynkin type of a positive definite quasi-Cartan matrix. Mathematics of Computation, 90(327):389– 412, January 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-327/
S0025-5718-2020-03559-5;
https://www.ams.org/journals/
mcom/2021-90-327/S0025-5718-
2020-03559-5/S0025-5718-2020-
03559-5.pdf.

Montardini:2020:STL

[MNST20] Monica Montardini, Matteo Negri, Giancarlo Sangalli, and Mattia Tani. Space–time least– squares isogeometric method and efficient solver for parabolic problems. Mathematics of Computation, 89(323):1193–1227, May 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

                                              /www.ams.org/AMSMathViewer;
                                               https://www.ams.org/journals/
                                              mcom/2020-89-323/S0025-5718-
                                              2019-03471-3;   https://www.
                                              ams.org/journals/mcom/2020-
                                              89-323/S0025-5718-2019-03471-
                                              3/S0025-5718-2019-03471-3.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Negri%2C%20Matteo;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1023408;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=1201262;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=663454.

Mossinghoff:2021:DFN

[MOT21] Michael J. Mossinghoff, Tomás Oliveira e Silva, and Timothy S. Trudgian. The distribution of k-free numbers. Mathematics of Computation, 90(328): 907–929, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-328/S0025-5718-2020-
                                              03581-9;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              328/S0025-5718-2020-03581-

9/S0025-5718-2020-03581-9.
pdf.

Melenk:2020:CVP

[MR20] J. M. Melenk and C. Rojik. On commuting p-version projection-based interpolation on tetrahedra. Mathematics of Computation, 89(321): 45–87, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03454-3;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03454-
3/S0025-5718-2019-03454-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Rojik%2C%20C.
;   https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
613978.

Magnusson:2023:CGV

[MR23] Tobias Magnusson and Martin Raum. On the computation of general vector-valued modular forms. Mathematics of Computation, 92(344):2861– 2891, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03847-9.

Murru:2024:NAA

[MR24] Nadir Murru and Giuliano Romeo. A new algorithm for padic continued fractions. Mathematics of Computation, 93(347):

1309–1331, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03890-X.

Matthes:2020:DFL

[MS20a] Daniel Matthes and Benjamin Söllner. Discretization of fluxlimited gradient flows: Γconvergence and numerical schemes. Mathematics of Computation, 89 (323):1027–1057, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03492-0;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03492-
0/S0025-5718-2019-03492-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Sollner%2C%
20Benjamin;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=722279.

Merai:2020:DSS

[MS20b] László Mérai and Igor E. Shparlinski. Distribution of short subsequences of inversive congruential pseudorandom numbers modulo 2^{t}. Mathematics of Computation, 89(322):911–922, March 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03467-1;   https://www.ams.
org/journals/mcom/2020-89-

322/S0025-5718-2019-03467-
1/S0025-5718-2019-03467-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Merai%2C%20Laszlo;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
192194.

Mercuri:2020:MFI

[MS20c] Pietro Mercuri and René Schoof.

Modular forms invariant under non-split Cartan subgroups. Mathematics of Computation, 89 (324):1969–1991, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03503-0;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03503-
0/S0025-5718-2020-03503-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Schoof%2C%
20Rene;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=1105728.

Ma:2022:EED

[MS22] Chupeng Ma and Robert Scheichl. Error estimates for discrete generalized FEMs with locally optimal spectral approximations. Mathematics of Computation, 91 (338):2539–2569, November 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03755-
8/S0025-5718-2022-03755-8.
pdf;   https://www.ams.org/

                                              mathscinet/search/authors.
                                              html?mrauthid=1251608;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=661163.

Merai:2024:DRM

[MS24] László Mérai and Igor E. Shparlinski. Distribution of recursive matrix pseudorandom number generator modulo prime powers. Mathematics of Computation, 93(347):1355–1370, May 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              347/S0025-5718-2023-03895-9.

McGown:2025:CLP

[MS25a] Kevin J. McGown and Jonathan P. Sorenson. Computation of the least primitive root. Mathematics of Computation, 94(352): 909–917, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              352/S0025-5718-2024-04003-6.

Miller:2025:LSM

[MS25b] Alexander Rossi Miller and Danny Scheinerman. Largescale Monte Carlo simulations for zeros in character tables of symmetric groups. Mathematics of Computation, 94(351): 505–515, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-03964-9.

Monsuur:2024:MRM

[MSS24] Harald Monsuur, Rob Stevenson, and Johannes Storn. Minimal residual methods in negative or fractional Sobolev norms. Mathematics of Computation, 93 (347):1027–1052, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2023-03904-7.

Mans:2023:FGF

[MSSS23] Bernard Mans, Min Sha, Igor E.

Shparlinski, and Daniel Sutantyo. Functional graphs of families of quadratic polynomials. Mathematics of Computation, 92(343): 2307–2331, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03838-8.

Murru:2020:FPA

[MT20] Nadir Murru and Lea Terracini.

On the finiteness and periodicity of the p-adic Jacobi– Perron algorithm. Mathematics of Computation, 89(326):2913– 2930, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03540-6;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03540-
6/S0025-5718-2020-03540-6.
pdf.

Marques:2022:EEC

[MT22] Diego Marques and Pavel Trojovský. Error estimates for a class of continuous Bonsetype inequalities. Mathematics of Computation, 91(337):2335– 2345, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03741-8;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03741-8/S0025-5718-2022-
03741-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Trojovsky%
2C%20Pavel;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=870578.

Maier:2022:MSN

[MV22] Roland Maier and Barbara Verfürth. Multiscale scattering in nonlinear Kerr-type media. Mathematics of Computation, 91 (336):1655–1685, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03722-4;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03722-4/S0025-5718-2022-
03722-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Verfurth%
2C%20Barbara;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1329180.

Mignotte:2024:KLF

[MV24a] Maurice Mignotte and Paul Voutier. A kit for linear forms in three logarithms. Mathematics of Computation, 93(348): 1903–1951, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2023-03908-4.

With an appendix by Michel Laurent.

Mitra:2024:PEE

[MV24b] K. Mitra and M. Vohralík. A posteriori error estimates for the Richards equation. Mathematics of Computation, 93(347): 1053–1096, May 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
347/S0025-5718-2024-03932-7.

Masdeu:2026:ECN

[MX26] Marc Masdeu and Xavier Xarles. Efficient computation of non-archimedean theta functions. Mathematics of Computation, 95(357):457–475, January 2026. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04066-3.

Ma:2021:OHM

[MZ21] Jiming Ma and Fangting Zheng.

Orientable hyperbolic 4-manifolds over the 120-cell. Mathematics of Computation, 90(331):2463– 2501, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-331/

                                              S0025-5718-2021-03625-X;
                                              https://www.ams.org/journals/
                                              mcom/2021-90-331/S0025-5718-
                                              2021-03625-X/S0025-5718-2021-
                                              03625-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Ma%
                                              2C%20Jiming;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Zheng%
                                              2C%20Fangting.

Nakatsukasa:2020:SEB

[Nak20] Yuji Nakatsukasa. Sharp error bounds for Ritz vectors and approximate singular vectors. Mathematics of Computation, 89(324): 1843–1866, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-324/S0025-5718-2020-
                                              03519-4;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              324/S0025-5718-2020-03519-
                                              4/S0025-5718-2020-03519-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=887438.

Nick:2024:NAE

[Nic24] Jörg Nick. Numerical analysis for electromagnetic scattering with nonlinear boundary conditions. Mathematics of Computation, 93 (348):1529–1568, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              348/S0025-5718-2023-03914-X.

Najman:2024:GMC

[NO24] Filip Najman and Petar Orlić.

Gonality of the modular curve X_{0}(N). Mathematics of Computation, 93(346):863–886, March 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2024-
93-346/S0025-5718-2023-03873-
X.

Notay:2020:ATG

[Not20] Yvan Notay. Analysis of two-grid methods: The nonnormal case. Mathematics of Computation, 89 (322):807–827, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03460-9;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03460-
9/S0025-5718-2019-03460-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=266926.

Notaris:2022:AGQ

[Not22]  Sotirios E. Notaris.  Anti-

Gaussian quadrature formulae of Chebyshev type. Mathematics of Computation, 91(338):2803– 2816, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03762-
5/S0025-5718-2022-03762-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Notaris%2C%
20Sotirios%20E..

Nuyens:2023:SLR

[NS23] Dirk Nuyens and Yuya Suzuki. Scaled lattice rules for integration on ℝ^{d} achieving higher-order convergence with error analysis in terms of orthogonal projections onto periodic spaces. Mathematics of Computation, 92(339): 307–347, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              339/S0025-5718-2022-03754-
                                              6/S0025-5718-2022-03754-6.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1348627;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=777310.

Najman:2023:QPB

[NV23] Filip Najman and Borna Vukorepa. Quadratic points on bielliptic modular curves. Mathematics of Computation, 92(342): 1791–1816, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              342/S0025-5718-2023-03805-4.

Ovall:2023:AIE

[OR23] Jeffrey S. Ovall and Robyn Reid. An algorithm for identifying eigenvectors exhibiting strong spatial localization. Mathematics of Computation, 92(341): 1005–1031, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              341/S0025-5718-2023-03734-

6/S0025-5718-2023-03734-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Reid%2C%20Robyn;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
728623.

Ostermann:2022:EEL

[ORS22] Alexander Ostermann, Frédéric Rousset, and Katharina Schratz. Error estimates at low regularity of splitting schemes for NLS. Mathematics of Computation, 91(333):169–182, January 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
333/S0025-5718-2021-03676-
5;   https://www.ams.org/journals/
mcom/2022-91-333/S0025-5718-
2021-03676-5/S0025-5718-2021-
03676-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Rousset%
2C%20Frederic;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=134575;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
990639.

Olshanskii:2024:EFE

[ORS24] Maxim A. Olshanskii, Arnold Reusken, and Paul Schwering. An Eulerian finite element method for tangential Navier–Stokes equations on evolving surfaces. Mathematics of Computation, 93(349):2031– 2065, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

org/journals/mcom/2024-93-
349/S0025-5718-2023-03931-X.

Olshanskii:2021:ISS

[ORZ21] Maxim A. Olshanskii, Arnold Reusken, and Alexander Zhiliakov. Inf-sup stability of the trace P_{2} − P_{1} Taylor–Hood elements for surface PDEs. Mathematics of Computation, 90(330): 1527–1555, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-330/
S0025-5718-2021-03551-6;
https://www.ams.org/journals/
mcom/2021-90-330/S0025-5718-
2021-03551-6/S0025-5718-2021-
03551-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1331093;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
147305;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=343398.

Obus:2021:SCM

[OS21] Andrew Obus and Tanush Shaska. Superelliptic curves with many automorphisms and CM Jacobians. Mathematics of Computation, 90(332):2951– 2975, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03639-X;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03639-
X/S0025-5718-2021-03639-X.
pdf;   https://www.ams.org/

mathscinet/search/authors.
html?mrauthid=678224;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=890287.

Orloski:2025:NLB

[OSS25] Bryce Joseph Orloski, Naser Talebizadeh Sardari, and Alexander Smith. New lower bounds for the Schur– Siegel–Smyth trace problem. Mathematics of Computation, 94 (354):2005–2040, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
354/S0025-5718-2024-04004-8.

Osting:2020:DGM

[OW20] Braxton Osting and Dong Wang.

A diffusion generated method for orthogonal matrix-valued fields. Mathematics of Computation, 89 (322):515–550, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03473-7;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03473-
7/S0025-5718-2019-03473-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Wang%2C%20Dong;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
876194.

Olver:2020:OPQ

[OX20] Sheehan Olver and Yuan Xu. Orthogonal polynomials in and on a quadratic surface of revolution. Mathematics of Computation, 89 (326):2847–2865, November 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03544-3;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03544-
3/S0025-5718-2020-03544-3.
pdf.

Pagani:2022:GCR

[Pag22] Lorenzo Pagani. Greenberg’s conjecture for real quadratic fields and the cyclotomic ℤ_{2}-extensions. Mathematics of Computation, 91 (335):1437–1467, May 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-335/
S0025-5718-2021-03712-6;
https://www.ams.org/journals/
mcom/2022-91-335/S0025-5718-
2021-03712-6/S0025-5718-2021-
03712-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Pagani%
2C%20Lorenzo.

Porzio:2025:TLS

[PNR25] Gian Maria Negri Porzio, Vanni Noferini, and Leonardo Robol. Tropical Laurent series, their tropical roots, and localization results for the eigenvalues of nonlinear matrix functions. Mathematics of Computation, 94(354): 1947–1975, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
354/S0025-5718-2024-04016-4.

Pan:2023:SPA

[PO23] Zexin Pan and Art B. Owen.

Super-polynomial accuracy of one dimensional randomized nets using the median of means. Mathematics of Computation, 92(340): 805–837, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03791-
1/S0025-5718-2022-03791-1.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1340663;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=134875.

Pan:2024:SPA

[PO24] Zexin Pan and Art B. Owen.

Super-polynomial accuracy of multidimensional randomized nets using the median-of-means. Mathematics of Computation, 93(349):2265–2289, September 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2024-
93-349/S0025-5718-2024-03880-
2.

Pfeiler:2020:DMM

[PP20] Carl-Martin Pfeiler and Dirk Praetorius. Dörfler marking with minimal cardinality is a linear complexity problem. Mathematics of Computation, 89(326): 2735–2752, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2020-89-326/
S0025-5718-2020-03553-4;

                                              https://www.ams.org/journals/
                                              mcom/2020-89-326/S0025-5718-
                                              2020-03553-4/S0025-5718-2020-
                                              03553-4.pdf.

Parker:2021:AFS

[PS21] Chris Parker and Jason Semeraro. Algorithms for fusion systems with applications to pgroups of small order. Mathematics of Computation, 90(331): 2415–2461, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-331/S0025-5718-2021-
                                              03634-0;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              331/S0025-5718-2021-03634-
                                              0/S0025-5718-2021-03634-0.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1048702;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=315689.

Palitta:2025:STP

[PSS25] Davide Palitta, Marcel Schweitzer, and Valeria Simoncini. Sketched and truncated polynomial Krylov subspace methods: Matrix Sylvester equations. Mathematics of Computation, 94(354): 1761–1792, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04002-4.

Peng:2025:OOE

[PSW25] Manting Peng, Zheng Sun, and Kailiang Wu. OEDG:

Oscillation-eliminating discontinuous Galerkin method for hyperbolic conservation laws. Mathematics of Computation, 94(353): 1147–1198, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03998-4.

Poor:2020:FAB

[PSY20] Cris Poor, Jerry Shurman, and David S. Yuen. Finding all Borcherds product paramodular cusp forms of a given weight and level. Mathematics of Computation, 89(325):2435– 2480, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03532-7;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03532-
7/S0025-5718-2020-03532-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=270719;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=291737;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
364614.

Platt:2021:ETP

[PT21] David J. Platt and Timothy S.

Trudgian. The error term in the prime number theorem. Mathematics of Computation, 90(328): 871–881, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-328/S0025-5718-2020-
                                              03583-2;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              328/S0025-5718-2020-03583-
                                              2/S0025-5718-2020-03583-2.
                                              pdf.

Pacetti:2022:CHC

[PT22] Ariel Pacetti and Lucas Villagra Torcomian. ℚ-curves, Hecke characters and some Diophantine equations. Mathematics of Computation, 91(338):2817– 2865, November 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2022-91-
                                              338/S0025-5718-2022-03759-
                                              5/S0025-5718-2022-03759-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Villagra%20Torcomian%
                                              2C%20Lucas;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=759256.

Pulaj:2020:CPF

[Pul20] Jonad Pulaj. Cutting planes for families implying Frankl’s conjecture. Mathematics of Computation, 89(322):829–857, March 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

                                              /www.ams.org/AMSMathViewer;
                                               https://www.ams.org/journals/
                                              mcom/2020-89-322/S0025-5718-
                                              2019-03461-0;   https://www.
                                              ams.org/journals/mcom/2020-
                                              89-322/S0025-5718-2019-03461-
                                              0/S0025-5718-2019-03461-0.
                                              pdf;   https://www.ams.org/

mathscinet/search/authors.
html?mrauthid=1182719.

Peterseim:2020:CHF

[PV20] Daniel Peterseim and Barbara Verfürth. Computational high frequency scattering from highcontrast heterogeneous media. Mathematics of Computation, 89 (326):2649–2674, November 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03529-7;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03529-
7/S0025-5718-2020-03529-7.
pdf.

Pang:2021:SCA

[PYAT21] Gang Pang, Yibo Yang, Xavier Antoine, and Shaoqiang Tang. Stability and convergence analysis of artificial boundary conditions for the Schrödinger equation on a rectangular domain. Mathematics of Computation, 90(332): 2731–2756, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-332/
S0025-5718-2021-03679-0;
https://www.ams.org/journals/
mcom/2021-90-332/S0025-5718-
2021-03679-0/S0025-5718-2021-
03679-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Pang%
2C%20Gang;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Tang%2C%20Shaoqiang;
 https://www.ams.org/mathscinet/

                                              search/authors.html?mrauthid=
                                              1183327;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=637398.

Quellmalz:2024:MDD

[QESS24] Michael Quellmalz, Peter Elbau, Otmar Scherzer, and Gabriele Steidl. Motion detection in diffraction tomography by common circle methods. Mathematics of Computation, 93(346): 747–784, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              346/S0025-5718-2023-03869-8.

Quesada-Herrera:2022:APC

[QH22] Oscar E. Quesada-Herrera. On the q-analogue of the Pair Correlation Conjecture via Fourier optimization. Mathematics of Computation, 91(337):2347–2365, September 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-337/
                                              S0025-5718-2022-03747-9;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-337/S0025-5718-
                                              2022-03747-9/S0025-5718-2022-
                                              03747-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Quesada-
                                              Herrera%2C%20Oscar%20E..

Qin:2025:LRQ

[QMSZ25] Zhenzhi Qin, Zhenyu Ming, Defeng Sun, and Liping Zhang. Low-rank quaternion tensor completion for color video inpainting via a novel factorization strategy. Mathematics of Computation, 94(355):2409– 2456, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
355/S0025-5718-2024-04025-5.

Qiu:2020:NMA

[QT20] Weifeng Qiu and Lan Tang.

A note on the Monge–Ampère type equations with general source terms. Mathematics of Computation, 89(326):2675– 2706, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03554-6;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03554-
6/S0025-5718-2020-03554-6.
pdf.

Rakvi:2024:CGM

[Rak24]  Rakvi.  A classification of

genus 0 modular curves with rational points. Mathematics of Computation, 93(348): 1859–1902, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2023-03907-2.

Reusken:2022:AFE

[Reu22] Arnold Reusken. Analysis of finite element methods for surface vector-Laplace eigenproblems. Mathematics of Computation, 91 (336):1587–1623, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-336/
                                              S0025-5718-2022-03728-5;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-336/S0025-5718-
                                              2022-03728-5/S0025-5718-2022-
                                              03728-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=147305.

MC-50240005

[Reu25] Arnold Reusken. Analysis of the Taylor–Hood surface finite element method for the surface Stokes equation. Mathematics of Computation, 94(354): 1701–1719, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04008-5.

Rickards:2022:ICF

[Ric22] James Rickards. Improved computation of fundamental domains for arithmetic Fuchsian groups. Mathematics of Computation, 91 (338):2929–2954, November 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2022-91-
                                              338/S0025-5718-2022-03777-
                                              7/S0025-5718-2022-03777-7.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=939651.

Rieger:2023:GGK

[Rie23]  Janosch Rieger.  Generalized

Gearhart–Koshy acceleration for the Kaczmarz method. Mathematics of Computation, 92(341):

1251–1272, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
341/S0025-5718-2023-03818-
2/S0025-5718-2023-03818-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=829842.

Rosenfeld:2020:HFM

[Ros20] Matthieu Rosenfeld. How far away must forced letters be so that squares are still avoidable? Mathematics of Computation, 89 (326):3057–3071, November 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03535-2;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03535-
2/S0025-5718-2020-03535-2.
pdf.

Rosenfeld:2022:ASW

[Ros22]  Matthieu Rosenfeld.  Avoid-

ing squares over words with lists of size three amongst four symbols. Mathematics of Computation, 91(337):2489– 2500, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03732-7;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03732-7/S0025-5718-2022-
03732-7.pdf;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1169147.

Rutka:2023:EEP

[RS23] Przemysław Rutka and Ryszard Smarzewski. The electrostatic equilibrium problem for classical discrete orthogonal polynomials. Mathematics of Computation, 92 (341):1331–1348, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              341/S0025-5718-2023-03808-
                                              X/S0025-5718-2023-03808-X.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=163855;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=890344.

Rosler:2024:CEL

[RS24] Frank Rösler and Alexei Stepanenko. Computing eigenvalues of the Laplacian on rough domains. Mathematics of Computation, 93(345):111–161, January 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              345/S0025-5718-2023-03827-3.

Rieder:2022:TDB

[RSM22] Alexander Rieder, Francisco-Javier Sayas, and Jens Markus Melenk. Time domain boundary integral equations and convolution quadrature for scattering by composite media. Mathematics of Computation, 91(337): 2165–2195, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03730-3;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03730-3/S0025-5718-2022-
03730-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1126143;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
613978;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=621885.

Reinke:2023:WDK

[RSS23] Bernhard Reinke, Dierk Schleicher, and Michael Stoll. The Weierstrass-Durand–Kerner root finder is not generally convergent. Mathematics of Computation, 92 (340):839–866, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03783-
2/S0025-5718-2022-03783-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1489442;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=325630;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
359328.

Rud:2022:ETN

[Rüd22] Thomas Rüd. Explicit Tamagawa numbers for certain algebraic tori over number fields. Mathematics of Computation, 91 (338):2867–2904, November 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

                                              org/journals/mcom/2022-91-
                                              338/S0025-5718-2022-03771-
                                              6/S0025-5718-2022-03771-6.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Rud%2C%20Thomas.

Saouter:2022:NRW

[Sao22] Yannick Saouter. New results for witnesses of Robin’s criterion. Mathematics of Computation, 91 (334):909–920, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03687-X;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03687-X/S0025-5718-2021-
                                              03687-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=306740.

Santos:2025:IKS

[SF25] Maria Corte-Real Santos and E. Victor Flynn. Isogenies on Kummer Surfaces. Mathematics of Computation, 94(355):2575– 2612, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              355/S0025-5718-2024-04036-X.

Sgobba:2023:DCO

[Sgo23] Pietro Sgobba. Divisibility conditions on the order of the reductions of algebraic numbers. Mathematics of Computation, 92(343): 2281–2305, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.

org/journals/mcom/2023-92-
343/S0025-5718-2023-03848-0.

Sheng:2021:UCU

[SH21] Qiwei Sheng and Cory D.

Hauck. Uniform convergence of an upwind discontinuous Galerkin method for solving scaled discrete-ordinate radiative transfer equations with isotropic scattering. Mathematics of Computation, 90(332):2645– 2669, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-332/
S0025-5718-2021-03670-4;
https://www.ams.org/journals/
mcom/2021-90-332/S0025-5718-
2021-03670-4/S0025-5718-2021-
03670-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1022258;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
748066.

Sablica:2022:BKF

[SH22] Lukas Sablica and Kurt Hornik.

On bounds for Kummer’s function ratio. Mathematics of Computation, 91(334):887–907, March 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2022-
91-334/S0025-5718-2021-03690-
X;   https://www.ams.org/journals/
mcom/2022-91-334/S0025-5718-
2021-03690-X/S0025-5718-2021-
03690-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Sablica%
2C%20Lukas;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=265899.

Sugiura:2025:EBG

[SH25] Hiroshi Sugiura and Takemitsu Hasegawa. Error bounds for Gauss–Jacobi quadrature of analytic functions on an ellipse. Mathematics of Computation, 94(351):359–379, January 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              351/S0025-5718-2024-03977-7.

Saito:2020:MAC

[SiTiY20] Asaki Saito, Jun ichi Tamura, and Shin ichi Yasutomi. Multidimensional p-adic continued fraction algorithms. Mathematics of Computation, 89(321): 351–372, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-321/S0025-5718-2019-
                                              03458-0;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              321/S0025-5718-2019-03458-
                                              0/S0025-5718-2019-03458-0.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=229479;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=306289;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              679931.

Sui:2024:NCP

[SJ24] Jiaming Sui and Junxiong Jia. Non-centered parametric variational Bayes’ approach for hierarchical inverse problems of partial differential equations. Mathematics of Computation, 93(348): 1715–1760, July 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
348/S0025-5718-2023-03906-0.

Sheng:2023:JTA

[SLN23] Zhou Sheng, Jianze Li, and Qin Ni. Jacobi-type algorithms for homogeneous polynomial optimization on Stiefel manifolds with applications to tensor approximations. Mathematics of Computation, 92(343): 2217–2245, September 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
343/S0025-5718-2023-03834-0.

Soga:2020:SVA

[Sog20] Kohei Soga. Stochastic and variational approach to finite difference approximation of Hamilton– Jacobi equations. Mathematics of Computation, 89(323): 1135–1159, May 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-323/S0025-5718-2019-
03437-3;   https://www.ams.
org/journals/mcom/2020-89-
323/S0025-5718-2019-03437-
3/S0025-5718-2019-03437-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=909684.

Sheng:2026:IEM

[SSX26] Changtao Sheng, Bihao Su, and Chenglong Xu. An implicit– explicit Monte Carlo method for semi-linear PDEs driven by α-stable Lévy process and its error estimates. Mathematics of Computation, 95(357): 263–291, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2026-95-
                                              357/S0025-5718-2024-04044-9.

Sauter:2023:ISS

[ST23] Stefan Sauter and Céline Torres. On the inf-sup stability of Crouzeix–Raviart Stokes elements in 3D. Mathematics of Computation, 92(341): 1033–1059, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              341/S0025-5718-2022-03793-
                                              5/S0025-5718-2022-03793-5.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Torres%2C%
                                              20Celine;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=313335.

Steinerberger:2021:WRK

[Ste21] Stefan Steinerberger. A weighted randomized Kaczmarz method for solving linear systems. Mathematics of Computation, 90(332): 2815–2826, November 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:

//www.ams.org/journals/mcom/
2021-90-332/S0025-5718-2021-
03644-3;   https://www.ams.
org/journals/mcom/2021-90-
332/S0025-5718-2021-03644-
3/S0025-5718-2021-03644-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=869041.

Sun:2021:LGP

[Sun21]  Wenchang Sun.  Local and

global phaseless sampling in real spline spaces. Mathematics of Computation, 90(330): 1899–1929, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-330/S0025-5718-2021-
03620-0;   https://www.ams.
org/journals/mcom/2021-90-
330/S0025-5718-2021-03620-
0/S0025-5718-2021-03620-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Sun%2C%20Wenchang.

MC-50080005

[Sun25] Zheng Sun. Reducing polynomial degree by one for innerstage operators affects neither stability type nor accuracy order of the Runge–Kutta discontinuous Galerkin method. Mathematics of Computation, 94(356): 2795–2838, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2024-04037-1.

Stevenson:2020:UPP

[SvV20] Rob Stevenson and Raymond van Venetië. Uniform preconditioners for problems of negative order. Mathematics of Computation, 89 (322):645–674, March 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-322/S0025-5718-2019-
03481-6;   https://www.ams.
org/journals/mcom/2020-89-
322/S0025-5718-2019-03481-
6/S0025-5718-2019-03481-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=van%20Venetie%
2C%20Raymond;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=310898.

Sorenson:2020:TAF

[SW20] Jonathan P. Sorenson and Jonathan Webster. Two algorithms to find primes in patterns. Mathematics of Computation, 89(324): 1953–1968, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03501-7;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03501-
7/S0025-5718-2020-03501-7.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=334195;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=903429.

Sun:2021:NAG

[SW21] Weiwei Sun and Chengda Wu.

New analysis of Galerkin-mixed FEMs for incompressible miscible flow in porous media. Mathematics of Computation, 90(327): 81–102, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-
03561-3;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03561-
3/S0025-5718-2020-03561-3.
pdf.

Sun:2025:UEE

[SW25] Weiran Sun and Li Wang. Uniform error estimate of an asymptotic preserving scheme for the Lévy–Fokker–Planck equation. Mathematics of Computation, 94 (352):681–725, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03975-3.

Schratz:2021:LRI

[SWZ21a] Katharina Schratz, Yan Wang, and Xiaofei Zhao. Lowregularity integrators for nonlinear Dirac equations. Mathematics of Computation, 90(327): 189–214, January 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-327/S0025-5718-2020-

03557-1;   https://www.ams.
org/journals/mcom/2021-90-
327/S0025-5718-2020-03557-
1/S0025-5718-2020-03557-1.
pdf.

Sun:2021:PCR

[SWZ21b] Xingping Sun, Zongmin Wu, and Xuan Zhou. On probabilistic convergence rates of stochastic Bernstein polynomials. Mathematics of Computation, 90(328): 813–830, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-328/S0025-5718-2020-
03589-3;   https://www.ams.
org/journals/mcom/2021-90-
328/S0025-5718-2020-03589-
3/S0025-5718-2020-03589-3.
pdf.

Sun:2021:OEE

[SX21] Zheng Sun and Yulong Xing. Optimal error estimates of discontinuous Galerkin methods with generalized fluxes for wave equations on unstructured meshes. Mathematics of Computation, 90 (330):1741–1772, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-330/S0025-5718-2021-
03605-4;   https://www.ams.
org/journals/mcom/2021-90-
330/S0025-5718-2021-03605-
4/S0025-5718-2021-03605-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.

html?authorName=Sun%2C%20Zheng;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
761305.

Steininger:2023:AAR

[SY23] Jakob Steininger and Sergey Yurkevich. An algorithmic approach to Rupert’s problem. Mathematics of Computation, 92 (342):1905–1929, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03831-5.

Tao:2025:LDG

[TCLS25] Qi Tao, Xiang-Ke Chang, Yong Liu, and Chi-Wang Shu. A local discontinuous Galerkin method for the Novikov equation. Mathematics of Computation, 94(354): 1603–1631, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
354/S0025-5718-2024-04018-8.

Tanguy:2025:PDS

[TFD25] Eloi Tanguy, Rémi Flamary, and Julie Delon. Properties of discrete sliced Wasserstein losses. Mathematics of Computation, 94 (353):1411–1465, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03994-7.

Tao:2025:CPC

[TG25] Zhengyu Tao and Xuejun Guo.

CM points, class numbers, and the Mahler measures of x^{3} + y^{3} + 1 − kxy. Mathematics of Computation, 94(351): 425–446, January 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
351/S0025-5718-2024-03961-3.

Tan:2025:EGS

[THW25] Jingyu Tan, Zhijian He, and Xiaoqun Wang. Extensible grid sampling for quantile estimation. Mathematics of Computation, 94 (352):763–800, March 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
352/S0025-5718-2024-03986-8.

Tang:2025:EAN

[TJW25] Peipei Tang, Bo Jiang, and Chengjing Wang. An efficient algorithm for the ℓ_{p} norm based metric nearness problem. Mathematics of Computation, 94(355): 2495–2532, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
355/S0025-5718-2024-04026-7.

MC-22170001

[Tra25] Ngoc Tien Tran. Finite element approximation for uniformly elliptic linear PDE of second order in nondivergence form. Mathematics of Computation, 94(353): 1043–1064, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03966-2.

Tao:2020:ULD

[TXS20] Qi Tao, Yan Xu, and Chi-Wang Shu. An ultraweak-local discontinuous Galerkin method for PDEs with high order spatial derivatives. Mathematics of Computation, 89(326):2753– 2783, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-326/S0025-5718-2020-
03562-5;   https://www.ams.
org/journals/mcom/2020-89-
326/S0025-5718-2020-03562-
5/S0025-5718-2020-03562-5.
pdf.

Trudgian:2025:TOE

[TY25] Timothy S. Trudgian and Andrew Yang. Toward optimal exponent pairs. Mathematics of Computation, 94(353): 1467–1502, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03993-5.

Ulas:2024:CDQ

[Ula24]  Maciej Ulas.  Construction

of diagonal quintic threefolds with infinitely many rational points. Mathematics of Computation, 93(349):2503–2511, September 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2024-93-349/
S0025-5718-2024-03953-4.

vanBommel:2020:EAI

[vBHM20] Raymond van Bommel, David Holmes, and J. Steffen Müller. Explicit arithmetic intersection theory and computation of Néron–Tate heights. Mathematics of Computation, 89(321): 395–410, January 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-321/S0025-5718-2019-
03441-5;   https://www.ams.
org/journals/mcom/2020-89-
321/S0025-5718-2019-03441-
5/S0025-5718-2019-03441-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Muller%2C%
20J.%20Steffen;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=1271408;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
972881.

vanLuijk:2023:CLP

[vLW23] Ronald van Luijk and Rosa Winter. Concurrent lines on del Pezzo surfaces of degree one. Mathematics of Computation, 92(339): 451–481, January 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
339/S0025-5718-2022-03779-
0/S0025-5718-2022-03779-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1471783;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=698553.

Voet:2026:MLO

[VSB26] Yannis Voet, Espen Sande, and Annalisa Buffa. Mass lumping and outlier removal strategies for complex geometries in isogeometric analysis. Mathematics of Computation, 95(357): 105–146, January 2026. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2026-95-
357/S0025-5718-2025-04060-2.

Walsh:1999:TCS

[Wal99] P. G. Walsh. On two classes of simultaneous Pell equations with no solutions. Mathematics of Computation, 68(225): 385–388, January 1999. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-
01048-0&u=/mcom/1999-68-225/

. See corrigendum [Wal21].

Walsh:2021:CTC

[Wal21]  P. G. Walsh.  Corrigendum

to “On two classes of simultaneous Pell equations with no solutions”. Mathematics of Computation, 90(331):2503– 2505, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-
03677-7;   https://www.ams.
org/journals/mcom/2021-90-
331/S0025-5718-2021-03677-
7/S0025-5718-2021-03677-7.

                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=230332.  See

[Wal99].

Wang:2024:CAL

[Wan24] Haiyong Wang. Convergence analysis of Laguerre approximations for analytic functions. Mathematics of Computation, 93(350): 2861–2884, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              350/S0025-5718-2024-03942-X.

MC-50080010

[Wan25] Xiao Wang. Complexity analysis of inexact cubic-regularized primal-dual methods for finding second-order stationary points. Mathematics of Computation, 94 (356):2961–3008, November 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              356/S0025-5718-2024-04029-2.

Wang:2024:CFP

[WD24] Zhaonan Wang and Yingpu Deng. Convergence, finiteness and periodicity of several new algorithms of p-adic continued fractions. Mathematics of Computation, 93 (350):2921–2942, November 2024. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              350/S0025-5718-2024-03948-0.

Wang:2020:DPS

[WDM20] Dongming Wang, Rina Dong, and Chenqi Mou. Decomposition of polynomial sets into characteristic pairs. Mathematics of Computation, 89(324): 1993–2015, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03504-2;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03504-
2/S0025-5718-2020-03504-2.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1242700;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=233946;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
906857.

Wiltshire-Gordon:2020:CEB

[WG20] John D. Wiltshire-Gordon. On computing the eventual behavior of an FI-module over the rational numbers. Mathematics of Computation, 89(326):2985– 3001, November 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2020-89-326/
S0025-5718-2020-03538-8;
https://www.ams.org/journals/
mcom/2020-89-326/S0025-5718-
2020-03538-8/S0025-5718-2020-
03538-8.pdf.

Wu:2025:FCC

[WGMZ25] Zhizhang Wu, Ivan G. Graham, Dingjiong Ma, and Zhiwen Zhang. A Filon–Clenshaw— Curtis–Smolyak rule for multidimensional oscillatory integrals with application to a UQ problem for the Helmholtz equation. Mathematics of Computation, 94 (354):1911–1946, July 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              354/S0025-5718-2024-04007-3.

Williams:2024:GKI

[WH24] David M. Williams and Qingguo Hong. Generalized Korn’s inequalities for piecewise H^{1} and H^{2} vector fields. Mathematics of Computation, 93(350):2587– 2609, November 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              350/S0025-5718-2024-03935-2.

Wang:2025:OEE

[WLF25] Kun Wang, Enlong Liu, and Xinlong Feng. Optimal error estimate of unconditionally positivitypreserving, mass-conserving and energy stable method for the Keller–Segel chemotaxis model. Mathematics of Computation, 94 (356):2761–2793, November 2025. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2025-94-
                                              356/S0025-5718-2024-04041-3.

Wang:2021:NSS

[WLL21] Xiang Wang, Junliang Lv, and Yonghai Li. New superconvergent structures developed from the finite volume element method in 1D. Mathematics of Computation, 90(329):1179–1205, May 2021.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-329/S0025-5718-2020-
03587-X;   https://www.ams.
org/journals/mcom/2021-90-
329/S0025-5718-2020-03587-
X/S0025-5718-2020-03587-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Li%2C%20Yonghai;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Lv%2C%20Junliang;   https:/
/www.ams.org/mathscinet/search/
authors.html?authorName=Wang%
2C%20Xiang.

Wang:2023:USL

[WLSZ23] Haijin Wang, Fengyan Li, Chi-Wang Shu, and Qiang Zhang. Uniform stability for local discontinuous Galerkin methods with implicit-explicit Runge– Kutta time discretizations for linear convection–diffusion equation. Mathematics of Computation, 92 (344):2475–2513, November 2023. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03842-X.

Wright:2020:ARP

[WpL20] Stephen J. Wright and Ching pei Lee. Analyzing random permutations for cyclic coordinate descent. Mathematics of Computation, 89(325):2217– 2248, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.

                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2020-89-325/S0025-5718-2020-
                                              03530-3;   https://www.ams.
                                              org/journals/mcom/2020-89-
                                              325/S0025-5718-2020-03530-
                                              3/S0025-5718-2020-03530-3.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1024505;   https:
                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=213980.

Wu:2024:GQR

[WW24] Menghan Wu and Haiyong Wang. Gaussian quadrature rules for composite highly oscillatory integrals. Mathematics of Computation, 93(346): 729–746, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2024-93-
                                              346/S0025-5718-2023-03878-9.

Wang:2021:TPA

[WWW21] Cong Wang, Jie Wu, and Qiang Wu. Totally positive algebraic integers with small trace. Mathematics of Computation, 90(331): 2317–2332, September 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/AMSMathViewer;   https:
                                              //www.ams.org/journals/mcom/
                                              2021-90-331/S0025-5718-2021-
                                              03636-4;   https://www.ams.
                                              org/journals/mcom/2021-90-
                                              331/S0025-5718-2021-03636-
                                              4/S0025-5718-2021-03636-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Wang%2C%20Cong;
                                               https://www.ams.org/mathscinet/

search/authors.html?authorName=
Wu%2C%20Jie;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Wu%
2C%20Qiang.

Wang:2022:FSA

[WX22] Yinkun Wang and Shuhuang Xiang. Fast and stable augmented Levin methods for highly oscillatory and singular integrals. Mathematics of Computation, 91 (336):1893–1923, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03725-X;
https://www.ams.org/journals/
mcom/2022-91-336/S0025-5718-
2022-03725-X/S0025-5718-2022-
03725-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Xiang%
2C%20Shuhuang;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1046342.

Wang:2022:DDE

[WY22a] Wansheng Wang and Lijun Yi.

Delay-dependent elliptic reconstruction and optimal L^{∞}(L^{2}) a posteriori error estimates for fully discrete delay parabolic problems. Mathematics of Computation, 91 (338):2609–2643, November 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2022-91-
338/S0025-5718-2022-03761-
3/S0025-5718-2022-03761-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Wang%2C%20Wansheng;
 https://www.ams.org/mathscinet/

                                              search/authors.html?mrauthid=
                                              883098.

Wu:2022:FOF

[WY22b] Yifei Wu and Fangyan Yao. A first-order Fourier integrator for the nonlinear Schrödinger equation on 𝕋 without loss of regularity. Mathematics of Computation, 91(335):1213– 1235, May 2022. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-335/
                                              S0025-5718-2021-03705-9;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-335/S0025-5718-
                                              2021-03705-9/S0025-5718-2021-
                                              03705-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Wu%
                                              2C%20Yifei;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1422543.

Wang:2023:PDP

[WY23] Larry X. W. Wang and Neil N. Y. Yang. Positivity of the determinants of the partition function and the overpartition function. Mathematics of Computation, 92 (341):1383–1402, May 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL https://www.ams.
                                              org/journals/mcom/2023-92-
                                              341/S0025-5718-2023-03810-
                                              8/S0025-5718-2023-03810-8.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Yang%2C%20Neil%
                                              20N.%20Y.;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=845775.

Wang:2022:SLR

[WZ22] Yan Wang and Xiaofei Zhao.

A symmetric low-regularity integrator for nonlinear Klein– Gordon equation. Mathematics of Computation, 91(337):2215– 2245, September 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-337/
S0025-5718-2022-03751-0;
https://www.ams.org/journals/
mcom/2022-91-337/S0025-5718-
2022-03751-0/S0025-5718-2022-
03751-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Wang%
2C%20Yan;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=1045425.

Widlund:2021:BFD

[WZSP21] O. B. Widlund, S. Zampini, S. Scacchi, and L. F. Pavarino. Block FETI-DP/BDDC preconditioners for mixed isogeometric discretizations of threedimensional almost incompressible elasticity. Mathematics of Computation, 90(330): 1773–1797, July 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2021-90-330/S0025-5718-2021-
03614-5;   https://www.ams.
org/journals/mcom/2021-90-
330/S0025-5718-2021-03614-
5/S0025-5718-2021-03614-5.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=182600;   https:

                                              //www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=291739;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              828985;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=923734.

Wang:2022:SEA

[WZZ22] Jihong Wang, Jiwei Zhang, and Chunxiong Zheng. Stability and error analysis for a second-order approximation of 1D nonlocal Schrödinger equation under DtNtype boundary conditions. Mathematics of Computation, 91(334): 761–783, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL  https://www.ams.org/
                                              journals/mcom/2022-91-334/
                                              S0025-5718-2021-03685-6;
                                              https://www.ams.org/journals/
                                              mcom/2022-91-334/S0025-5718-
                                              2021-03685-6/S0025-5718-2021-
                                              03685-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Wang%
                                              2C%20Jihong;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Zhang%
                                              2C%20Jiwei;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Zheng%
                                              2C%20Chunxiong.

Xu:2024:DVT

[XB24] Wei-Wei Xu and Zheng-Jian Bai. Double-variable trace maximization for extreme generalized singular quartets of a matrix pair: a geometric method. Mathematics of Computation, 93(349): 2331–2359, September 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
349/S0025-5718-2024-03936-4.

Xiang:2023:PEE

[XKLW23] Shuhuang Xiang, Desong Kong, Guidong Liu, and Li-Lian Wang. Pointwise error estimates and local superconvergence of Jacobi expansions. Mathematics of Computation, 92(342): 1747–1778, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03835-2.

Xing:2024:UAF

[XLL24] Xin Xing, Xiaoxu Li, and Lin Lin. Unified analysis of finite-size error for periodic Hartree–Fock and second order Møller–Plesset perturbation theory. Mathematics of Computation, 93(346): 679–727, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03877-7.

Xia:2025:SPR

[XXX25] Yu Xia, Zhiqiang Xu, and Zili Xu.

Stability in phase retrieval: Characterizing condition numbers and the optimal vector set. Mathematics of Computation, 94(356): 2931–2960, November 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
356/S0025-5718-2024-04042-5.

Yang:2020:OCS

[YCQ20] Yang Yang, Xiaofeng Cai, and Jing-Mei Qiu. Optimal convergence and superconvergence of semi-Lagrangian discontinuous Galerkin methods for linear convection equations in one space dimension. Mathematics of Computation, 89(325):2113– 2139, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03527-3;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03527-
3/S0025-5718-2020-03527-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1008921;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=1167587;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
843920.

Yin:2025:TDL

[YEHS25] Peimeng Yin, Eirik Endeve, Cory D. Hauck, and Stefan R. Schnake. Towards dynamical lowrank approximation for neutrino kinetic equations. Part I: Analysis of an idealized relaxation model. Mathematics of Computation, 94 (353):1199–1233, May 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
353/S0025-5718-2024-03997-2.

Yu:2024:FPG

[YLLO24] Jiajia Yu, Rongjie Lai, Wuchen Li, and Stanley Osher. A fast proximal gradient method and convergence analysis for dynamic mean field planning. Mathematics of Computation, 93(346): 603–642, March 2024. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2024-93-
346/S0025-5718-2023-03879-0.

Zhang:2021:NTA

[ZC21] Liping Zhang and Chiyu Chen.

A Newton-type algorithm for the tensor eigenvalue complementarity problem and some applications. Mathematics of Computation, 90(327):215–231, January 2021. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/journals/mcom/2021-
90-327/S0025-5718-2020-03558-
3;   https://www.ams.org/journals/
mcom/2021-90-327/S0025-5718-
2020-03558-3/S0025-5718-2020-
03558-3.pdf.

Zahm:2022:CDR

[ZCL^{+}22] Olivier Zahm, Tiangang Cui, Kody Law, Alessio Spantini, and Youssef Marzouk. Certified dimension reduction in nonlinear Bayesian inverse problems. Mathematics of Computation, 91(336): 1789–1835, July 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-336/
S0025-5718-2022-03737-6;
https://www.ams.org/journals/

mcom/2022-91-336/S0025-5718-
2022-03737-6/S0025-5718-2022-
03737-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1084083;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1084144;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=1106417;   https:
//www.ams.org/mathscinet/search/
authors.html?mrauthid=761311;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
869711.

Zeng:2023:NIC

[Zen23] Zhonggang Zeng. A Newton’s iteration converges quadratically to nonisolated solutions too. Mathematics of Computation, 92(344): 2795–2824, November 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
344/S0025-5718-2023-03657-2.

Zhao:2022:SCE

[Zha22] Weifeng Zhao. Strictly convex entropy and entropy stable schemes for reactive Euler equations. Mathematics of Computation, 91 (334):735–760, March 2022. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2022-91-334/
S0025-5718-2022-03721-2;
https://www.ams.org/journals/
mcom/2022-91-334/S0025-5718-
2022-03721-2/S0025-5718-2022-
03721-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Zhao%
2C%20Weifeng.

Zimmermann:2021:GAS

[Zim21]  Ralf Zimmermann.  A geo-

metric approach to subspace updates and orthogonal matrix decompositions under rankone modifications. Mathematics of Computation, 90(328): 671–688, March 2021. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  https://www.ams.org/
journals/mcom/2021-90-328/
S0025-5718-2020-03574-1;
https://www.ams.org/journals/
mcom/2021-90-328/S0025-5718-
2020-03574-1/S0025-5718-2020-
03574-1.pdf.

Zhang:2020:BAT

[ZL20] Qinghai Zhang and Zhixuan Li. Boolean algebra of twodimensional continua with arbitrarily complex topology. Mathematics of Computation, 89(325): 2333–2364, September 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-
03539-X;   https://www.ams.
org/journals/mcom/2020-89-
325/S0025-5718-2020-03539-
X/S0025-5718-2020-03539-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Li%2C%20Zhixuan;
https://www.ams.org/mathscinet/
search/authors.html?authorName=
Zhang%2C%20Qinghai.

Zhao:2023:IPV

[ZMZW23] Jikun Zhao, Shipeng Mao, Bei Zhang, and Fei Wang. The interior penalty virtual element method for the biharmonic problem. Mathematics of Computation, 92(342): 1543–1574, July 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
342/S0025-5718-2023-03828-5.

Zhang:2020:LCM

[ZWZ20] Ning Zhang, Jia Wu, and Liwei Zhang. A linearly convergent majorized ADMM with indefinite proximal terms for convex composite programming and its applications. Mathematics of Computation, 89(324): 1867–1894, July 2020. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/AMSMathViewer;   https:
//www.ams.org/journals/mcom/
2020-89-324/S0025-5718-2020-
03506-6;   https://www.ams.
org/journals/mcom/2020-89-
324/S0025-5718-2020-03506-
6/S0025-5718-2020-03506-6.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Wu%2C%20Jia;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Zhang%2C%20Liwei;   https:/
/www.ams.org/mathscinet/search/
authors.html?authorName=Zhang%
2C%20Ning.

Zhang:2025:CDP

[ZYYZ25] Lei-Hong Zhang, Linyi Yang, Wei Hong Yang, and Ya-Nan Zhang. A convex dual problem for the rational minimax approximation and Lawson’s iteration. Mathematics of Computation, 94(355):2457– 2494, September 2025. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2025-94-
355/S0025-5718-2024-04021-8.

Zhang:2023:RLB

[ZZY23] Hui Zhang, Lu Zhang, and Hao-Xing Yang. Revisiting linearized Bregman iterations under Lipschitz-like convexity condition. Mathematics of Computation, 92(340): 779–803, March 2023. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/2023-92-
340/S0025-5718-2022-03792-
3/S0025-5718-2022-03792-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Yang%2C%20Hao-
Xing;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Zhang%2C%20Hui;
 https://www.ams.org/mathscinet/
search/authors.html?authorName=
Zhang%2C%20Lu.